### v1

# Rethinking Arrow–Debreu: A New Framework for Exchange, Time, and Uncertainty

Rethinking Arrow–Debreu: A New Framework for Exchange, Time, and Uncertainty
Nizar Riane
Université Mohammed V de Rabat, Maroc111nizar.riane@gmail.com
Keywords: General equilibrium – Welfare – Network – Convergence – Quantity theory of money – Uncertainty – Open economy.
JEL Classification:
Acknowledgements: The author would like to thank Michel Lapidus and Michael Maroun for their insightful comments, which greatly contributed to improving this work.
1 Introduction
The Arrow–Debreu Model (ADM) [AD54] has long been a cornerstone of modern economic theory, providing a rigorous mathematical framework for analyzing microeconomic decisions. It offers a formal resolution to an old question posed by Léon Walras in [Wal74], leveraging advancements in the theory of multifunctions and fixed points.
While the ADM is meticulously formulated, it has not been universally embraced by economists. Criticism often targets its assumptions, including perfect competition, the existence of forward markets for every commodity and all conceivable contingencies, and the absence of money as a store of value. This has led some scholars, such as Mark Blaug, to express strong reservations. Blaug famously characterized Arrow and Debreu’s seminal paper [AD54] as marking "the beginning of what has since become a cancerous growth in the very center of microeconomics" (see [Bla98]). He further argued that "by the time we got to Arrow and Debreu, general equilibrium theory had ceased to make any descriptive claim about actual economic systems and had become a purely formal apparatus about a quasi-economy". His criticism extended to Gérard Debreu’s book, The Theory of Value [Deb59], which he deemed "probably the most arid and pointless book in the entire literature of economics".
Beyond Blaug’s critiques of its theoretical abstraction, it is important to recognize that general equilibrium theory has undergone significant refinements since the publication of Arrow and Debreu’s 1954 paper. Debreu’s The Theory of Value [Deb59] further formalized the framework, particularly in economies facing uncertainty. The role of money within the model has been extensively debated, with insightful contributions from Robert Clower [Clo67], Frank Hahn [Hah89], and Joseph M. Ostroy and Ross M. Starr [OS74] and others. The question of equilibrium convergence was explored using nontraditional techniques and more realistic approaches (see, for example, [HN62]). Additionally, time was explicitly incorporated into the system [Rad72], and market imperfections-such as equilibrium with indivisible goods-were examined to better align the theory with economic realities [Sve84].
In his book The Methodology of Economics: Or, How Economists Explain [Bla92], Mark Blaug provocatively asks whether general equilibrium theory should be regarded as an empirical theory or merely as a conceptual framework. To explore this question, one can draw an analogy with classical physics: just as Newtonian mechanics assumes the existence of a Galilean reference frame–a perfectly inertial system in which the laws of motion take their simplest form–general equilibrium theory is grounded in idealized assumptions of complete markets and perfect competition. However, in practice, the Earth is a non-inertial frame of reference: its rotation gives rise to apparent forces, such as the Coriolis effect, which must be accounted for to correctly describe physical phenomena. Likewise, real-world markets are subject to imperfections and incompleteness, forming what one might call a "non-Galilean" economic environment. In such contexts, the neat predictions of general equilibrium theory must be modified to incorporate frictions, constraints, and institutional complexities. Thus, while general equilibrium provides a powerful and elegant framework for understanding economic coordination, its empirical relevance depends on our ability to account for the deviations from its foundational assumptions.
Contemplating the ADM, one should recognize that it fundamentally relies on the concepts of theoretical demand and supply—idealized representations of preferences and resource endowments. However, these abstractions often diverge from real-world market transactions, where effective demand and supply, shaped by constraints and actual economic behavior, determine outcomes. A critical reassessment of these distinctions is essential for a deeper understanding of exchange systems and money.
In the real economy, effective demand and supply, determined by what individuals can actually transact, take precedence. Transactions reflect tangible constraints, such as budgets, production capabilities, and specifically, the other party capacities and needs, making it a bilateral consent.
Utility in real-world exchanges depends on effective allocations-what individuals receive or consume through transactions-rather than theoretical allocations. This distinction underscores the importance of addressing the outcomes of exchanges rather than unattainable aspirations. Markets are not arenas for idealized redistribution but mechanisms for facilitating feasible transactions shaped by constraints and negotiation.
Real-world exchanges also separate the roles of buying and selling. Transactions are inherently bilateral: they occur between two parties negotiating terms and quantities, rather than through collective redistribution. The ADM’s implicit assumption of a pooled tribal distribution system-akin to the split necklace problem-fails to capture the sequential and decentralized nature of actual market interactions. In practice, individuals bring goods to the market not for pooling but for specific exchanges, governed by mutually acceptable terms.
Moreover, the quantity exchanged in a transaction is inherently constrained. It is determined by the minimum that both parties can support, reflecting their respective limitations. This contrasts with theoretical models that emphasize optimal allocations without considering such practical constraints.
Equilibrium is not a pure quantity adjustment process but a dynamic interplay between prices and quantities. Individuals do not merely optimize quantities based on given prices; they actively set prices and negotiate quantities in response to their circumstances and opportunities. This dynamic adjustment process better reflects how markets operate, with equilibrium emerging from these decentralized interactions.
In the monetary sphere, the bilateral nature of exchange also provides a lens to trace the circulation of money in the economy from a local perspective. Rather than relying on a global analysis dictated by the quantitative theory of money, this approach emphasizes localized flows of currency, capturing the intricacies of individual transactions and their cumulative effects, and also the impact of the topology of the network relationships.
Under this vision, monetary policy takes on a more nuanced role, exerting exotic impacts that simultaneously influence quantities, prices, velocity, and redistribution. By addressing these interconnected factors, monetary interventions could reshape not only aggregate outcomes but also the microeconomic structures underpinning economic activity.
The Rational Expectations Hypothesis, which assumes that agents form beliefs based on the true statistical expectation of future variables, has long served as a foundational assumption in anticipation modeling. However, this framework abstracts away from key psychological and informational constraints faced by real-world decision-makers. In practice, individuals do not bet on the future by averaging over full probability distributions. Instead, their anticipations are shaped by bounded rationality, limited attention, memory constraints, and heuristic reasoning. And one should distinguish between anticipation an uncertain future and gambling over a set of lotteries. In light of these limitations, we propose a departure from the traditional use of expectations and instead define anticipation as the conditional mode—that is, the most likely realization of a variable given the agent’s information, memory, and subjective beliefs. This approach better reflects the role of instinctive judgment, focal outcomes, and perceived salience, aligning with Keynes’ notion of "animal spirits" [Key36]. By grounding anticipations in the mode rather than the mean, this formulation introduces behavioral realism without sacrificing formal rigor. It allows for agent heterogeneity, evolving beliefs, and more nuanced responses to uncertainty—features that are essential in dynamic, strategic environments.
Our paper is organized in the following scheme:
  1. 1.
    In the first part, we introduce the effective trade model, establish the key properties of the price-demand correspondence, and prove the existence of Nash equilibria. We analyze welfare properties, market convergence, and imperfections arising from indivisible goods and market topology. Additionally, we incorporate production and money, deriving the quantity equation of money, and conclude with a numerical illustration.
  2. 2.
    Next, we examine the role of time in the system, highlighting the emergence of a time market where loanable funds suppliers interact with production plans that consider time as a production factor.
  3. 3.
    In the third part, we introduce uncertainty into the system and demonstrate how trade equilibrium is achieved despite the presence of rationing in supply and demand.
  4. 4.
    Finally, we extend the analysis to open economies, exploring the emergence of exchange rates and their implications.
2 New way to consider the Consumer Problem
Consider an economy $\mathcal{E}=\left(w^{i},u^{i},\mathscr{X}^{i}\right)_{1\leqslant i\leqslant n}$ consisting of $L$ goods and $n$ consumption units. Each unit $i$ is endowed with a real-valued utility function $u^{i}$ and an initial endowment vector $w^{i}=(w^{i}_{1},\ldots,w^{i}_{L})\in\mathscr{X}^{i}$. The consumption possibilities set $\mathscr{X}^{i}$ is a closed convex subset of ${\mathbb{R}}_{+}^{L}$.
We suppose that each consumption unit (or customer, for simplicity) selects its own supply price vector (not a price taker)
$p^{i}\in P=\left\{p\mid p\in\mathbb{R}_{+}^{L},\,p\neq 0,\,\sum_{j=1}^{L}p_{j}=1\right\}$ (1)
and its potential bilateral transactions matrix (representing theoretical demands and supplies)
$X^{i}=(x^{i}_{ji},x^{i}_{ij})_{1\leqslant j\leqslant n}\in(\mathbb{R}_{+}^{L})^{2n}$ (2)
with $x^{i}_{ii}=0$, where $x^{i}_{ij}$ denotes the quantity that customer $i$ wants to sell to customer $j$, and $x^{i}_{ji}$ denotes the quantity that customer $i$ wants to buy from customer $j$, interpreted in the sense of a directed multigraph.
The effective supply (transaction) from customer $i$ to customer $j$ is given by
$q_{ij}=\min\left(x^{i}_{ij},x^{j}_{ij}\right).$ (3)
Conversely, the effective demand of customer $i$ from customer $j$ is defined as
$q_{ji}=\min\left(x^{i}_{ji},x^{j}_{ji}\right).$ (4)
Define the final holding of customer $i$ after trade as
$x^{i}=w^{i}+\sum_{j\neq i}q_{ji}-\sum_{j\neq i}q_{ij}.$ (5)
We introduce the notations $X=(X^{1},\dots,X^{n})$ and $p=(p^{1},\dots,p^{n})$. The transactions of customer $i$ must satisfy the transaction balance condition
$\sum_{j\neq i}p^{j}\cdot q_{ji}=\sum_{j\neq i}p^{i}\cdot q_{ij}.$ (6)
Along with the condition that the final holding must be non-negative $0\leqslant x^{i}$ and belongs to the feasible set $\mathscr{X}^{i}$.
Remark 2.1.
On the Budget Constraint
The transaction balance condition should not be replaced by an inequality of the form
$\sum_{j\neq i}p^{j}\cdot q_{ji}\leqslant\sum_{j\neq i}p^{i}\cdot q_{ij}.$
A strict inequality would imply that the customer accepted an exchange in which higher-valued goods were traded for lower-valued ones, which effectively equivalent to setting a price lower than $p^{i}$. One might consider a condition similar to that in the Arrow–Debreu model [AD54]:
$p^{j}\cdot x^{i}\leqslant p^{i}\cdot w^{i}.$
This is equivalent to
$p^{i}\cdot\sum_{j\neq i}q_{ji}\leqslant p^{i}\cdot\sum_{j\neq i}q_{ij}.$
However, this constraint lacks any justification in our context of exchange.
The customer problem is
$\max_{(p^{i},X^{i})\in\mathbf{B}^{i}(p^{\neq i},X^{\neq i})}u^{i}(x^{i})$ (7)
Under the budget constraint
$\mathbf{B}^{i}(p^{\neq i},X^{\neq i})$ $=\left\{(p^{i},X^{i})\in P\times({\mathbb{R}}_{+}^{L})^{2n}\ \mid\ \sum_{j\neq i}p^{j}\cdot q_{ji}=\sum_{j\neq i}p^{i}\cdot q_{ij}\ ,\ w^{i}+\sum_{j\neq i}q_{ji}-\sum_{j\neq i}q_{ij}\in\mathscr{X}^{i}\right\}$ (8)
The notation $\neq i$} refers to indices $j\in\{1,\dots,n\}$ such that $j\neq i$. An equivalent way to express the maximization problem is to redefine $x^{i}=w^{i}+\sum_{j\neq i}x^{i}_{ji}-\sum_{j\neq i}x^{i}_{ij}$ and then solve
$\max_{(p^{i},X^{i})\in\mathcal{B}^{i}(p^{\neq i},X^{\neq i})}u^{i}(x^{i})$ (9)
where the budget constraint becomes
$\begin{split}\mathcal{B}^{i}(p^{\neq i},X^{\neq i})=\left\{\right.&(p^{i},X^{i})\in P\times({\mathbb{R}}_{+}^{L})^{2n}\ \mid\ \sum_{j\neq i}p^{j}\cdot x_{ji}^{i}=\sum_{j\neq i}p^{i}\cdot x_{ij}^{i}\ ,\ w^{i}+\sum_{j\neq i}x_{ji}^{i}-\sum_{j\neq i}x_{ij}^{i}\in\mathscr{X}^{i}\ ,\ x_{ij}^{i}\leqslant x_{ij}^{j}\ ,\\ &\left.x_{ji}^{i}\leqslant x_{ji}^{j}\right\}\end{split}$ (10)
Remark 2.2.
The equivalence of problems 7 and 9 reflects the fact that the supplier has no interest in exceeding demand, and the demander has no interest in exceeding supply. This is because the exchange cannot exceed the minimum of both, making any disequilibrium in the transaction logically inconceivable. However, in the first program, a possible disequilibrium could emerge between theoretical demands and supplies.
The set-valued correspondence $\mathcal{B}^{i}(p^{\neq i},X^{\neq i})$ is non-empty, convex valued, closed and bounded; hence, it is compact valued. Moreover, it is continuous by a classical argument:
  • Upper semicontinuity: Fix $(p^{\neq i},X^{\neq i})\in P^{n-1}\times{\mathbb{R}}_{+}^{2n\times n-1}$, and let $(p^{\neq i,m},X^{\neq i,m})_{m\geqslant 1}$ be a sequence converging to $(p^{\neq i},X^{\neq i})$. Since $\mathcal{B}^{i}$ is non-empty for all $i$, there exists a sequence $(p^{i,m},X^{i,m})_{m\geqslant 1}$ with $(p^{i,m},X^{i,m})\in\mathcal{B}^{i}(p^{\neq i,m},X^{\neq i,m})$ for all $m$. Since $(p^{\neq i,m},X^{\neq i,m})\to(p^{\neq i},X^{\neq i})$, there is a closed and bounded set $\Theta\subseteq P^{n-1}\times{\mathbb{R}}_{+}^{2n\times n-1}$, such that, for some $M\geqslant 1$, all the $(p^{\neq i,m},X^{\neq i,m})$ with $m\geqslant M$, and $(p^{\neq i},X^{\neq i})$ are contained in $\Theta$. Moreover, the structure of $\mathcal{B}^{i}$ implies that all of the $(p^{i,m},X^{i,m})\in\mathcal{B}^{i}(p^{\neq i,m},X^{\neq i,m})$ for $m\geqslant M$ lies in a closed and bounded subset of $P^{n-1}\times{\mathbb{R}}_{+}^{2n\times n-1}$. Thus, for all $m\geqslant M$, all the elements of the sequence $(p^{m},X^{m})=(p^{1,m},\ldots,p^{n,m},X^{1,m},\ldots,X^{n,m})$ lie in a closed and bounded subset of $P^{n-1}\times{\mathbb{R}}_{+}^{2n\times n-1}$. By the Bolzano-Weierstrass theorem, this sequence has a convergent subsequence $(p^{m_{s}},X^{m_{s}})_{s\geqslant 1}$ with limit point $(p,X)=(p^{1},\ldots,p^{n},X^{1},\ldots,X^{n})$. And since each element of this convergent subsequence satisfies $(p^{i,m},X^{i,m})\in\mathcal{B}^{i}(p^{\neq i,m},X^{\neq i,m})$, the limit point will also have to satisfy $(p^{i},X^{i})\in\mathcal{B}^{i}(p^{\neq i},X^{\neq i})$, for all $i$.
  • Lower semicontinuity: Fix $(p^{\neq i},X^{\neq i})\in P^{n-1}\times{\mathbb{R}}_{+}^{2n\times n-1}$ and $(p^{i},X^{i})\in\mathcal{B}^{i}(p^{\neq i},X^{\neq i})$. Let $(p^{\neq i,m},X^{\neq i,m})\to(p^{\neq i},X^{\neq i})$. If $X^{i}=0$ or $p^{i}=0$, then $(p^{i,m},X^{i,m})=(p^{i},X^{i})\in\mathcal{B}^{i}(p^{\neq i,m},X^{\neq i,m})$ converges to $(p^{i},X^{i})$. If not, consider the sequence
    $(p^{i,m},X^{i,m})$ $=\displaystyle\pi_{\mathcal{B}^{i,m}}(p^{i},X^{i})$ (11)
    where
    $\begin{split}\mathcal{B}^{i,m}=\displaystyle\left\{\right.&(p^{i},X^{i})\in P\times({\mathbb{R}}_{+}^{L})^{2n}\ \mid\ \sum_{j\neq i}p^{j,m}\cdot x_{ji}=\sum_{j\neq i}p^{i,m}\cdot x_{ij}\ ,\ w^{i}+\sum_{j\neq i}x_{ji}-\sum_{j\neq i}x_{ij}\in\mathscr{X}^{i}\ ,\ x_{ij}\leqslant x^{j,m}_{ij}\ ,\\ &\left.x_{ji}\leqslant x^{j,m}_{ji}\right\}\end{split}$ (12)
    One can verify that $(p^{i,m},X^{i,m})\in\mathcal{B}^{i}(p^{\neq i,m},X^{\neq i,m})$ and it does converge to $(p^{i},X^{i})$.
Let us assume the following properties of the utility functions:
  1. a
    Continuity: $u^{i}(\cdot)$ is continuous on $\mathscr{X}^{i}$ .
  2. b
    Strict monotonicity: $u^{i}(\cdot)$ is increasing on $\mathscr{X}^{i}$ in each argument.
  3. c
    Quasi-concavity: $u^{i}(\cdot)$ is quasi-concave on $\mathscr{X}^{i}$.
Continuity ensures that consumers do not experience sudden jumps in utility, concavity implies diminishing marginal utility, and strict monotonicity expresses the principle that more is always preferred to less. We recall Berge’s Maximum Theorem from Theorem 14.1 in the appendix.
Remark 2.3.
Even under the assumption of strict concavity, the utility function $u^{i}$ remains only concave with respect to the elements of $X^{i}$. This can lead to non-uniqueness in the agents’ optimal strategies. One potential remedy involves introducing a regularization term—such as a penalty on larger values of $x$ or on the entropy induced by price fluctuations—to encourage sharp and unique solutions. However, we will not impose such conditions in the sequel, preserving the generality of the formulation. An alternative and potentially richer approach for the quantities involves incorporating preferences over the connections between agents, capturing human relationships, marketing influences, customer loyalty, and other social phenomena. This extension would reflect how external social dynamics shape individual decisions and, consequently, the equilibrium structure of the economy.
Berge’s maximum theorem ensures the existence of the price-demand correspondences $p^{i}(p^{\neq i},X^{\neq i})$ and $X^{i}(p^{\neq i},X^{\neq i})$ which are upper semi-continuous. Moreover, $p^{i}(p^{\neq i},X^{\neq i})$ and $X^{i}(p^{\neq i},X^{\neq i})$ are convex sets.
At the maximum: Suppose $x_{ji}^{i}<x_{ji}^{j}$. Then, consumer $i$ could improve their utility by increasing $x_{ji}^{i}$ through an increase in $p^{i}$ until $x_{ji}^{i}=x_{ji}^{j}$. However, increasing $x_{ij}^{i}$ may have a negative compensatory effect on utility.
One should also note that assumption b can be replaced with an alternative condition:
  1. b
    Component-wise local non-satiation: $\forall x\in\mathscr{X}^{i}$, $\forall\varepsilon>0$ and $\forall k\in\{1,\ldots,L\}$, $\exists y\in\mathscr{X}^{i}$, with $\parallel x-y\parallel=\left|x_{k}-y_{k}\right|\leqslant\varepsilon$, such that $u^{i}(x)<u^{i}(y)$.
Directional Local Non-Satiation ensures that for any given consumption bundle, no matter how small the adjustment, there always exists a perturbation in the quantity of any single good that strictly improves the agent’s utility, guaranteeing that preferences remain locally expandable in every individual dimension without requiring full monotonicity.
At the maximum, we have $x^{i,\star}=w^{i,\star}+\sum_{j\neq i}x_{ji}^{i,\star}-\sum_{j\neq i}x_{ij}^{i,\star}$, and still satisfy $x_{ji}^{i,\star}=x_{ji}^{j}$. Otherwise, suppose $0\leqslant x_{ji}^{i,\star}-x_{ji}^{j}=\delta$ and define $L_{+}=\{k\in\{1,\dots,L\}\mid\delta_{k}>0\}$. By assumption b’, for every $\varepsilon_{k}>0$, there exists $y\in\mathscr{X}^{i}$ such that $\parallel x^{i,\star}-y\parallel\leqslant\varepsilon_{k}$, and $u^{i}(x^{i,\star})<u^{i}(y)$. For each $k\in L_{+}$, one can choose $\varepsilon_{k}=\delta_{k}$ and construct $y$ by setting $\tilde{x}_{ji,k}^{i}=x_{ji,k}^{i}+\varepsilon_{k}$ and $\tilde{x}_{ji,l}^{i}=x_{ji,l}^{i}$ for $l\neq k$, while increasing $p^{i}$ and keeping $x_{ij}^{i}$ fixed, until the budget constraint $p^{j}\cdot\sum_{j\neq i}\tilde{x}_{ji}^{i}=p^{i}\cdot\sum_{j\neq i}x_{ij}^{i}$ is satisfied. Since $u^{i}(x^{i,\star})<u^{i}(y)$, this contradicts the optimality of $x^{i,\star}$.
Remark 2.4.
Note that assumption a, together with assumption c, implies assumption b. To see this, consider the one-dimensional case (fixing the other components). Let $0<x<y$. Applying DLNS to 0, we obtain that for all $0<x$, $u(0)<u(x)$, which implies that 0 is the unique global minimum of $u$ on $\mathbb{R}_{+}$. Now, on the interval $[0,x]$, continuity and quasi-concavity guarantee the existence of a global maximum $z\in[0,x]$. Suppose, by contradiction, that $z\in]0,x[$. Then, by DLNS, there exists $a\in]0,x[$ with $a\neq z$ such that $u(z)<u(a)$, contradicting the optimality of $z$. Hence, we must have $z=x$, the unique global maximum. By the same reasoning, applying DLNS to the interval $[x,y]$ implies that the global maximum $z=x$ or $z=y$. However, if $z=x$, applying DLNS at $x$ leads to a contradiction, since $x$ is the global maximum of $[0,y]$ by hypothesis. Therefore, we conclude that $z=y$ is the unique global maximum on that interval. Thus, $u$ is strictly increasing since $u(x)<u(y)$.
Remark 2.5.
Given the definition of the budget constraint $\mathcal{B}^{i}$, if we multiply every price in $p^{\neq i}$ by $\lambda>0$, the possible adjustments to maintain balance at the optimum are as follows: increase $x_{ij}^{i}$ (if possible), decrease $x_{ji}^{i}$ (both of which negatively affect utility in their respective configurations), or increase $p^{i}$. Therefore, we can conclude that $\left(X^{i}(p^{\neq i},X^{\neq i}),\lambda p^{i}(p^{\neq i},X^{\neq i})\right)$ is a solution to the modified problem for all $i$. This justify the normalization in equation 1.
An interesting feature is to compare our configuration with the Arrow–Debreu vision of the economy. If we adopt a completely endogenous price system, the consumer’s problem would be
$\tilde{u}^{i,\star}(p)=\max_{X^{i}\in\tilde{\mathcal{B}}^{i}(p)}u^{i}(x^{i}).$ (13)
Under the budget constraint
$\tilde{\mathcal{B}}^{i}(p)=\left\{X^{i}\in(\mathbb{R}_{+}^{L})^{2n}\mid p\cdot\sum_{j\neq i}x_{ji}^{i}\leqslant p\cdot\sum_{j\neq i}x_{ij}^{i}\ ,\ w^{i}+\sum_{j\neq i}x_{ji}^{i}-\sum_{j\neq i}x_{ij}^{i}\in\mathscr{X}^{i}\right\}.$ (14)
Our maximization problem could be rewritten as
$u^{i,\star}(p)$ $=\max_{X^{i}\in\mathcal{B}^{i}(p)}u^{i}(x^{i})$ (15)
where
$\mathcal{B}^{i}(p)$ $=\left\{X^{i}\in({\mathbb{R}}_{+}^{L})^{2n}\ \mid\ p\cdot\sum_{j\neq i}x_{ji}^{i}=p\cdot\sum_{j\neq i}x_{ij}^{i}\ ,\ w^{i}+\sum_{j\neq i}x_{ji}^{i}-\sum_{j\neq i}x_{ij}^{i}\in\mathscr{X}^{i}\ ,\ x_{ij}^{i}\leqslant x_{ij}^{j}\ ,\ x_{ji}^{i}\leqslant x_{ji}^{j}\right\}$ (16)
One could then remark that $\mathcal{B}^{i}(p)\subseteq\tilde{\mathcal{B}}^{i}(p)$, and that
$u^{i,\star}(p)\leqslant\tilde{u}^{i,\star}(p)$ (17)
When adopting a unique price, one can deduce the following result
Proposition 2.1.
Optimality – First Comparison
The consumer’s maximization problem is suboptimal compared to the maximum in the Arrow–Debreu model, that is, $u^{\star}\leqslant\tilde{u}^{\star}$.
Let us now recall the notion of a generalized game.
Definition 2.1.
Definition of a Generalized Game[Kre12]
A $n$-player generalized game $G=\{A^{i},\mathcal{F}^{i},u^{i}\}_{1\leqslant i\leqslant n}$ for a finite $n\in{\mathbb{N}}$ consists of, for $i=1,\ldots,n$,
  1. 1.
    A set of strategies $A^{i}$,
  2. 2.
    A constraint correspondence $\mathcal{F}^{i}\ :\ \prod_{j\neq i}A^{j}\to A_{i}$,
  3. 3.
    A payoff function $u^{i}$.
A Nash equilibrium for this generalized game is a strategy profile $(a^{i})_{1\leqslant i\leqslant n}\in\prod_{i=1}^{n}A^{i}$ such that, for $i=1,\ldots,n$
  1. 1.
    $a^{i}\in\mathcal{F}^{i}(a^{\neq i})$.
  2. 2.
    $a^{i}$ maximizes $u^{i}(\cdot,a^{\neq i})$ over $\mathcal{F}^{i}(a^{\neq i})$.
The following proposition outlines the conditions under which a Nash equilibrium exists in the generalized game.
Proposition 2.2.
Equilibrium in Effective Trade Economy[Kre12]
Suppose that $G=\{A^{i},\mathcal{F}^{i},u^{i}\}_{1\leqslant i\leqslant n}$ is a generalized game for which
  1. 1.
    Each $A^{i}$ is a non-empty, compact, convex set,
  2. 2.
    Each $\mathcal{F}^{i}$ is a continuous, non-empty valued, and convex-valued correspondence,
  3. 3.
    Each $u^{i}$ is jointly continuous in the full vector of actions and quasi-concave in $a^{i}$ for each fixed $a^{\neq i}$.
Then $G$ has a Nash equilibrium.
The proposition below ensures the existence of a Nash equilibrium of the game $G=\{\mathbf{X}^{i},\mathcal{B}^{i},u^{i}\}_{1\leqslant i\leqslant n}$, where $\mathbf{X}^{i}$ is the strategy set of player $i$ defined by $\mathcal{B}^{i}$. We will refer to this equilibrium as a transaction equilibrium.
One notable transaction equilibrium is the autarky situation: $X=(X^{1},\dots,X^{n})=0$, for all $p=(p^{1},\dots,p^{n})\in P^{n}$. Let $x\sim(p,X)$ denote an allocation $x=(x^{1},\dots,x^{n})$ derived from the realizable price-quantity matrix $(p,X)$. Based on the properties of the price-demand correspondence and Remark 2.5, the following can be deduced.
Corollary 2.3.
Homogeneity at Equilibrium
If $(p,X)$ is a transaction equilibrium of the economy, then $(\lambda p,X)$ is also a transaction equilibrium for this economy for any $\lambda>0$.
We will denote by $\mathbf{E}$ the set of transaction equilibria of the economy $\mathcal{E}$ and $\mathcal{B}=(\mathcal{B}^{1},\ldots,\mathcal{B}^{n})$. We have the following result.
Proposition 2.4.
Compactness of $\mathbf{E}$ [LRS19]
Under the assumptions a, b and c, the set $\mathcal{E}$ of transaction equilibria is non empty and compact.
We recall next the definition of Pareto optimal allocation.
Definition 2.2.
Pareto Optimal Allocation [Sta12]
Let $\mathcal{B}^{i}\subseteq\mathbb{R}^{L}$ be the feasible set of allocations of customer $i$ and $u^{i}:\mathcal{B}^{i}\to\mathbb{R}$ be the utility function of the same agent, for $i=1,\dots,n$. An allocation $x^{\star}\sim(p^{\star},X^{\star})$ such that $(p^{\star},X^{\star})\in\mathcal{B}$, is Pareto optimal if and only if
[…]
Due to the autarky situation, one could deduce the following result.
Proposition 2.5.
The first welfare theorem does not hold, that is, not all the transaction equilibria of the economy are Pareto optimal.
Definition 2.3.
Let $\mathfrak{P}(A)$ denote the set of Pareto optimal allocations within an allocation set $A$. We define $\mathfrak{P}(\mathbf{E})$ as the set of Nash–Pareto optimal allocations, also called Nash–Pareto equilibria.
It is straightforward that if an equilibrium allocation is Pareto optimal, then it is also Nash–Pareto optimal. Moreover, if at least one equilibrium allocation is Pareto optimal, the set $\mathfrak{P}(\mathbf{E})$ coincides with the set of allocations that are both transaction equilibria and Pareto optimal. We have the following result:
Proposition 2.6.
The second welfare theorem does not hold, that is, not all Pareto optimal allocations of the economy are transaction equilibria.
A counter-example could be found in Section 8. Moreover, the set of transaction equilibria and Pareto optimal solution could be disjoint.
3 KKT characterization
In this section we make the following assumption:
  1. d
    Differentiability: $u^{i}(\cdot)$ is differentiable on $\mathscr{X}^{i}$.
And for simplicity, let $\mathscr{X}^{i}={\mathbb{R}}_{+}^{L}$ and denote $x^{i}$ as $x$ and $x^{j}$ as $y$, respectively. Define the Lagrangian function $\mathcal{L}(p^{i},X,\lambda,\mu,\eta,\nu,\theta,\kappa)$ as
$\begin{split}\mathcal{L}(p^{i},X,\lambda,\mu,\eta,\nu,\theta,\kappa)&=u(x)+\lambda\left(\sum_{k=1}^{L}p^{i}_{k}\sum_{j\neq i}x_{ij,k}-\sum_{k=1}^{L}\sum_{j\neq i}p^{j}_{k}x_{ji,k}\right)+\sum_{k=1}^{L}\sum_{j\neq i}\mu_{ij,k}x_{ij,k}+\sum_{k=1}^{L}\sum_{j\neq i}\mu_{ji,k}x_{ji,k}\\ &+\sum_{k=1}^{L}\sum_{j\neq i}\eta_{ij,k}(y_{ij,k}-x_{ij,k})+\sum_{k=1}^{L}\sum_{j\neq i}\eta_{ji,k}(y_{ji,k}-x_{ji,k})\\ &+\sum_{k=1}^{L}\theta_{k}\left(w^{i}_{k}+\sum_{j\neq i}x_{ji,k}-\sum_{j\neq i}x_{ij,k}\right)+\sum_{k=1}^{L}\nu_{k}p^{i}_{k}+\kappa\left(\sum_{k=1}^{L}p^{i}_{k}-1\right),\end{split}$ (18)
The KKT conditions for this problem are as follows, for all $i,j,k$:
$\begin{split}&0\leqslant X\leqslant Y\ ,\ 0\leqslant p^{i}\ ,\ \sum_{k=1}^{L}p^{i}_{k}=1\ ,\ \sum_{j\neq i}p^{j}\cdot x_{ji}=p^{i}\cdot\sum_{j\neq i}x_{ij}\ ,\ 0\leqslant w^{i}+\sum_{j\neq i}x_{ji}-\sum_{j\neq i}x_{ij}\ ,\\ &\ (\mu_{ij,k}\ ,\ \eta_{ij,k}\ ,\ \mu_{ji,k}\ ,\ \eta_{ji,k}\ ,\ \nu_{k}\ ,\ \theta_{k})\geqslant 0.\end{split}$ (19)
and for all $i,j,k$:
$\begin{split}\mu_{ij,k}x_{ij,k}&=0\ ,\ \mu_{ji,k}x_{ji,k}=0\ ,\ \nu_{k}p^{i}_{k}=0\ ,\ \eta_{ij,k}(y_{ij,k}-x_{ij,k})=0\ ,\\ \ \eta_{ij,k}(y_{ij,k}-x_{ij,k})&=0\ ,\ \theta_{k}(w^{i}_{k}+\sum_{j\neq i}x_{ji,k}-\sum_{j\neq i}x_{ij,k})=0\ ,\ \nu_{k}p^{i}_{k}=0.\end{split}$ (20)
We have the following stationary condition
$\frac{\partial\mathcal{L}}{\partial p^{i}_{k}}$ $=\lambda\sum_{j\neq i}x_{ij,k}+\nu_{k}+\kappa=0$
$\frac{\partial\mathcal{L}}{\partial x_{ij,k}}$ $=-\frac{\partial u}{\partial x_{k}}+\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}=0$
$\frac{\partial\mathcal{L}}{\partial x_{ji,k}}$ $=\frac{\partial u}{\partial x_{k}}-\lambda p^{j}_{k}+\mu_{ji,k}-\eta_{ji,k}+\theta_{k}=0\,.$
We obtain the following situations:
$\sum_{k=1}^{L}p^{i}_{k}\sum_{j\neq i}x_{ij,k}=\sum_{k=1}^{L}\sum_{j\neq i}p^{j}_{k}x_{ji,k}$: thus,
$\lambda\sum_{j\neq i}x_{ij,k}$ $=-\nu_{k}-\kappa$
$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}>0$
$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{j}_{k}-\mu_{ji,k}+\eta_{ji,k}-\theta_{k}>0\,.$
This implies, for $i\neq j\neq l$:
$\lambda(p^{i}_{k}-p^{j}_{k})$ $=\eta_{ij,k}+\eta_{ji,k}-(\mu_{ij,k}+\mu_{ji,k})$
$\mu_{ij,k}-\eta_{ij,k}$ $=\mu_{il,k}-\eta_{il,k}$
$\eta_{ji,k}-\mu_{ji,k}$ $=\eta_{li,k}-\mu_{li,k}\,.$
  1. 1.
    If $\lambda=0$, then $\mu_{ij,k}=\eta_{ji,k}>0$, which implies $x_{ij,k}=0$ and $x_{ji,k}=y_{ji,k}$ for all $i,j,k$. This is only possible if $x_{ij,k}=x_{ji,k}=y_{ji,k}=0$, a situation that will be excluded in the remainder of the analysis.
  2. 2.
    $p^{i}_{k}>p^{j}_{k}$: then
    $\eta_{ij,k}+\eta_{ji,k}>(\mu_{ij,k}+\mu_{ji,k})\,.$
    The possible ways (or combination) are
    1. (a)
      $\eta_{ij,k}>\mu_{ji,k}\geqslant 0$ and $\mu_{ij,k}=0$: $x_{ij,k}=y_{ij,k}$.
    2. (b)
      $\eta_{ji,k}>\mu_{ij,k}\geqslant 0$ and $\mu_{ji,k}=0$: $x_{ji,k}=y_{ji,k}$.
  3. 3.
    $p^{j}_{k}>p^{i}_{k}$: then
    $\eta_{ij,k}+\eta_{ji,k}<(\mu_{ij,k}+\mu_{ji,k})\,.$
    The possible ways (or combination) are
    1. (a)
      $0\leqslant\eta_{ji,k}<\mu_{ij,k}$ and $\eta_{ij,k}=0$: $x_{ij,k}=0$.
    2. (b)
      $0\leqslant\eta_{ij,k}<\mu_{ji,k}$ and $\eta_{ji,k}=0$: $x_{ji,k}=0$.
  4. 4.
    $p^{j}_{k}=p^{i}_{k}\geqslant 0$: then
    $\eta_{ij,k}+\eta_{ji,k}=(\mu_{ij,k}+\mu_{ji,k})\,.$
    $\eta_{ij,k}$, $\eta_{ji,k}$, $\mu_{ij,k}$ and $\mu_{ji,k}$ cannot all be positive (a situation previously isolated). The possible ways (or combinations) are
    1. (a)
      $\mu_{ij,k}=\eta_{ji,k}>0$ and $\eta_{ij,k}=\mu_{ji,k}=0$: $x_{ij,k}=0$ and $x_{ji,k}=y_{ji,k}$.
    2. (b)
      $\mu_{ji,k}=\eta_{ij,k}>0$ and $\eta_{ji,k}=\mu_{ij,k}=0$: $x_{ji,k}=0$ and $x_{ij,k}=y_{ij,k}$.
    3. (c)
      $\eta_{ij,k}=\eta_{ji,k}=\mu_{ij,k}=\mu_{ji,k}=0$ and the constraints on $X$ are not satiated.
Now let us analyze a Nash equilibrium situation: Consider the market of good $k$. Given a customer $i$, let designate $\mathbb{E}^{+}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}>p^{i}_{k}\}$, $\mathbb{E}^{-}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}<p^{i}_{k}\}$ and $\mathbb{E}^{=}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}=p^{i}_{k}\}$.
  1. 1.
    For $j\in\mathbb{E}^{-}_{k}$ (resp. $j\in\mathbb{E}^{+}_{k}$), the KKT condition for both customers implies one of the three situations:
    1. (a)
      $x^{i}_{ij,k}=x^{j}_{ij,k}=0$ and $x^{i}_{ji,k}=x^{j}_{ji,k}$.
    2. (b)
      $x^{i}_{ij,k}=x^{j}_{ij,k}$ and $x^{i}_{ji,k}=x^{j}_{ji,k}=0$.
    3. (c)
      $x^{i}_{ij,k}=x^{j}_{ij,k}=0=x^{j}_{ji,k}=x^{i}_{ji,k}$.
  2. 2.
    For $j\in\mathbb{E}^{=}_{k}$, the Nash equilibrium condition implies:
    1. (a)
      $0\leqslant x^{i}_{ij,k}=x^{j}_{ij,k}$ and $0\leqslant x^{i}_{ji,k}=x^{j}_{ji,k}$.
Remark 3.1.
To avoid the indeterminacy caused by the situation $\mathbb{E}^{=}_{k}$, particularly the occurrence of unnecessary transactions in both directions ($0<x^{i}_{ij,k}$ and $0<x^{i}_{ji,k}$) when prices are equal, one could impose the condition $x^{i}_{ij}\cdot x^{i}_{ji}=0$ for all $i,j$. This condition ensures that no simultaneous flows occur in opposite directions and will be implicitly assumed when considering the joint feasible set to guarantee boundedness.
4 Convergence
Next, we present three convergence processes toward Pareto optimal and Nash–Pareto states.
4.1 Gradient direction
Suppose the economy starts in the autarky situation $X=0$ for all $p\in P$, or any other non-equilibrium situation. Define $U(x)=\sum_{i=1}^{n}u^{i}(x^{i})$. A way to improve his situation, a player (customer) could decide to update his strategy $(p^{i},X^{i})$ in the direction of increasing utility (the gradient direction), without breaking the budget constraint (projected gradient). Such information could be observed in the magnitude of demand and supply expressed by the other players.
More precisely, let $\left(p,X\right)=\left(p^{1},\ldots,p^{n},X^{1},\ldots,X^{n}\right)$ be the price-quantity matrix, and let us recall the budget set expression
$\begin{split}\mathcal{B}^{i}(p,X)=\left\{\right.&(p^{i},X^{i})\in P\times({\mathbb{R}}_{+}^{L})^{2n}\ \mid\ \sum_{j\neq i}p^{j}\cdot x_{ji}^{i}=\sum_{j\neq i}p^{i}\cdot x_{ij}^{i}\ ,\ w^{i}+\sum_{j\neq i}x_{ji}^{i}-\sum_{j\neq i}x_{ij}^{i}\in\mathscr{X}^{i}\ ,\\ &\left.x_{ij}^{i}\leqslant x_{ij}^{j}\ ,\ x_{ji}^{i}\leqslant x_{ji}^{j}\right\}\,.\end{split}$ (21)
The define
$\mathcal{B}(p,X)$ $=\prod_{i=1}^{n}\mathcal{B}^{i}(p,X)\,.$ (22)
We introduce the following assumption:
  1. c
    Concavity: $u^{i}(\cdot)$ is concave on $\mathscr{X}^{i}$.
During the interaction process, each player adjusts their strategy in the direction of the supergradient, as the utility functions $u^{i}$ are concave, meaning
$\left(p_{t+1},X_{t+1}\right)$ $=\pi_{\mathcal{B}(p_{t+1},X_{t+1})}\left(\left(p_{t},X_{t}\right)+\mu_{t}\,\partial U(x_{t})\right)\,$, (23)
where $\partial U$ is the supergradient of $U$ with respect to $(p,X)$ and $0<\mu_{t}$ is a step size. The projection $\pi_{\mathcal{B}(p_{t+1},X_{t+1})}(z)$ refers to the unique closest element to $z$ in the feasible set.
A classical proof of the projected supergradient descend convergence could be found in [Sho98]. The convergence point $\left(p^{\star},X^{\star}\right)=\left(p^{1,\star},X^{1,\star},\ldots,p^{n,\star},X^{n,\star}\right)$ satisfies
$\left(p^{\star},X^{\star}\right)$ $=\pi_{\mathcal{B}(p^{\star},X^{\star})}\left(\left(p^{\star},X^{\star}\right)+\mu_{t}\,\partial U(x^{\star})\right)\,.$ (24)
The concave nature of $u^{i}$ and the convexity of $\mathcal{B}^{i}$, for all $i$, ensure the existence of a convergence point $(p^{\star},X^{\star})$. This point represents a situation where no increase in utilities is possible given the constraints (optimality under constraint conditions). It is Pareto optimal since no collective improvement is possible under the budget constraint. However, this construction does not guarantee a stable monotonicity of the utilities during the process, and a customer may be forced to abandon a better situation for the benefit of the community. Moreover, by Proposition 2.6, the optimal situation may not be a transaction equilibrium.
4.2 Non-tâtonnement process
In this configuration, we assume that at each step $t$, the customer proceeds with the exchange provided that a Pareto improvement is possible, that is,
[…]
At each step $t$, the new endowment of each agent $i$ becomes $w^{i}_{t}=x^{i}_{t-1}=w^{i}_{t-1}+\sum_{j\neq i}x^{i}_{ji,t-1}-\sum_{j\neq i}x^{i}_{ij,t-1}$, and the sequence of utilities $(u^{1}_{t},\dots,u^{n}_{t})_{t\in\mathbb{N}}$ is non-decreasing and bounded by $\left(u^{1}\left(\sum_{i=1}^{n}w^{i}_{t}\right),\dots,u^{n}\left(\sum_{i=1}^{n}w^{i}_{t}\right)\right)$, so it converges to a Pareto optimal allocation (with null Nash equilibrium) since
[…]
4.3 Non-tâtonnement process with Pareto suboptimality
Regarding Proposition 2.6, instead of forcing the customer to proceed with Pareto optimal exchanges, we require that the exchange is in addition a Nash equilibrium, if there is one. At each step $t$, the customers has non-decreasing utilities (since each one is maximizing their utility, and the status quo situation of autarky is always possible). The new endowment of each agent $i$ is given as before by $w^{i}_{t}=x^{i}_{t-1}=w^{i}_{t-1}+\sum_{j\neq i}x^{i}_{ji,t-1}-\sum_{j\neq i}x^{i}_{ij,t-1}$, and the sequence of utilities $(u^{1}_{t},\dots,u^{n}_{t})_{t\in\mathbb{N}}$ is non-decreasing and bounded by $\left(u^{1}\left(\sum_{i=1}^{n}w^{i}_{t}\right),\dots,u^{n}\left(\sum_{i=1}^{n}w^{i}_{t}\right)\right)$, so it converges to an autarky equilibrium corresponding to a Nash–Pareto allocation.
5 Indivisible goods
In the presence of indivisible goods, the existence of equilibrium may be questioned. However, the continuity of the utility function and the compactness of the budget set ensure that a solution to the maximization problem exists. Berge’s theorem still guarantees the upper semicontinuity of the price-demand correspondence, as $\mathcal{B}^{i}(p^{\neq i},X^{\neq i})$ remains compact and continuous, following the same reasoning as in Section 2.
Notably, the equilibrium set is nonempty and includes the autarky situation. However, convergence to a more favorable outcome, as described in Section 4, is no longer guaranteed, since the existence of a nontrivial equilibrium is not assured.
Note that the convex framework in Section 2 can be interpreted as a mixed-strategy version of the game, where the game is played repeatedly, and each player’s strategy follows a probability distribution while satisfying the transaction balance condition.
6 Market topology
What happens when there are barriers to exchange, and the bilateral exchange graph is incomplete? This occurs when some customers do not have access to all suppliers and vice versa, due to factors such as wholesale-retail inadequacy, geographic barriers, or information asymmetry.
In this situation, the economy $\mathcal{E}=\left(w^{i},u^{i},\mathscr{X}^{i},C^{i}\right)_{1\leqslant i\leqslant n}$ will be characterized with exchange capacities $C^{i}=(c^{i}_{ji},c^{i}_{ij})_{1\leqslant j\leqslant n}\in({\mathbb{R}}_{+}^{L})^{2n}$ such that $X^{i}\leqslant C^{i}$. Define the set of topological constraints $\mathfrak{T}^{i}=\left\{X^{i}\in({\mathbb{R}}_{+}^{L})^{2n}\ \mid\ X^{i}\leqslant C^{i}\right\}$. The new customer problem is
$\max_{(p^{i},X^{i})\in\mathcal{B}^{i}(p^{\neq i},X^{\neq i})\cap\mathfrak{T}^{i}}u^{i}(x^{i})$
The introduction of these new constraints alters the problem, and a direct comparison with the initial formulation leads to the following result (the middleman imperfection).
Proposition 6.1.
Optimality – Second Comparison
The topologically constrained maximization problem is suboptimal compared to the maximum in the initial effective trade model.
The new configuration does not prevent the existence of a transaction equilibrium. However, in the topologically constrained setting, the equilibrium is suboptimal compared to the unrestricted case.
7 Topological Influences on Consumer Choice and Market Coordination
In the following, we address the problem of demand indeterminacy by analyzing consumer preferences within an exchange network. Our focus is on exchange processes structured by network interactions: each agent $i$ may hold intrinsic preferences over goods, as captured by classical demand theory (see, for instance, [MCWG95]), while also developing affinities toward specific trading partners. These affinities—shaped by factors such as marketing, customer loyalty, market positioning, or interpersonal relationships—can lead agents to favor transactions with certain peers over others. We proceed to formalize and examine this layered preference structure and explore its implications for decentralized exchange.
Given a binary relation $R$ on a set $S$, we recall the following properties:
  1. 1.
    Reflexivity: $\forall x\in S,\ xRx$.
  2. 2.
    Transitivity: $\forall x,y,z\in S,\ (xRy\ \wedge\ yRz)\Rightarrow xRz$.
  3. 3.
    Completeness: $\forall x,y\in S,\ x\neq y\Rightarrow(xRy\ \vee\ yRx)$.
Now, define the binary relation $\succeq$ on a set of alternatives $Z$, where $x\succeq y$ is interpreted as "$x$ is at least as good as yy". From this, we derive two related notions:
  1. 1.
    Strict preference: $x\succ y$ if and only if $(x\succeq y)\wedge\neg(y\succeq x)$, meaning $x$ is strictly preferred to $y$.
  2. 2.
    Indifference: $x\sim y$ if and only if $(x\succeq y)\wedge(y\succeq x)$, meaning $x$ and $y$ are considered equally preferable.
Definition 7.1.
Rational Preferences
A preference relation $\succeq$ on a set $S$ is said to be rational if it satisfies the following two properties:
  • Completeness: For all $x,y\in S$, either $x\succeq y$ or $y\succeq x$ (or both).
  • Transitivity: For all $x,y,z\in S$, if $x\succeq y$ and $y\succeq z$, then $x\succeq z$.
We now introduce the concept of a utility function that represents a preference relation:
Definition 7.2.
Utility Function
A function $u:S\to\mathbb{R}$ is called a utility function representing the preference relation $\succeq$ if, for all $x,y\in S$,
$x\succeq y\ \Rightarrow\ u(x)\geqslant u(y).$
It is well known that rational preferences can be represented by a utility function:
Proposition 7.1.
[MCWG95]
A preference relation $\succeq$ can be represented by a utility function only if it is rational.
Definition 7.3.
Continuous Preferences
A preference relation $\succeq$ on $S$ is said to be continuous if for every pair of sequences $(x_{m})$ and $(y_{m})$ in $S$ such that $x_{m}\succeq y_{m}$ for all $m$, and $x_{m}\to x$, $y_{m}\to y$, it follows that $x\succeq y$.
The following fundamental result is due to Debreu:
Proposition 7.2.
Debreu’s Theorem [Kre12]
If a continuous function $u$ represents $\succeq$, then $\succeq$ is continuous. Conversely, if $\succeq$ is continuous, then there exists a continuous utility function $u$ that represents it.
Definition 7.4.
Convex Preferences
Let $S$ be a convex subset of a vector space. A preference relation $\succeq$ on $S$ is said to be:
  1. 1.
    Convex if for all $x,y\in S$ such that $x\succeq y$, and for all $\alpha\in[0,1]$, we have
    $\alpha x+(1-\alpha)y\succeq y.$
  2. 2.
    Strictly convex if for all $x,y\in S$, with $x\neq y$ and $x\succeq y$, and for all $\alpha\in(0,1)$, we have
    $\alpha x+(1-\alpha)y\succ y.$
  3. 3.
    Semi-strictly convex if they are convex and for all $x,y\in S$ such that $x\succ y$, and for all $\alpha\in(0,1)$, we have
    $\alpha x+(1-\alpha)y\succ y.$
Convexity is similarly reflected in the properties of utility representations:
Proposition 7.3.
[Kre12]
  1. 1.
    If preferences $\succeq$ are represented by a concave utility function $u$, then $\succeq$ is convex. If $u$ is strictly concave, then $\succeq$ is strictly convex.
  2. 2.
    Suppose that $u$ represents preferences $\succeq$. Then:
    • $u$ is quasi-concave if and only if $\succeq$ is convex.
    • $u$ is strictly quasi-concave if and only if $\succeq$ is strictly convex.
    • $u$ is semi-strictly quasi-concave if and only if $\succeq$ is semi-strictly convex.
Let us now consider the two sets $\{1,\dots,n\}$ and $\{1,\dots,L\}$, and define the corresponding preference relations $\succeq_{n}$ and $\succeq_{L}$. Natural questions arise when these relations are considered simultaneously, i.e., when a consumer is faced with a decision to choose between goods owned by different agents. A first question that comes to mind is: which of these relations takes priority? Is it always the case? Can an agent prefer a good solely because it was marketed by a particular agent? What happens when considering the dimensions of preferences? Do the usual smooth properties such as continuity and convexity still hold?
Let us now approach this by discussing separable preferences:
Definition 7.5.
Separable Preferences
Let $J_{1},\dots,J_{N}$ partition $\{1,\dots,L\}$. That is, $J_{m}\cap J_{n}=\emptyset$ for $n\neq m$, and $J_{1}\cup\dots\cup J_{N}=\{1,\dots,L\}$. Preferences $\succeq$ are strongly separable in $J_{1},\dots,J_{N}$ if for every $K=J_{n_{1}},\dots,J_{n_{k}}$, for some set of indices $\{n_{1},\dots,n_{k}\}$ drawn from $\{1,\dots,L\}$, $(x_{K},x_{K^{c}})\succeq(x^{\prime}_{K},x_{K^{c}})$ for some $x_{K^{c}}$ implies that $(x_{K},x^{\prime}_{K^{c}})\succeq(x^{\prime}_{K},x^{\prime}_{K^{c}})$ for all $x^{\prime}_{K^{c}}$.
We have the following result:
Proposition 7.4.
[Kre12]
Suppose preferences $\succeq$ are continuous and strongly separable in $J_{1}$ through $J_{N}$. Suppose further that $\succeq$ is nontrivial on at least three of the commodity index sets $J_{1}$ through $J_{N}$ (Therefore, $N\geqslant 3$ is certainly required). Then we can find continuous functions $u_{n}:\mathbb{R}_{+}^{J_{n}}\to\mathbb{R}$ such that
$u(x)$ $=\sum_{n=1}^{N}u_{n}(X_{J_{n}})$
is a utility representation of $\succeq$. Conversely, if preferences are represented by a utility function $u$ taking the previous form, then preferences $\succeq$ are strongly separable.
One might suppose some sort of separability between the preferences of agents and the preferences of goods, and imagine a utility function of the form
$\tilde{u}(X)$ $=u(x)+v(X)\,.$
To guarantee strict concavity and the uniqueness of choice, we must ensure that $v$ is strictly concave.
Definition 7.6.
Equivalent Classes of Utility
Define $T(X)=\left(w^{i}+\sum_{j\neq i}x_{ji}-\sum_{j\neq i}x_{ij}\right)_{1\leqslant i\leqslant n}$, and the kernel set $\ker(T)=\left\{X\in\mathbb{R}^{2nL}:T(X)=0\right\}$. A point $\dot{x}\in\mathbb{R}^{2nL}/\ker(T)$ is a representative of an equivalence class of allocations that all correspond to the same utility level of good consumption $u$.
The utility function can be written as
$\tilde{u}(X)$ $=u(T(X))+v(X)\,.$
where $u:\mathbb{R}_{+}^{L}\to\mathbb{R}$ is a continuous and concave function, and $v:\mathbb{R}_{+}^{nL}\to\mathbb{R}$ is a continuous and strictly concave function. Strict concavity of $v$ represents a desire for diversification by the partners. By Debreu’s Theorem 7.2 and Proposition 7.3, $\tilde{u}(X)$ represents continuous and strictly convex preferences $\tilde{\succeq}$.
Given this construction, what does $X\tilde{\succeq}Y$ mean for $X,Y\in\mathbb{R}^{nL}$?
  1. 1.
    Better net consumption bundle: $u(T(X))\geqslant u(T(Y))$ and $v(X)=v(Y)$.
  2. 2.
    Better net consumption bundle: $u(T(X))=u(T(Y))$ and $v(X)\geqslant v(Y)$.
  3. 3.
    Better net bundle, but worse flow pattern: $u(T(X))\geqslant u(T(Y))$ and $v(X)\leqslant v(Y)$.
  4. 4.
    Worse net bundle, but better flow pattern: $u(T(X))\leqslant u(T(Y))$ and $v(X)\geqslant v(Y)$.
In practice, the function $v$ can be made more explicit. For example, one could consider $v$ to be some measure of centrality in the network, as there are many possibilities. From a trade perspective, one could imagine that the seller, depending on their position in the market, could have different preferences. For instance, a large seller might seek to gain more influence in the market and thus maximize their spectral influence (see [Ria24]), while another agent might aim to maximize their prestige by forming relationships with larger sellers and thus maximize their eigenvector centrality. For more details on centrality measures, consult [BJT23]. Marketing and social determinants can also influence the form of $v$.
Dealing with pricing, one should also consider the agent’s perception of reality. An agent may choose a price vector $p\in P$, where $P$ is the $L$-dimensional unit simplex. However, the multitude of possible price choices can lead to a loss of decision-making uniqueness. To address this, one could imagine a strictly concave component to the pricing decision, say $h:\mathbb{R}_{+}^{L}\to\mathbb{R}$, so that the utility of the customer is given by
$\tilde{u}(p,X)$ $=u(T(X))+v(X)+h(p)\,.$
A possible choice for $h(p)$ is:
$h(p)=-\frac{1}{2}\sum_{j\neq i}\sum_{k=1}^{L}(p_{k}^{j}-p_{k})^{2}\,.$
This represents a natural tendency to avoid excessively high prices relative to the market (the crowd).
Under those new assumptions, Corollary 14.2 applies and the price-demand function is a single-valued continuous function.
8 Numerical simulation: Exchange with indivisible goods
In the following, we construct an economy consisting of two markets for indivisible goods, labeled $l=1,2$, and three consumers, denoted by $i=a,b,c$. Each consumer $i$ is endowed with an initial wealth $w^{i}=(w^{i}_{1},w^{i}_{2})$ and has a constant elasticity of substitution (CES) utility function of the form $u^{i}(x^{i}_{1},x^{i}_{2})=\left(\alpha_{i}(x^{i}_{1})^{r}+(1-\alpha_{i})(x^{i}_{2})^{r}\right)^{\frac{1}{r}}$. The allocation of consumer $i$ is defined as $x^{i}=(x^{i}_{1},x^{i}_{2})=\left(w^{i}_{1}+\sum_{j\neq i}x^{i}_{ji,1}-\sum_{j\neq i}x^{i}_{ij,1},\quad w^{i}_{2}+\sum_{j\neq i}x^{i}_{ji,2}-\sum_{j\neq i}x^{i}_{ij,2}\right)$. For numerical analysis, we set the parameters as follows:
$w^{a}=(3,1),\quad w^{b}=(2,2),\quad w^{c}=(1,3),\quad r=0.3,\quad\alpha_{a}=0.2,\quad\alpha_{b}=0.4,\quad\alpha_{c}=0.8.$
We begin with a counter-example to disprove Proposition 2.6. In the case of a continuous good, we can compute a Pareto-optimal allocation, which is given by the following table:
$(u^{a},u^{b},u^{c})$ $(p^{a}_{1},p^{b}_{1},p^{c}_{1})$ $(x_{12,1},x_{13,1},x_{23,1},x_{21,1},x_{31,1},x_{32,1})$ $(x_{12,2},x_{13,2},x_{23,2},x_{21,2},x_{31,2},x_{32,2})$
$(4.03,0,3.68)$ $(0.49,0.73,0)$ $(1.03,1.12,3.03,0,0,0)$ $(0,0.66,0,5.09,0,3.09)$
Table 1: Pareto-optimal allocation in the continuous good case
However, this solution does not constitute a Nash equilibrium.
Now, consider the discrete case. For each feasible quantity vector $X\in\mathcal{B}^{a}\cap\mathcal{B}^{b}\cap\mathcal{B}^{c}$, we assign a unique price vector (although multiple price vectors may correspond to the same quantity vector—for instance, in the case of the null solution, any price in $[0,1]$ satisfies the conditions). The total number of feasible solutions is $3334$, of which $771$ are Pareto optimal.
We present the set of Nash equilibria, including the price vectors $(p^{a}_{1},p^{b}_{1},p^{c}_{1})$, the allocation of good 1 $(x_{12,1},x_{13,1},x_{23,1},x_{21,1},x_{31,1},x_{32,1})$, the allocation of good 2 $(x_{12,2},x_{13,2},x_{23,2},x_{21,2},x_{31,2},x_{32,2})$, and the corresponding utilities $(u^{a},u^{b},u^{c})$:
$(u^{a},u^{b},u^{c})$ $(p^{a}_{1},p^{b}_{1},p^{c}_{1})$ $(x_{12,1},x_{13,1},x_{23,1},x_{21,1},x_{31,1},x_{32,1})$ $(x_{12,2},x_{13,2},x_{23,2},x_{21,2},x_{31,2},x_{32,2})$
$(1.29,2,1.29)$ $(0.27,0,1)$ $(0,0,0,0,0,0)$ $(0,0,0,0,0,0)$
$(1.29,2,1.29)$ $(0.23,0.23,0.23)$ $(0,0,0,0,0,0)$ $(0,1,0,1,0,1)$
$(1.29,2,1.29)$ $(0.65,0.65,0.65)$ $(0,0,0,0,0,0)$ $(1,0,1,0,1,0)$
$(1.29,2,1.29)$ $(0.22,0.22,0.22)$ $(0,1,0,1,0,1)$ $(0,0,0,0,0,0)$
$(1.90,2,1.90)$ $(1,0.79,0)$ $(0,3,0,0,0,0)$ $(0,0,0,0,3,0)$
$(2,2,2)$ $(1,0,0)$ $(0,1,0,0,0,0)$ $(0,0,0,0,1,0)$
$(2,2,2)$ $(0.34,0.34,0.66)$ $(1,0,1,0,0,0)$ $(0,0,0,0,1,0)$
$(2,2,2)$ $(0.59,0.41,0.41)$ $(0,1,0,0,0,0)$ $(0,0,0,1,0,1)$
$(1.29,2.02,2)$ $(0,1,0)$ $(0,0,1,0,0,0)$ $(0,0,0,0,0,1)$
$(1.29,2.02,2)$ $(0.26,0.74,0.26)$ $(0,0,1,0,0,0)$ $(1,0,0,0,1,0)$
$(1.29,2.02,2)$ $(0.94,0.94,0.07)$ $(0,1,0,1,0,0)$ $(0,0,0,0,0,1)$
(2.47,2,2.47)(2.47,2,2.47)* (1,0,0)(1,0,0)* (0,2,0,0,0,0)(0,2,0,0,0,0)* (0,0,0,0,2,0)(0,0,0,0,2,0)*
(2.47,2,2.47)(2.47,2,2.47)* (0.61,0.39,0.39)(0.61,0.39,0.39)* (0,2,0,0,0,0)(0,2,0,0,0,0)* (0,0,0,2,0,2)(0,0,0,2,0,2)*
(2.47,2,2.47)(2.47,2,2.47)* (0.57,0.43,0.43)(0.57,0.43,0.43)* (0,2,0,0,0,0)(0,2,0,0,0,0)* (0,0,0,1,1,1)(0,0,0,1,1,1)*
(2.47,2,2.47)(2.47,2,2.47)* (0.51,0.51,0.49)(0.51,0.51,0.49)* (2,0,2,0,0,0)(2,0,2,0,0,0)* (0,0,0,0,2,0)(0,0,0,0,2,0)*
(2,2.02,2.47)(2,2.02,2.47)* (0.97,0.97,0.03)(0.97,0.97,0.03)* (0,1,1,0,0,0)(0,1,1,0,0,0)* (0,0,0,0,1,1)(0,0,0,0,1,1)*
(2,2.02,2.47)(2,2.02,2.47)* (0.53,0.53,0.47)(0.53,0.53,0.47)* (0,2,0,1,0,0)(0,2,0,1,0,0)* (0,0,0,0,1,1)(0,0,0,0,1,1)*
(2,2.02,2.47)(2,2.02,2.47)* (0.58,0.58,0.42)(0.58,0.58,0.42)* (1,0,2,0,0,0)(1,0,2,0,0,0)* (0,0,0,0,1,1)(0,0,0,0,1,1)*
Table 2: Nash equilibria in the discrete goods case: Pareto-optimal equilibria are highlighted in red, while asterisk is used to mark Nash–Pareto equilibria.
There are eighteen Nash equilibria corresponding to six utility profiles, seven of which are Nash–Pareto equilibria (and Pareto optimal).
Now, suppose we introduce topological constraints—for example, players 1 and 3 cannot trade. Under these restrictions, the number of feasible solutions is reduced to $662$, of which $135$ are Pareto optimal. However, the set of Nash equilibria is reduced to two situations, none of them is Pareto optimal.
$(u^{a},u^{b},u^{c})$ $(p^{a}_{1},p^{b}_{1},p^{c}_{1})$ $(x_{12,1},x_{13,1},x_{23,1},x_{21,1},x_{31,1},x_{32,1})$ $(x_{12,2},x_{13,2},x_{23,2},x_{21,2},x_{31,2},x_{32,2})$
$(1.29,2,1.29)$ $(0.27,0,1)$ $(0,0,0,0,0,0)$ $(0,0,0,0,0,0)$
(1.29,2.02,2)(1.29,2.02,2)* (0.87,1,0)(0.87,1,0)* (0,0,1,0,0,0)(0,0,1,0,0,0)* (0,0,0,0,0,1)(0,0,0,0,0,1)*
Table 3: Nash equilibria in the discrete goods case under topological constraints: Nash–Pareto equilibria are marked with an asterisk.
This numerical example is crucial for understanding the price vector $p$. In our model, price is subjective: a null price $p^{i}_{k}$ for agent $i$ indicates that agent $i$ is not involved in transactions of good $k$. In other words, an agent assigns a price only to goods they intend to trade. This is natural, as price is a market-related concept determined by supply and demand forces. It should not be confused with intrinsic value. [^2] The distinction between use value and exchange value can be traced back to Adam Smith [Smi76]. Furthermore, the topology plays a crucial role in shaping equilibria, influencing welfare losses, and highlighting imperfections caused by intermediaries.
9 Production
In this section, we consider an economy $\mathcal{E}=\left(w^{i_{d}},u^{i_{d}},\mathscr{X}^{i_{d}},\mathscr{Y}^{i_{s}}\right)_{1\leqslant{i_{d}}\leqslant n_{d},1\leqslant{i_{s}}\leqslant n_{s}}$ consisting of $L$ goods. The economy includes $n_{d}$ consumption units, each characterized by an endowment $w^{i_{d}}$, a utility function $u^{i_{d}}$, and a consumption set $\mathscr{X}^{i_{d}}$. Additionally, there are $n_{s}$ production units, each defined by its production-possibility set $\mathscr{Y}^{i_{s}}$. The total number of agents in the economy is $n=n_{d}+n_{s}$.
The producer problem is to maximize the profit
$\max_{(p^{i_{s}},X^{i_{s}})\in\mathcal{P}^{i_{s}}(p^{\neq{i_{s}}},X^{\neq{i_{s}}})}\pi^{i_{s}}(p^{i_{s}},X^{i_{s}})=\sum_{j\neq{i_{s}}}p^{i_{s}}\cdot x_{{i_{s}}j}^{i_{s}}-\sum_{j\neq{i_{s}}}p^{j}\cdot x_{j{i_{s}}}^{i_{s}}$ (25)
under the budget constraint
$\mathcal{P}^{i_{s}}(p^{\neq{i_{s}}},X^{\neq{i_{s}}})$ $=\left\{(p^{i_{s}},X^{i_{s}})\in P\times({\mathbb{R}}_{+}^{L})^{2n}\ \mid\ \sum_{j\neq{i_{s}}}x^{i_{s}}_{{i_{s}}j}-\sum_{j\neq{i_{s}}}x^{i_{s}}_{j{i_{s}}}\in\mathscr{Y}^{i_{s}}\ ,\ x_{{i_{s}}j}^{i_{s}}\leqslant x_{{i_{s}}j}^{j}\ ,\ x_{j{i_{s}}}^{i_{s}}\leqslant x_{j{i_{s}}}^{j}\right\}\,.$ (26)
Here, $\mathscr{Y}^{i_{s}}$ is the production-possibility set of producer $i_{s}$}. It satisfies the classical assumptions:
  1. 1.
    No free lunch: $\mathscr{Y}^{i_{s}}\cap{\mathbb{R}}_{+}^{L}\subseteq\{0\}$.
  2. 2.
    Free disposal: if $y^{i_{s}}\in\mathscr{Y}^{i_{s}}$ and $y^{\prime i_{s}}\leqslant y^{i_{s}}$ then $y^{\prime i_{s}}\in\mathscr{Y}^{i_{s}}$.
  3. 3.
    Possibility of inaction: $0\in\mathscr{Y}^{i_{s}}$.
  4. 4.
    Convexity: $\mathscr{Y}^{i_{s}}$ is a convex set of ${\mathbb{R}}^{L}$.
  5. 5.
    Closedness: $\mathscr{Y}^{i_{s}}$ is a closed set of ${\mathbb{R}}^{L}$.
We allow each consumption unit $i_{d}$} to receive a share $\pi_{{i_{s}}{i_{d}}}$ of the $i_{s}$}th production unit’s profit, where
$\sum_{j=1}^{n_{d}}\pi_{{i_{s}}j}=\pi^{i_{s}}.$ (27)
Thus, the new transaction condition of each consumer $i_{d}$} is
$\sum_{j\neq{i_{d}}}p^{j}\cdot x^{i_{d}}_{j{i_{d}}}=\sum_{j\neq{i_{d}}}p^{i_{d}}\cdot x^{i_{d}}_{{i_{d}}j}+\sum_{l=1}^{n_{s}}\pi_{l{i_{d}}}.$ (28)
Given the properties of the sets $\mathcal{B}$ and $\mathcal{P}$, the application of Berge’s maximum Theorem 14.1 and the generalized game Proposition 2.2 ensures the existence of solutions to both the consumer and producer problems, with an upper semicontinuous supply-price (demand-price) correspondence, and the existence of a transaction equilibrium in the economy with production.
10 Exchange with money
Let us introduce money. We consider a monetary economy in which money functions both as a medium of exchange and a store of value. The economy consists of $L+1$ markets, corresponding to $L$ goods plus money, with each market containing $n$ consumption units (no production).
10.1 Equilibrium with money
Recall that the effective transaction of goods from customer $i$ to customer $j$ is defined as $q_{ij}=\min\left(x^{i}_{ij},x^{j}_{ij}\right)$, it generates an equivalent monetary flow from customer $j$ to customer $i$: $m_{ji}=p^{i}\cdot q_{ij}$. Conversely, the effective demand of customer $i$ from customer $j$ is given by $q_{ji}=\min\left(x^{i}_{ji},x^{j}_{ji}\right)$, with an equivalent monetary flow in the opposite direction: $m_{ij}=p^{j}\cdot q_{ji}$.
As before, following the alternative definition of the customer problems 9 and 10, define the final holding of goods for customer $i$ after trade as $x^{i}=w^{i}+\sum_{j\neq i}x^{i}_{ji}-\sum_{j\neq i}x^{i}_{ij}$, and the final holding of money as $m^{i}=\underline{m}^{i}-\sum_{j\neq i}m_{ij}+\sum_{j\neq i}m_{ji}$. Here, $w^{i}$ and $\underline{m}^{i}$ represent customer $i$’s initial holdings of goods and money, respectively.
The realistic nature of transactions makes the Clower cash-in-advance hypothesis unnecessary [Clo67], since each individual only demands goods that they can afford, and every transaction is equivalent to a monetary flow. When analyzing the monetary flows, the customer satisfies the following transaction balance condition
$\sum_{j\neq i}p^{j}\cdot x_{ji}^{i}+\sum_{j\neq i}m_{ji}$ $=\sum_{j\neq i}p^{i}\cdot x_{ij}^{i}+\sum_{j\neq i}m_{ij}\,.$ (29)
along with the following monetary counterpart identities
$\sum_{j\neq i}p^{j}\cdot x_{ji}^{i}=\sum_{j\neq i}m_{ij}\quad,\quad\sum_{j\neq i}p^{i}\cdot x_{ij}^{i}=\sum_{j\neq i}m_{ji}\,.$ (30)
Remark 10.1.
Equation 29 is the equivalent of the classical transaction balance condition in a price system $\tilde{p}$, with a unique price $\tilde{p}_{L+1}$ for money, accepted by all customers. That is,
$\sum_{j\neq i}\tilde{p}^{j}\cdot x_{ji}^{i}+\sum_{j\neq i}\tilde{p}_{L+1}\,m_{ji}$ $=\sum_{j\neq i}\tilde{p}^{i}\cdot x_{ij}^{i}+\sum_{j\neq i}\tilde{p}_{L+1}\,m_{ij}\,.$ (31)
The price system $p$ is derived by normalizing each individual price by the price of money, which we will discuss later in more details.
The consumer’s utility maximization problem in a monetary economy is expressed as follows
$\max_{\displaystyle(p^{i},X^{i})\in\mathcal{M}^{i}(p^{\neq i},X^{\neq i})}u_{i}(x^{i},m^{i})\,.$ (32)
The utility function $u^{i}$ retains its previously specified properties and is also assumed to depend on the real cash balance $\frac{m^{i}}{\mathbb{P}(p)}$, where $\mathbb{P}(p)$ is a price index. The variable $m^{i}$ is constrained to lie within a closed convex subset $\mathscr{M}^{i}$ of $\mathbb{R}$$_{+}$.
To prove the existence of equilibrium in this monetary economy, one could apply a simplex method to the prices and then normalize the money price to one, leveraging Corollary 2.3. However, the following points should be noted:
  1. 1.
    The money price must be strictly positive to permit normalization to one.
  2. 2.
    Variations in equilibrium prices across different customers create inconsistencies in the value of money between individuals, especially when choosing the numéraire (a global unit of account).
To resolve these issues, an additional condition must be imposed. Since money functions as a universally accepted medium of exchange within the economy, it should have the same positive price for all customers. Thus, we start with a system of prices $\tilde{p}$ and impose the condition that
$\tilde{p}^{1}_{L+1}=\tilde{p}^{2}_{L+1}=\dots=\tilde{p}^{n}_{L+1}=\tilde{p}_{L+1}>0\,.$ (33)
Under this new condition, the customer’s problem is given by
$\max_{\displaystyle(\tilde{p}^{i},X^{i})\in\mathcal{M}^{i}(\tilde{p}^{\neq i},X^{\neq i})}u_{i}(x^{i},m^{i})\,.$ (34)
The customer monetary budget constraint becomes
$\begin{split}\mathcal{M}^{i}(\tilde{p}^{\neq i},X^{\neq i})=\Bigg\{&(\tilde{p}^{i},X^{i})\in P\times(\mathbb{R}_{+}^{L})^{2n}\ \Bigg|\ \tilde{p}^{j}\cdot x_{ji}^{i}=\tilde{p}^{j}_{L+1}\,m_{ij}\ ,\ \tilde{p}^{i}\cdot x_{ij}^{i}=\tilde{p}^{i}_{L+1}\,m_{ji}\ ,\ w^{i}+\sum_{j\neq i}x_{ji}^{i}-\sum_{j\neq i}x_{ij}^{i}\in\mathscr{X}^{i}\ ,\\ &\underline{m}^{i}-\sum_{j\neq i}m_{ij}+\sum_{j\neq i}m_{ji}\in\mathscr{M}^{i}\ ,\ x_{ij}^{i}\leqslant x_{ij}^{j}\ ,\ x_{ji}^{i}\leqslant x_{ji}^{j}\ ,\ \tilde{p}^{i}_{L+1}=\tilde{p}^{j}_{L+1}\ ,\ j\neq i\Bigg\}\,.\end{split}$ (35)
Note that the set-valued correspondence $\mathcal{M}^{i}$ is non-empty, convex-valued, and compact-valued, since the corresponding set is both closed and bounded. The continuity of the utility function $u^{i}$ ensures the existence of the price-demand correspondence. The classical results still hold, allowing us to invoke Theorem 2.2 to establish the existence of equilibrium and Corollary 2.3 to normalize the equilibrium price system by setting $p^{1}_{L+1}=\dots=p^{n}_{L+1}=1$.
Why must $\tilde{p}_{L+1}>0$? Suppose instead that $\tilde{p}_{L+1}=0$. In this case, the equations $\tilde{p}^{j}\cdot x_{ji}^{i}=\tilde{p}_{L+1},m_{ij}$ and $\tilde{p}^{i}\cdot x_{ij}^{i}=\tilde{p}_{L+1},m_{ji}$ for all $i,j$ would imply a null system of prices, resulting in an economy where all goods are free. Given the increasing nature of the utility function, customers would have an incentive to increase their effective demands indefinitely without violating the budget constraint. This would lead to unbounded demands, making the existence of transaction equilibrium impossible in an economy with a finite supply.
By treating money as the $(L+1)$th good with a unitary price, one can observe that $\mathcal{M}^{i}(p^{\neq i},X^{\neq i})\subseteq\mathcal{B}^{i}(p^{\neq i},X^{\neq i})$ for all $i$, leading to the following result.
Proposition 10.1.
Optimality – Third Comparison
The monetary customer maximization problem is suboptimal relative to the maximum achieved in the effective trade model.
A final remark concerns the following identity, which holds for all $i$
$\sum_{j\neq i}p^{j}\cdot x_{ji}^{i}+\sum_{j\neq i}m_{ji}$ $=\sum_{j\neq i}p^{i}\cdot x_{ij}^{i}+\sum_{j\neq i}m_{ij}\,.$
By considering the net flows we get the local net quantity of money equation
$\sum_{j\neq i}\left(p^{j}\cdot x_{ji}^{i}-p^{i}\cdot x_{ij}^{i}\right)$ $=\sum_{j\neq i}\left(m_{ij}-m_{ji}\right)\,.$ (36)
If we sum the identity $m^{i}=\underline{m}^{i}+\sum_{j\neq i}\left(m_{ji}-m_{ij}\right)$ over $i$, we get the money market equilibrium condition (demand equals supply)
$M=\sum_{i=1}^{n}m^{i}=\sum_{i=1}^{n}\underline{m}^{i}=\underline{M}\,.$ (37)
And if we consider one direction flows and sum over $i$ we get the quantity of money equation
$p\cdot X\cdot\mathbf{1}=\sum_{i=1}^{n}\sum_{j\neq i}p^{j}\cdot x_{ji}^{i}=\sum_{i=1}^{n}\sum_{j\neq i}m_{ij}=\underline{M}\,v\,.$ (38)
where $v$ is the money velocity $v=\displaystyle\frac{\sum_{i=1}^{n}\sum_{j\neq i}m_{ij}}{\underline{M}}$.
10.2 Marginal effects
In this section, we analyze the effects of variations in prices and money on the economy. Given a price matrix $p$, what happens when $p^{j}_{k}$ increases for some customer $j$ and some real good $k$? From equations 29 and 30 one could expect one (or a combination) of the following reactions of the customer to occur, depending on the configuration of the economy:
  1. 1.
    Money disbursement: Decrease of $m^{i}$ causing $u^{i}$ to decrease.
  2. 2.
    Income effect: Decrease of $x^{i}_{ji,k}$ causing $u^{i}$ to decrease.
  3. 3.
    Substitute seller: Decrease of $x^{i}_{ji,k}$ and increase of $x^{i}_{li,k}$ for $l\neq j$.
  4. 4.
    Substitute product: Decrease of $x^{i}_{ji,k}$ and increase of $x^{i}_{ji,\neq k}$ for $j\neq i$.
  5. 5.
    Money hoarding: Decrease of $x^{i}_{ji,k}$ and increase of $m^{i}$.
  6. 6.
    Substitute other products (sell): $x^{i}_{ji,k}$ is kept and $x^{i}_{ij,\neq k}$ increases.
  7. 7.
    Inflation effect: $x^{i}_{ji,k}$ is kept and $p^{i}$ increases.
  8. 8.
    Substitute other products (buy): $x^{i}_{ji,k}$ is kept and $x^{i}_{ji,\neq k}$ decreases.
Now, consider an increase $\underline{m}_{i}+\delta_{m}>\underline{m}_{i}$ in the money endowment of an individual $i$. With the customer becoming wealthier (distribution disruption), the individual’s monetary budget balance suggests three conceivable equivalent effects (even at equilibrium!) depending on their utility:
  1. 1.
    Hoarding: $m^{i}$ absorbs the increase in $m$, and nothing changes in the economy (reducing money velocity).
  2. 2.
    Quantity adjustment: The customer decides either to reduce $x_{ij}^{i}$ or to increase $x_{ji}^{i}$ (if other agents are interested in holding money), leading to an effect on trade.
  3. 3.
    Price adjustment: Reducing $p^{i}$ to maintain balance has no impact on utility. Other customers could also reduce the relative price of money by increasing their own commodity prices to maintain balance (inflation effect).
One should also highlight the redistribution effect of this policy in creating inequalities.
11 Intertemporal equilibrium
Incorporating time into the system introduces new subtleties. Next, we describe a deterministic dynamic economy:
$\mathcal{E}=\left(w^{i_{d}}_{t},u^{i_{d}}_{t},\mathscr{X}^{i_{d}}_{t},\mathscr{M}^{i_{d}}_{t},\mathscr{Y}^{i_{s}}_{t},C^{i_{d}}_{t},C^{i_{s}}_{t},F^{i_{f}}_{t}\right)_{1\leqslant i_{d}\leqslant n_{d},1\leqslant i_{s}\leqslant n_{s},1\leqslant i_{f}\leqslant n_{f},t_{0}\leqslant t\leqslant T},$
where $t$ represents discrete time periods. As usual, we define $n=n_{d}+n_{s}+n_{f}+1$ as the total number of agents, consisting of $n_{d}$ consumers, $n_{s}$ suppliers, $n_{f}$ financial institutions and a central bank $\mathfrak{C}$.
Each individual is endowed with an initial holdings vector $W^{i}=\left((w^{i}(t_{0}),\underline{m}^{i}(t_{0})),\dots,(w^{i}(T),\underline{m}^{i}(T))\right)$, which represents the quantities of commodities and money received at the start of each period. The function $u^{i}_{t}$ denotes the utility of consumer $i$ at time $t$, which we assume to be continuous and increasing.
We denote by $p^{i}=(p^{i}(t_{0}),\dots,p^{i}(T))$ the price vectors for consumer $i$ at each period (with the price of money normalized to $1$ at equilibrium). At time $t$, the matrix of potential goods transactions is given by $X^{i}(t)=(x^{i}_{ji}(t),x^{i}_{ij}(t))_{1\leqslant j\leqslant n_{d}+n_{s},\,j\neq i,\,t_{0}\leqslant t\leqslant T}$, while the exchange capacities are represented as $C^{i}_{t}=(c^{i}_{ji}(t),c^{i}_{ij}(t))_{1\leqslant j\leqslant n_{d}+n_{s},\,t_{0}\leqslant t\leqslant T}\in(\mathbb{R}_{+}^{L})^{2n}$. The individual money holding at time $t$ is denoted by $m^{i}(t)$. We reintroduce the set of exchange topological constraints as $\mathfrak{T}^{i}=\left\{X^{i}\in(\mathbb{R}_{+}^{L})^{2(n_{d}+n_{s})}\ \middle|\ X^{i}(t)\leqslant C^{i}_{t}\right\}$.
Similarly, we define the financial transactions matrix for each agent $i$ as
$D^{i}(t)=(d^{i}_{ji}(t),d^{i}_{ij}(t))_{1\leqslant j\leqslant n_{f},\,j\neq i,\,t_{0}\leqslant t\leqslant T}\,$,
with financial capacities given by $F^{i}_{t}=(f^{i}_{ji}(t),f^{i}_{ij}(t))_{1\leqslant j\leqslant n_{f},\,t_{0}\leqslant t\leqslant T}\in(\mathbb{R}_{+}^{L})^{2n}$. Finally, we introduce the set of financial topological constraints: $\mathfrak{D}^{i}=\left\{D^{i}\in(\mathbb{R}_{+}^{L})^{2n_{f}}\ \middle|\ D^{i}(t)\leqslant F^{i}(t)\right\}$.
We introduce $0\leqslant r^{i}(t)\leqslant 1$ to be the interest rate of agent $i$, while $P$ represents the $(L+1)$-simplex. In equilibrium, the money price is normalized to one for each period $t$.
11.1 Transaction equilibrium in a simple financial market
In the presence of a financial system, the interest rate serves as the price of time, determined by the dynamics of supply and demand. Each agent $j$ can incur a debt $d^{j}(t)$ at time $t$, arising from trade deficits and outstanding debts from previous periods. At time $t$, agent $j$ may submit a loanable funds demand $d^{j}_{{i_{f}}j}(t)$ or a loanable funds supply $d^{j}_{j{i_{f}}}(t)$ to a financial institution $i_{f}$}. These satisfy the identities
$[d^{j}(t)]^{+}=\sum_{i=1}^{n_{f}}d^{j}_{ij}(t),\quad\text{(resp. }[d^{j}(t)]^{-}=\sum_{i=1}^{n_{f}}d^{j}_{ji}(t)\text{),}$ (39)
where $\cdot$$^{+}$ and $\cdot$$^{-}$ denote the positive and negative parts, respectively. The debt $d^{i}(t)$ for a consumer (or for a producer, excluding profit shares) can be expressed in detail as
$\begin{split}d^{i}(t)&=\sum_{j=1}^{n_{f}}\left[d^{i}_{ji}(t-1)-d^{i}_{ij}(t-1)\right]\\ &=\left(\sum_{j\neq i}p^{j}(t)\cdot x_{ji}^{i}(t)-\sum_{j\neq i}p^{i}(t)\cdot x_{ij}^{i}(t)-\sum_{j=1}^{n_{s}}\pi_{ji,t}-\sum_{l=1}^{n_{f}}\pi_{li,t}\right)\\ &+\sum_{j=1}^{n_{f}}\left[(1+r^{j}(t))d^{i}_{ji}(t-1)-(1+r^{i}(t))d^{i}_{ij}(t-1)\right]\\ &=p^{i}_{L+1}(t)\left(\sum_{j\neq i}m_{ij}(t)-\sum_{j\neq i}m_{ji}(t)\right)+\sum_{j=1}^{n_{f}}\left[(1+r^{j}(t))d^{i}_{ji}(t-1)-(1+r^{i}(t))d^{i}_{ij}(t-1)\right]\end{split}$ (40)
Under the no-Ponzi condition
$d^{i}(T)$ $\leqslant 0.$ (41)
The interest rate $r^{i}(t)$ represents the price set by player $i$ at time $t$ for lending funds, where the $(L+2)$ market corresponds to the market for loanable funds. Each consumption unit $i_{d}$ receives a share of the profits from both production units and financial institutions.
Let $r^{i}(t_{0})=0$ for all $i$. The financial institution’s problem is
$\max_{\displaystyle(r^{i_{f}},D^{i_{f}})\in\mathcal{F}^{i_{f}}(r^{\neq{i_{f}}},D^{\neq{i_{f}}})\cap\mathfrak{D}^{i_{f}}}\ \sum_{t=t_{0}}^{T}\beta^{i_{f}}(t)\,\pi^{i_{f}}_{t}(r^{i_{f}},D^{i_{f}})\,$, (42)
where $\pi^{i_{f}}_{t}(r^{i_{f}},D^{i_{f}})=r^{i_{f}}(t)\sum_{j\neq{i_{f}}}d^{i_{f}}_{{i_{f}}j}(t-1)-\sum_{j\neq{i_{f}}}r^{j}(t)\,d^{i_{f}}_{j{i_{f}}}(t-1)$, and $\beta^{i}(t)=\prod_{k=t_{0}}^{t}\frac{1}{1+r^{i}(k)}$ is the discount factor of the agent $i$. The financial institution’s constraint is
$\begin{split}\mathcal{F}^{i_{f}}(r^{\neq{i_{f}}},D^{\neq{i_{f}}})=\Bigg\{&(r^{i_{f}},D^{i_{f}})\in\left([0,1]\times(\mathbb{R}_{+}^{L})^{2n}\right)^{T+1}\ \Bigg|\ d^{i_{f}}(T)\leqslant 0\ ,\ \sum_{j\neq{i_{f}}}d^{i_{f}}_{{i_{f}}j}(t)\leqslant g_{t}\left({d^{i_{f}}(t)}^{-}\right),\\ &d^{i_{f}}_{j{i_{f}}}(t)\leqslant d^{j}_{j{i_{f}}}(t),\quad d^{i_{f}}_{{i_{f}}j}(t)\leqslant d^{j}_{j{i_{f}}}(t)\Bigg\}\,.\end{split}$ (43)
We allow interbank flows $(d^{i}_{ji}(t),d^{i}_{ij}(t),d^{j}_{ji}(t),d^{j}_{ij}(t))_{1\leqslant t\leqslant T,\ 1\leqslant i,j\leqslant n_{f},\ i\neq j}$ and flows to and from the central bank $(d^{i}_{ci}(t),d^{i}_{ic}(t),d^{c}_{ci}(t),d^{c}_{ic}(t))_{1\leqslant t\leqslant T,\ 1\leqslant i\leqslant n_{f}}$. Here, $g_{t}$ represents a money creation rule (e.g., compulsory reserves, credit supervision, etc.) at time $t$, determined by the central bank $\mathfrak{C}$. We assume that $g_{t}(\cdot)$ is continuous.
The producer now confronts two key decisions: how to produce and how to exchange. The new optimization problem is
$\max_{(p^{i_{s}},r^{i_{s}},X^{i_{s}},y^{i_{s}})\in\mathcal{P}^{i_{s}}(p^{\neq{i_{s}}},r^{\neq{i_{s}}},X^{\neq{i_{s}}},D^{\neq{i_{s}}})\cap\mathfrak{T}^{i_{s}}\cap\mathfrak{D}^{i_{s}}}\sum_{t=t_{0}}^{T}\beta^{i_{s}}(t)\,\pi^{i_{s}}_{t}(p^{i_{s}},r^{i_{s}},X^{i_{s}})\,$, (44)
where
[…]
The production constraint is
$\begin{split}\mathcal{P}^{i_{s}}(p^{\neq{i_{s}}},r^{\neq{i_{s}}},X^{\neq{i_{s}}},D^{\neq{i_{s}}})=\Bigg\{&(p^{i_{s}},r^{i_{s}},X^{i_{s}},y^{i_{s}})\in\left(P\times[0,1]\times(\mathbb{R}_{+}^{L})^{2(n_{d}+n_{s})}\times{\mathbb{R}}^{L}\right)^{T+1}\ \mid\ d^{i_{s}}(T)\leqslant 0\ ,\\ &\left(y^{i_{s}}(t+\tau),-\tau\right)\in\mathscr{Y}^{i_{s}}_{t}\ ,\ S^{i_{s}}_{t}\in\mathscr{S}^{i_{s}}_{t}\ ,\ x_{{i_{s}}j}^{i_{s}}(t)\leqslant x_{{i_{s}}j}^{j}(t)\ ,x_{j{i_{s}}}^{i_{s}}(t)\leqslant x_{j{i_{s}}}^{j}(t)\ ,\\ &d_{{i_{s}}j}^{i_{s}}(t)\leqslant d_{{i_{s}}j}^{j}(t)\ ,\ d_{j{i_{s}}}^{i_{s}}(t)\leqslant d_{j{i_{s}}}^{j}(t)\ ,\ p^{i_{s}}_{L+1}(t)=p^{j}_{L+1}(t)\ ,\ j\neq i_{s}\Bigg\}\,.\end{split}$ (46)
where $S^{i_{s}}_{t}=\sum_{j\neq i_{s}}\left(p^{i_{s}}(t)\cdot x^{i_{s}}_{i_{s}j}(t)-p^{j}(t)\cdot x^{i_{s}}_{ji_{s}}(t)\right)+(1-\delta)S^{i_{s}}_{t-1}$ denotes the producer’s stock of commodities at time $t$; $\delta$ is the depreciation rate of the stock; $\mathscr{S}^{i_{s}}_{t}$ is a closed, bounded subset of $\mathbb{R}_{+}^{L}$ representing storage capacity constraints; and $\tau=(\tau_{1},\dots,\tau_{L})$ is a vector of time requirements for producing the outputs $y^{i_{s}}(t+\tau)$, with time being treated as a production factor. The time gap between production and exchange results in an intertemporal transfer [^3] A similar idea was introduced by Keynes under the notion of effective demand in his book [Key36].. The production-possibility set $\mathscr{Y}^{i_{s}}_{t}$ of producer $i_{s}$} satisfies, at each time $t$:
  1. 1.
    No free lunch: $\mathscr{Y}^{i_{s}}_{t}\cap{\mathbb{R}}_{+}^{L+1}\subseteq\{0\}$.
  2. 2.
    Free disposal: if $y^{i_{s}}\in\mathscr{Y}^{i_{s}}_{t}$ and $y^{\prime i_{s}}\leqslant y^{i_{s}}$ then $y^{\prime i_{s}}\in\mathscr{Y}^{i_{s}}_{t}$.
  3. 3.
    Possibility of inaction: $0\in\mathscr{Y}^{i_{s}}_{t}$.
  4. 4.
    Convexity: $\mathscr{Y}^{i_{s}}_{t}$ is a convex set of ${\mathbb{R}}^{L+1}$.
  5. 5.
    Closedness: $\mathscr{Y}^{i_{s}}_{t}$ is a closed set of ${\mathbb{R}}^{L+1}$.
Remark 11.1.
Additional assumptions could be made about $\mathscr{Y}^{i_{s}}_{t}$, such as network effects on competition and technology. One could assume that $\mathscr{Y}^{i_{s}}_{t}$ varies according to the position of the producer in the network.
The consumer’s new inter-temporal utility maximization problem is
$\max_{\displaystyle(p^{i_{d}},r^{i_{d}},X^{i_{d}})\in\mathcal{T}^{i_{d}}(p^{\neq{i_{d}}},r^{\neq{i_{d}}},X^{\neq{i_{d}}},D^{\neq{i_{d}}})\cap\mathfrak{T}^{i_{d}}\cap\mathfrak{D}^{i_{d}}}\ \sum_{t=t_{0}}^{T}\beta^{i_{d}}(t)\,u^{i_{d}}_{t}(x^{i_{d}},m^{i_{d}}).$ (47)
where $u^{i_{d}}_{t}$ denotes the intertemporal utility function at time $t$. This function retains its previously specified properties and is assumed to depend on the real cash balance $\frac{m^{i_{d}}(t)}{\mathbb{P}(p(t))}$, with $\mathbb{P}(p(t))$ serving as a price index. Moreover, it captures consumer preferences regarding future consumption. The intertemporal budget set is given by
$\begin{split}\mathcal{T}^{i_{d}}(p^{\neq{i_{d}}},r^{\neq{i_{d}}},X^{\neq{i_{d}}},D^{\neq{i_{d}}})=\Bigg\{&(p^{i_{d}},r^{i_{d}},X^{i_{d}},D^{i_{d}})\in\left(P\times[0,1]\times({\mathbb{R}}_{+}^{L})^{2(n_{d}+n_{s})}\times({\mathbb{R}}_{+})^{2{n_{f}}}\right)^{T+1}\ \Bigg|\ d^{i_{d}}(T)\leqslant 0\ ,\\ &x^{i_{d}}(t)\in\mathscr{X}^{i_{d}}_{t}\ ,\ m^{i_{d}}(t)\in\mathscr{M}^{i_{d}}_{t}\ ,\ x_{{i_{d}}j}^{i_{d}}(t)\leqslant x_{{i_{d}}j}^{j}(t)\ ,\ x_{j{i_{d}}}^{i_{d}}(t)\leqslant x_{j{i_{d}}}^{j}(t)\ ,\\ &d_{{i_{d}}j}^{i_{d}}(t)\leqslant d_{{i_{d}}j}^{j}(t)\ ,\ d_{j{i_{d}}}^{i_{d}}(t)\leqslant d_{j{i_{d}}}^{j}(t)\ ,\ p^{i_{d}}_{L+1}(t)=p^{j}_{L+1}(t)\ ,\ j\neq i_{d}\Bigg\}\,.\end{split}$ (48)
Considering the actions of the central bank $\mathfrak{C}$ as exogenous, the sets $\mathcal{F}$, $\mathcal{P}$, and $\mathcal{T}$ possess the necessary properties for the application of Berge’s Maximum Theorem 14.1. Therefore, the generalized game Proposition 2.2 applies, ensuring the existence of a transaction equilibrium in the economy.
11.2 Loanable funds market and interest rate
The loanable funds market operates as follows: at time $t_{0}$, agents offer and demand loanable funds to be repaid at time $t_{1}$. Given the equilibrium price matrix $p(t_{0})$ at time $t_{0}$, the future value of one monetary unit for the supplier $i$ is described by the following differential equation:
$\begin{cases}\displaystyle\frac{d{v^{i}}}{dt}(s)&=r^{i}(s)v^{i}(s),\\ v^{i}(t_{0})&=p_{L+1}(t_{0})\,.\end{cases}$ (49)
The solution is known to be
$v^{i}(t_{1})$ $=p_{L+1}(t_{0})\exp\left(\int_{t_{0}}^{t_{1}}r^{i}(s)\,ds\right).$ (50)
The interest paid at period $t_{1}$ is:
$v^{i}(t_{1})-v^{i}(t_{0})$ $=p_{L+1}(t_{0})\left(\exp\left(\int_{t_{0}}^{t_{1}}r^{i}(s)\,ds\right)-1\right)$
$=p_{L+1}(t_{0})\sum_{k=1}^{\infty}\frac{\left(\int_{t_{0}}^{t_{1}}r^{i}(s)\,ds\right)^{k}}{k!}\,.$
The interest earned by the agent can be decomposed into two components:
  1. 1.
    The present price of the good, $p^{i}_{k}(t_{0})$.
  2. 2.
    An exponential factor that captures the cumulative effect of the price of time, $r^{i}$. This factor reflects the initial amount plus the cost over the period $t_{0},t_{1}]$, along with the additional cost incurred from waiting until $t_{1}$ to receive the extra cost, and so on. The recursive accumulation of these costs leads to exponential growth.
Taking into account the variation in the value of money, and normalizing its price to one, we write
$v^{i}(t_{1})=\exp\left(\int_{t_{0}}^{t_{1}}r^{i}(s)\,ds\right).$ (51)
Thus, the lender exchanges an actual money flow $v^{i}(t_{0})=1$ for the borrower’s future money flow $v^{i}(t_{1})=\exp\left(\int_{t_{0}}^{t_{1}}r^{i}(s)\,ds\right)$, where the interest rate represents the price of time in this transaction.
A transaction in this market introduces an intertemporal distortion. Consider a two-period economy at times $t_{0}$ and $t_{1}$. First, in an intertemporal equilibrium of an exchange economy without a loanable funds market, the equilibrium values $(p,X)$ result in allocations $(x,m)$. Now, suppose we introduce a loanable funds market, enabling two consumers, $i$ and $j$, to engage in exchange. Specifically, agent $j$ incurs debt $v^{i}(t_{0})$ from agent $i$ at an interest rate $r^{i}=r^{i}(t_{1})$. The exchange will only be viable if it benefits both parties. What happens next?
  1. 1.
    At time $t_{0}$, agent $j$ receives additional money as a temporary wealth transfer, which alters the distribution of wealth. This results in an increase (or decrease) in the utility of agent $j$ (or agent $i$), and affects prices with transactions based on their network positions, wealth, and preferences.
  2. 2.
    At time $t_{1}$, agent $i$ receives money plus interest, reversing the wealth transfer. This increases (or decreases) the utility of agent $i$ (or agent $j$), and further influences prices.
This illustrates how the loanable funds market induces intertemporal interactions across markets. Compared to a situation without a loanable funds market (a strict intertemporal equilibrium where $d^{i}(t)\leqslant 0$ for all $t$), we can state:
Proposition 11.1.
Optimality – Fourth Comparison
The strict intertemporal maximization problem is suboptimal when compared to the intertemporal maximization problem with a loanable funds market.
A question remains: why would agents $i$ and $j$ agree to an intertemporal wealth transfer in an exchange economy? One explanation lies in the discount factor $\beta$, which represents price of time, and time preferences. A difference between $\beta^{i}$ and $\beta^{j}$, or shifts in future preferences in $u^{i}_{t}$ and $u^{j}_{t}$, could justify such transactions. Otherwise, with identical $\beta$ and $u$, agents would be indifferent between present and future goods in a linear utility framework. Differences in wealth distribution at $t_{1}$ could also justify the transfer.
When production is considered, a producer’s interest in wealth transfer depends on their transformation process. Since production requires time, borrowing wealth becomes necessary, serving as the primary motivation for a time market.
A producer requires goods for production and also time, as production is not instantaneous. Time thus becomes a production factor. Consider a one-shot production process starting at $t_{0}$ with duration $\tau$. Assuming a continuous framework, the producer’s profit maximization problem at time $t=t_{0}+\tau$ is given by
$\max\left(\sum_{j\neq{i_{s}}}p^{i_{s}}(t)\cdot x_{{i_{s}}j}^{i_{s}}(t)-\sum_{j\neq{i_{s}}}\int_{t_{0}}^{t}p^{j}(s)e^{\int_{s}^{t}r^{j}(v)\,dv}\cdot x_{j{i_{s}}}^{i_{s}}(s)\,ds\right).$ (52)
The product price can thus be decomposed into input prices $x_{j{i_{s}}}^{i_{s}}(t)$ and the price (rent) of time.
11.3 Quantity of money equation
Reviewing the debt equation at equilibrium (after normalization by the price of money), we begin by summing the debts:
$0$ $=\sum_{i=1}^{n}d^{i}(t)=d^{c}(t)+\sum_{{i_{f}}=1}^{n_{f}}d^{i_{f}}(t)+\sum_{{i_{s}}=1}^{n_{s}}d^{i_{s}}(t)+\sum_{{i_{d}}=1}^{n_{d}}d^{i_{d}}(t)$
$=\sum_{l=1}^{n_{f}}\left[d_{lc}(t)-d_{cl}(t)\right]+\sum_{{i_{f}}=1}^{n_{f}}\sum_{l=1}^{n}\left[d_{l{i_{f}}}(t)-d_{{i_{f}}l}(t)\right]+\sum_{{i_{s}}=1}^{n_{s}}\sum_{l=1}^{n_{f}}\left[d_{l{i_{s}}}(t)-d_{{i_{s}}l}(t)\right]+\sum_{{i_{d}}=1}^{n_{d}}\sum_{l=1}^{n_{f}}\left[d_{l{i_{d}}}(t)-d_{{i_{d}}l}(t)\right]$
$=\sum_{l=1}^{n_{f}}\left[d_{lc}(t)-d_{cl}(t)\right]+\sum_{{i_{f}}=1}^{n_{f}}\sum_{l=1}^{n}\left[d_{l{i_{f}}}(t)-d_{{i_{f}}l}(t)\right]$
$+\sum_{{i_{s}}=1}^{n_{s}}\left(\sum_{j\neq{i_{s}}}p^{j}(t)\cdot x_{j{i_{s}}}(t)-\sum_{j\neq{i_{s}}}p^{i_{s}}(t)\cdot x_{{i_{s}}j}(t)\right)+\sum_{{i_{s}}=1}^{n_{s}}\sum_{j=1}^{n_{f}}\left[(1+r^{j}(t))d_{j{i_{s}}}(t-1)-(1+r^{i_{s}}(t))d_{{i_{s}}j}(t-1)\right]$
$+\sum_{{i_{d}}=1}^{n_{d}}\left(\sum_{j\neq{i_{d}}}p^{j}(t)\cdot x_{j{i_{d}}}(t)-\sum_{j\neq{i_{d}}}p^{i_{d}}(t)\cdot x_{{i_{d}}j}(t)-\sum_{j=1}^{n_{s}}\pi_{j{i_{d}},t}-\sum_{l=1}^{n_{f}}\pi_{l{i_{d}},t}\right)$
$+\sum_{{i_{d}}=1}^{n_{d}}\sum_{j=1}^{n_{f}}\left[(1+r^{j}(t))d_{j{i_{d}}}(t-1)-(1+r^{i_{d}}(t))d_{{i_{d}}j}(t-1)\right]\,.$
Designate by $m_{ij}$ (resp. $m_{ji}$) the monetary flow from $i$ to $j$ (resp. from $j$ to $i$) resulting from trade and profit transfer, we can rewrite
$\sum_{i=1}^{n_{d}+n_{s}}\sum_{j\neq i}p^{j}(t)\cdot x_{ji}(t)+\sum_{i=1}^{n_{d}+n_{s}}\sum_{j=1}^{n_{f}}\left[(1+r^{j}(t))d_{ji}(t-1)-(1+r^{i}(t))d_{ij}(t-1)\right]$
$=$ $\sum_{l=1}^{n_{f}}\left[d_{cl}(t)-d_{lc}(t)\right]+\sum_{{i_{f}}=1}^{n_{f}}\sum_{l=1}^{n}\left[d_{{i_{f}}l}(t)-d_{l{i_{f}}}(t)\right]+\sum_{i=1}^{n_{d}+n_{s}}\sum_{j\neq i}m_{ji}(t)\,.$
We can rearrange the terms to get the quantity of money equation
$p(t)\cdot X(t)\cdot\mathbf{1}+(1+r(t))\cdot D^{-}(t-1)\cdot\mathbf{1}=(1+r(t))\cdot D^{+}(t-1)\cdot\mathbf{1}+\overline{M}(t)\left(v_{m}(t)+v_{d}(t)\right)\,$, (53)
where $D^{-}(t-1)$ represents the non-financial agents’ debts, $D^{+}(t-1)$ represents the non-financial agents’ receivables, $v_{m}(t)=\frac{\sum_{i=1}^{n_{d}+n_{s}}\sum_{j\neq i}m_{ji}(t)}{\overline{M}(t)}$ is the money velocity, $v_{d}(t)=\frac{-\left(d^{c}(t)+\sum_{i_{f}=1}^{n_{f}}d^{i_{f}}(t)\right)}{\overline{M}(t)}$ is the debt velocity, and $\overline{M}(t)=\underline{M}(t)-\left(d^{c}(t)+\sum_{i_{f}=1}^{n_{f}}d^{i_{f}}(t)\right)=\underline{M}(t)+\underline{D}(t)$, where $\underline{D}(t)$ represents the credits to the economy. One immediately observes that the equation now involves two additional factors: the debts and receivables from the previous period, along with their interest rates, and the new net debts. The analysis in Section 10.2 could lead to additional effects in the case of an expansionary monetary policy:
A perfectly expected (no uncertainty) increase in the money endowment $\underline{M}(t)$ or the credits to the economy $\underline{D}(t)$ could be equivalent to:
  1. 1.
    A negative effect on each other.
  2. 2.
    A negative effect on their velocities (hoarding).
  3. 3.
    An increase in prices (inflation).
  4. 4.
    An increase in quantities (production).
  5. 5.
    An increase in last period debts (as future money becomes less valuable), accompanied by an increase in interest rates (resulting from increased demand for last period debts).
  6. 6.
    A reduction in last period receivables (as future money becomes less valuable and lending becomes less attractive), accompanied by a decrease in the interest rates on last period receivables (due to reduced demand for last period debts).
11.4 Monetary policy and dynamics
We now examine monetary policy within an intertemporal equilibrium under the certainty framework. The central authority influences the economy through two primary channels of monetary policy transmission:
  1. 1.
    The money endowment: $\underline{M}(t)=\sum_{i=1}^{n}\underline{m}^{i}(t)$,
  2. 2.
    The credits to the economy: $\underline{D}(t)=-\left(d^{c}(t)+\sum_{i_{f}=1}^{n_{f}}d^{i_{f}}(t)\right)$.
To understand the impact of these mechanisms, consider an exchange economy across two periods, $t_{0}$ and $t_{1}$:
  1. 1.
    Money Supply: Suppose the central authority $\mathfrak{C}$ increases the money endowment $\underline{m}^{i}(t_{0})$ by $\delta_{m}$. This intervention has the following effects:
    1. (a)
      It decreases the value of money at both $t_{0}$ and $t_{1}$ in equilibrium, increases the wealth of agent $i$ by $\delta_{m}$, and alters the system of prices. This also reduces the relative value of the wealth of agents $j\neq i$, leading to redistributive effects and changes in purchasing power.
    2. (b)
      In the case of hoarding, there is no immediate change at $t_{0}$ beyond the increase in agent $i$’s wealth by $\delta_{m}$. The effects on prices and redistribution are instead deferred to period $t_{1}$.
    3. (c)
      It reduces future net debts of agent $i$ and decreases the value of time by lowering his incentive to borrow.
    4. (d)
      The updated money supply after the intervention is given by
      $\underline{M}^{new}(t_{0})$ $=\underline{M}(t_{0})+\delta_{m},$
      $\underline{M}^{new}(t_{1})$ $=\underline{M}(t_{1})+\delta_{m}.$
  2. 2.
    Credit Supply: Suppose a financial institution $i_{f}$ increases the credit $d^{i_{f}}_{{i_{f}}i}(t_{0})$ extended to agent $i$. This generates the following consequences:
    1. (a)
      At $t_{0}$, it decreases the value of money at equilibrium, increases the wealth of agent $i$ by $d^{i_{f}}_{{i_{f}}i}(t_{0})$, and adjusts the price system. This diminishes the relative wealth of other agents $j\neq i$. At $t_{1}$, agent $i$’s wealth is reduced by the repayment $d^{i_{f}}_{{i_{f}}i}(t_{0})(1+r^{i_{f}})$, resulting in intertemporal redistribution.
    2. (b)
      It decreases the value of time by increasing the availability of credit and reducing the cost of borrowing.
    3. (c)
      At $t_{1}$, the destruction of $d^{i_{f}}_{{i_{f}}i}(t_{0})$ restores the value of money, but the effects on distribution and purchasing power persist.
    4. (d)
      The updated money supply after the intervention becomes
      $\underline{M}^{new}(t_{0})$ $=\underline{M}(t_{0})+d^{i_{f}}_{{i_{f}}i},$
      $\underline{M}^{new}(t_{1})$ $=\underline{M}(t_{1})\,.$
In conclusion, both channels of monetary policy—money endowments and credit supply—affect the value of money, wealth distribution, and intertemporal dynamics. An increase in the money endowment tends to devalue money across periods and reduce future borrowing costs. In contrast, increased credit supply also reduces the value of money in the short run but reverses its impact as debts are repaid, generating complex intertemporal redistributive effects. These mechanisms highlight the complex interplay between monetary policy and economic equilibrium over time, with the system’s response depending on its specific configuration—such as the structure of utilities, production technologies, and the design of monetary policy.
12 Uncertainty
Under uncertainty, agents are assumed to form expectations about future events to guide both their current and future decisions. In contrast to the rational expectations hypothesis ([Mut61]), which is grounded in conditional expectations, we argue that expectations should instead be formed based on the mode. The rationale is that individuals are more inclined to act upon the most likely outcome rather than an average. Relying on the mean of two distinct possibilities may result in decisions that are suboptimal for both, whereas empirical evidence suggests that individuals typically focus on the most probable event as the one most worth preparing for. While situations of indecision may arise—where multiple outcomes appear equally plausible—individuals often resolve this ambiguity by relying on personal beliefs, social norms, cognitive simplification, or external influences. In practice, these mechanisms contribute to the emergence of a unimodal distribution.
A second essential feature of our framework is the subjectivity of expectations. We model expectations as conditional probabilities based on the specific information available to each agent at a given moment. This captures the idea that agents have only partial knowledge of the world, and hence their beliefs are represented through subjective probability measures.
Lastly, our formulation accounts for the bounded memory of agents, acknowledging their tendency to place greater weight on recent information when forming expectations.
12.1 The mode as an anticipation criterion
Given $x\in{\mathbb{R}}^{d}$, a random variable $y$ with general law $\mathbb{P}$, and a measurable function $f:\mathbb{R}^{d}\times{\mathbb{R}}\to{\mathbb{R}}$, the mode of $f$ is defined as follows
$\mathbb{M}(f(x,y))=\arg\sup_{z\in\mathcal{Z}}\limsup_{r\to 0}\frac{\mathbb{P}_{x}(B_{r}(z))}{\lambda(B_{r}(z))}\,$, (54)
where $\mathbb{P}_{x}$ is the push-forward probability measure defined by $\mathbb{P}_{x}(D)=\mathbb{P}\{y\,\mid\,u(x,y)\in D\}$, $\lambda$ is the Lebesgue measure, and $B_{r}(z)$ the unit ball of radius $r$ and center $z$.
When $y$ has density $p(y)$, then $\mathbb{P}_{x}=p\circ f(x,\cdot)^{-1}$ and one could reformulate this problem as follows
$\mathbb{M}(f(x,y))=\arg\max_{z\in F(x)}g(z,x)\,$, (55)
where $F(x)=\{z\,\mid\,\exists y\ \text{s.t.}\ u(x,y)=z\}$ is the image of $f(x,\cdot)$ and $g(z,x)=\mathbb{P}_{x}(z)$. One could suppose that
  1. 1.
    $f$ is jointly continuous in $x$ and $y$.
  2. 2.
    for each $x$, the map $y\to f(x,y)$ is a $C^{1}$-diffeomorphism onto $\mathbb{R}$}.
  3. 3.
    $p(y)$ is continuous on $\mathbb{R}$}.
So, $g(z,x)=p(f^{-1}(x,z))\left|\frac{df^{-1}}{dz}\right|$ is continuous in both $x$ and $z$, and
  1. 1.
    $\|f^{-1}(x,z)\|\to\infty$ as $\|z\|\to\infty$.
  2. 2.
    $\exists C>0$ such that $|\det(D_{z}f^{-1}(x,z))|<C$ for all $z$.
  3. 3.
    $p(y)\to 0$ as $\|y\|\to\infty$.
So $g(z,x)\to 0$ as $\|z\|\to\infty$. One could then apply the generalized maximum Theorem 14.4 (as proved in [FKV14] - theorem 1.2), to prove that the mode is an an upper semicontinuous map $\mathbb{M}(x)$, since $F$ is a constant set-valued map. Moreover, under the strict quasiconcavity of $g(x,\cdot)$, one could end up with a continuous function.
When $y$ has a discrete infinite values, $f:\mathbb{R}^{d}\times E\to{\mathbb{R}}$ is supposed to be continuous in $x$, the feasible set $F(x)=\{f(x,y_{i})\ :\ i\in{\mathbb{N}}\}$ is countable, and $g(z,x)=\displaystyle\sum_{i=0}^{\infty}\mathbb{P}(y=y_{i})\,\mathbf{1}_{\{f(x,y_{i})=z\}}$.
One could check first that $F(x)$ is non empty, and that $F$ is lower semicontinuous: Let $x_{0}\in\mathbb{R}^{d}$ be arbitrary, and let $U\subset{\mathbb{R}}$ be an open set such that $F(x_{0})\cap U\neq\emptyset$. Then there exists some index $i_{0}\in{\mathbb{N}}$ such that $f(x_{0},y_{i_{0}})\in U$. Since $x\mapsto f(x,y_{i_{0}})$ is continuous, and $U$ is open, there exists a neighborhood $V(x_{0})$ of $x_{0}$ such that for all $x\in V$, $f(x,y_{i_{0}})\in U$. But then, for all $x\in V$, we have $F(x)\cap U\neq\emptyset$, because $f(x,y_{i_{0}})\in F(x)\cap U$. Hence, by definition, $F$ is lower semicontinuous at $x_{0}$. Since $x_{0}$ was arbitrary, $F$ is lower semicontinuous on ${\mathbb{R}}^{d}$.
Second, $g$ is upper semicontinuous: Fix $z\in{\mathbb{R}}$, and let $x_{0}\in{\mathbb{R}}^{d}$. Let $\varepsilon>0$ be arbitrary. Define the index set
$I_{z}:=\{i\in\mathbb{N}\mid f(x_{0},y_{i})=z\}.$
Note that $g(z,x_{0})=\sum_{i\in I_{z}}p_{i}$. Let $I_{z}^{\delta}\subset I_{z}$ be a finite subset such that
$\sum_{i\in I_{z}\setminus I_{z}^{\delta}}p_{i}<\frac{\varepsilon}{2}.$
Since each $f(x,y_{i})$ is continuous, and $f(x_{0},y_{i})=z$ for all $i\in I_{z}^{\delta}$, there exists a neighborhood $V_{i}(x_{0})$ such that for all $x\in V_{i}$,
$z-\delta<f(x,y_{i})<z+\delta,$
and hence $f(x,y_{i})\neq z$ for small $\delta>0$, unless $f(x,y_{i})\equiv z$. Let $V:=\bigcap_{i\in I_{z}^{\delta}}V_{i}$. Then for all $x\in V$, we have
$g(z,x)\leqslant\sum_{i\in I_{z}^{\delta}}p_{i}+\sum_{i\in I_{z}\setminus I_{z}^{\delta}}p_{i}<g(z,x_{0})+\varepsilon.$
Thus, $g(z,x)\leqslant g(z,x_{0})+\varepsilon$ for all $x\in V$, which shows upper semicontinuity at $x_{0}$. Since $x_{0}$ was arbitrary, $g(z,x)$ is upper semicontinuous on ${\mathbb{R}}^{d}$.
In order to apply Theorem 14.4, one needs to additionally show that, for every compact set $K\subset{\mathbb{R}}^{d}$, the graph
$\operatorname{Gr}_{K}(F):=\{(x,z)\in K\times\mathbb{R}\mid z\in F(x)\}$ (56)
is closed.
Indeed, let $K\subset{\mathbb{R}}^{d}$ be a compact set. Consider a sequence $(x_{n},z_{n})\in\operatorname{Gr}_{K}(F)$ such that
$(x_{n},z_{n})\to(x,z)\in K\times{\mathbb{R}}\,.$
Since $z_{n}\in F(x_{n})$, there exists $i_{n}\in{\mathbb{N}}$ such that $z_{n}=f(x_{n},y_{i_{n}})$. Because $\mathbb{N}$} is countable, we may extract a subsequence (still denoted $i_{n}$) such that $i_{n}=i^{\star}$ is constant along this subsequence. Then we have $z_{n}=f(x_{n},y_{i^{\star}})\to z$. By continuity of $f(x,y_{i^{\star}})$ in $x$, we conclude that $f(x_{n},y_{i^{\star}})\to f(x,y_{i^{\star}})$, so $z=f(x,y_{i^{\star}})\in F(x)$. Thus, $(x,z)\in\operatorname{Gr}_{K}(F)$, and the graph is closed.
Since lower semicontinuity and closedness of the graph over compacts are satisfied, $F$ is $K$-inf-compact.
Suppose additionally that:
  1. 1.
    $\sum_{i=0}^{\infty}\mathbb{P}(y=y_{i})=1$.
One can deduce that the mode is an upper semicontinuous correspondence. Moreover, if one can guarantee the uniqueness of the mode of $\mathbb{P}_{x}$ — for example, by assuming that:
  1. 1.
    $f(x,\cdot)$ is injective for each $x$.
  2. 2.
    The distribution of $y$ is unimodal with mode $y_{i^{\star}}$.
Then $\mathbb{M}(f(x,y))=f(x,y_{i^{\star}})$ is a continuous function of $x$.
12.2 Subjective probabilities, memory and expectations
  1. 1.
    Let $(\psi_{s})_{s\in\mathbb{N}}$ be a discrete-time stochastic process taking values in a measurable space $\Psi$, representing the state of the world, which is not directly observable by the agents.
  2. 2.
    Let $\varpi\in\mathbb{N}^{\star}$ be the agents’ memory window (limited memory).
  3. 3.
    Define the total information at time $t$ over the horizon $t-\varpi,t\,]$ by
    $I_{t}\;=\;\sigma\Bigl(Z_{s}^{\psi_{s}},\;A_{s},\;\mathcal{I}_{s}^{\psi_{s}}:\;s\in[t-\varpi,\,t]\Bigr)\,$, (57)
    where
    • $Z_{s}^{\psi_{s}}$ is the state–process of the economy,
    • $A_{s}$ is the history of past aggregate actions,
    • $\mathcal{I}_{s}^{\psi_{s}}=(\mathcal{I}_{s}^{1},\dots,\mathcal{I}_{s}^{n})$ collects all public and private information signals (announcements, rumours, etc.),
    • each of these processes is a measurable function of the underlying state of the world $\psi_{s}$.
  4. 4.
[…]
    $\sigma\bigl(A_{s},\;\mathcal{I}_{s}^{i}:s\in[\,t-\varpi,\,t]\bigr)\;\subseteq\;I^{i}_{t}\,.$
  5. 5.
    Each agent $i$ holds a subjective probability measure $\mathbb{P}^{i}$ on the measurable space $\Psi$, in the spirit of Keynes’s notion of subjective probability [^4] J. M. Keynes, A Treatise on Probability, 1921.. These probabilities are updated via conditioning on the agent’s personal information $\sigma$–algebra $I^{i}_{t}$, and need not agree with any objective law.
  6. 6.
    Each function $f\,:\,\psi\to f(\psi):=f^{\psi}$ epresenting an aspect of the state of the world has finite preimages: $f^{-1}(\psi)$ is finite. This ensures that the agent’s conditional belief $\mathbb{P}^{i}\!\bigl(\cdot\mid I^{i}_{t}\bigr)$ has finite support over the state space $\Psi$, reflecting the agent’s limited set of conceivable world states.
  7. 7.
    For any $I^{i}_{t}$–measurable function $y=y^{\psi}:=y(\psi)$, define agent $i$’s anticipation of $y$ at time $t$ by
    $\mathbb{M}\bigl(y\mid I^{i}_{t}\bigr):=\underset{y^{\prime}}{\arg\max}\;\mathbb{P}^{i}\!\bigl(y=y^{\prime}\mid I^{i}_{t}\bigr)\,;$ (58)
    i.e., the conditional mode of $y$ given the limited‐memory information $I^{i}_{t}$ under the subjective law $\mathbb{P}^{i}$.
12.3 Decision under uncertainty
Given the random structure of the economy, each agent seeks to maximize an expected objective function, based on their beliefs. The decision process is intertemporal: agents choose deterministic strategies that depend on their anticipations of future outcomes. While each agent’s own strategy is determined through optimization, the future strategies of other agents are not directly observed and are instead replaced by subjective anticipations.
Let agent $i$ be given. We assume that their current strategy $a^{i,\psi_{t_{0}}}(t_{0}\mid t_{0})$ is known almost surely at time $t_{0}$, and is thus observable by all other agents. In contrast, future strategies are unknown ex ante, as they depend on the evolution of the state. These are denoted by $a^{j,\psi_{t}}(t\mid t_{0})$ for $t_{0}+1\leqslant t\leqslant T$, reflecting agent $j$’s planned action at time $t$ conditional on the initial time $t_{0}$ and future state $\psi_{t}$.
At each time $t$, agents make decisions based on the information available to them, incorporating expectations over future states and actions. These strategies are updated dynamically over time, as new information becomes available. Accordingly, we denote by $a^{i}(t+s\mid t)$ the deterministic action of agent $i$ at time $t+s$, as decided at time $t$, and by $a^{j,\psi_{t+s}}(t+s\mid t)$ the corresponding anticipated actions of other agents conditional on the future state $\psi_{t+s}$.
Let $\mathbb{M}_{i,t}(y):=\mathbb{M}(y\mid I^{i}_{t})$ denote the conditional mode, that is, the subjective anticipation of agent $i$ at time $t$ given their information set $I^{i}_{t}$.
12.4 Random economy setting
The stochastic dynamic economy is defined by the tuple: $\mathcal{E}=\left(w^{\psi_{t}}_{t},u^{\psi_{t}}_{t},\mathscr{X}^{\psi_{t}}_{t},\mathscr{M}^{\psi_{t}}_{t},\mathscr{Y}^{\psi_{t}}_{t},C^{\psi_{t}}_{t},F^{\psi_{t}}_{t},\Psi\right)$, with the same notation previously introduced in Section 11. As usual, we define $n=n_{d}+n_{s}+n_{f}+1$ being the total number of agents, where $n_{d}$ represents the number of consumers, $n_{s}$ the number of suppliers, $n_{f}$ the number of financial institutions, and $\mathfrak{C}$ denotes the central bank.
Let $W^{i,\psi_{t}}=\left(w^{i,\psi_{t_{0}}}(t_{0}),\underline{m}^{i,\psi_{t_{0}}}(t_{0}),\dots,w^{i,\psi_{T}}(T),\underline{m}^{i,\psi_{T}}(T)\right)$ represent the holdings of commodities by agent $i$ at the beginning of each period. The utility of consumer $i$ at time $t$ is denoted by $u^{i,\psi_{t}}_{t}$, which is assumed to be continuous and increasing.
At each time $t$, let $p^{i}=(p^{i}(t_{0}),\dots,p^{i}(T))$ represent the price vectors for consumer $i$ across all periods, and define $X^{i}(t)=(x^{i}_{ji}(t),x^{i}_{ij}(t))$ as the matrix of potential goods transactions for agent $i$. The corresponding exchange capacities are given by: $C^{i,\psi_{t}}_{t}=(c^{i,\psi_{t}}_{ji}(t),c^{i,\psi_{t}}_{ij}(t))\in(\mathbb{R}_{+}^{L})^{2(n_{d}+n_{s})}$, for $1\leqslant j\leqslant n_{d}+n_{s}$, $j\neq i$, and $t_{0}\leqslant t\leqslant T$.
Recall the goods allocation $x^{i}(t)$ and the money holdings $m^{i}(t)$. The feasible set of transactions is characterized by the topological constraint: $\mathfrak{T}^{i}=\left\{X^{i}\in(\mathbb{R}_{+}^{L})^{2(n_{d}+n_{s})}\mid X^{i}(t)\leqslant\mathbb{M}_{i,t_{0}}\left(C^{i,\psi_{t}}_{t}\right)\right\}$.
Similarly, the financial transactions matrix for agent $i$ is defined as: $D^{i}(t)=(d^{i}_{ji}(t),d^{i}_{ij}(t))$, with financial capacities: $F^{i,\psi_{t}}_{t}=(f^{i,\psi_{t}}_{ji}(t),f^{i,\psi_{t}}_{ij}(t))\in(\mathbb{R}_{+}^{L})^{2n_{f}}$, and the corresponding financial constraint: $\mathfrak{D}^{i}=\left\{D^{i}\in(\mathbb{R}_{+}^{L})^{2n_{f}}\mid D^{i}(t)\leqslant\mathbb{M}_{i,t_{0}}\left(F^{i,\psi_{t}}_{t}\right)\right\}$, for $1\leqslant j\leqslant n_{f}$, $j\neq i$, and $t_{0}\leqslant t\leqslant T$.
We allow both interbank flows: $(d^{i}_{ji}(t),d^{i}_{ij}(t),d^{j}_{ji}(t),d^{j}_{ij}(t))_{t_{0}\leqslant t\leqslant T,\ 1\leqslant i,j\leqslant n_{f},\ i\neq j}$, and flows between financial institutions and the central bank $\mathfrak{C}$: $(d^{i}_{ci}(t),d^{i}_{ic}(t),d^{c}_{ci}(t),d^{c}_{ic}(t))_{t_{0}\leqslant t\leqslant T,\ 1\leqslant i\leqslant n_{f}}$, where $g_{t}$ is a continuous function representing a money creation rule determined by the central bank at time $t$.
As before, each agent $j\neq i$ can incur debt, which, in the case of consumers and producers, is explicitly detailed as
[…]
where $r^{i}(t)$ represents the interest rate applied to agent $i$’s debt at time $t$.
12.5 Anticipation and non-uniqueness of equilibrium
In the sequel, the information set is given by $I_{t}\;=\;\sigma\Bigl(Z_{s}^{\psi_{s}},\;A_{s},\;\mathcal{I}_{s}^{\psi_{s}}:\;t-\varpi\leqslant s\leqslant t\Bigr)$ where $Z_{s}^{\psi_{s}}$ represents the stochastic process
$Z^{\psi_{t}}_{t}$ $=(w^{\psi_{t}}_{t},\underline{m}^{\psi_{t}}_{t},u^{\psi_{t}}_{t},\mathscr{X}^{\psi_{t}}_{t},\mathscr{M}^{\psi_{t}}_{t},\mathscr{Y}^{\psi_{t}}_{t},\mathscr{S}^{\psi_{t}}_{t},C^{\psi_{t}}_{t},F^{\psi_{t}}_{t})$ (60)
which characterizes the economy. Let $z^{i,\psi_{t}}_{t}\subset Z^{\psi_{t}}_{t}$ be a subset that characterizes agent $i$’s own optimization program. Specifically:
  • For financial institutions: $z^{i,\psi_{t}}_{t}=F^{\psi_{t}}_{t}$,
  • For producers: $z^{i,\psi_{t}}_{t}=(\mathscr{Y}^{\psi_{t}}_{t},\mathscr{S}^{\psi_{t}}_{t},C^{\psi_{t}}_{t},F^{\psi_{t}}_{t})$,
  • For consumers: $z^{i,\psi_{t}}_{t}=(w^{\psi_{t}}_{t},\underline{m}^{\psi_{t}}_{t},u^{\psi_{t}}_{t},\mathscr{X}^{\psi_{t}}_{t},\mathscr{M}^{\psi_{t}}_{t},C^{\psi_{t}}_{t},F^{\psi_{t}}_{t})$.
The subjective information satisfies $\sigma(z^{i,\psi_{s}}_{s}:t-\varpi\leqslant s\leqslant t)\subseteq I_{i,t}$, which implies that each agent is fully aware of their own actual and past characteristics relevant to their optimization problem. As a consequence, we have the following properties:
$\begin{cases}\mathbb{M}_{i,t}(a^{j,\psi_{t}}(t\mid t))&=a^{j,\psi_{t}}(t\mid t),\\ \mathbb{M}_{i,t}(z^{i,\psi_{t}}_{t})&=z^{i,\psi_{t}}_{t}.\end{cases}$ (61)
Given a state $\psi$, what is the anticipated value of the future action $a^{j,\psi}(t\mid s)$ from the perspective of another agent $i\neq j$ at time $s$? One could interpret $a^{j,\psi}(t\mid s)$ as an element of the Nash equilibrium set at state $\psi$ and time $t$, denoted by $\mathbf{E}^{\psi}_{t}$, then apply the anticipation operator $\mathbb{M}_{i,t}$.
12.6 Games and Nash equilibria
When analyzing trade as a game, several structural aspects should be taken into account. Nature selects the actual state of the world, denoted by $\psi_{t_{0}}$. At time $t_{0}$, each agent $i$ observes their own type $z^{i,\psi_{t_{0}}}_{t_{0}}$, reducing the game at this stage to a classical complete-information game, as previously analyzed.
However, agents remain uncertain about future states of the world and future types—including their own. Each agent forms inferences about these uncertain elements based on their private information $I^{i}_{t}$ and their subjective belief system $\mathbb{P}^{i}$, operationalized through the conditional mode $\mathbb{M}_{i,t}$.
Under this framework, the game can be viewed as involving $T\,\sum_{i=1}^{n}\lvert\Psi_{i}\rvert$ players, where $\Psi_{i}$ denotes the finite set of states of the world that agent $i$ assigns positive probability to and $T$ the number of periods. This formulation preserves the structural conditions required for the existence of Nash equilibria.
12.7 The trade equilibrium
Now we could establish the conditions of transactions. We will make the following additional assumptions to ensure the continuity and the uniqueness of the mode:
  1. a’
    $u^{i,\psi}(\cdot)$ is continuous, increasing and quasi-concave on $\mathscr{X}^{i,\psi}$ for all $\psi\in\Psi$.
  2. e
    The subjective probability $\mathbb{P}^{i}$ has finite support $\Psi^{i}$ for all $1\leqslant i\leqslant n$.
  3. f
    The maps $\psi\to\pi^{\psi}$ and $\psi\to u^{\psi}$ are injective, for each fixed context of evaluation.
  4. g
    The conditional distribution of $\psi$ to the subjective information is unimodal.
Let consider the financial institution’s optimization problem:
$\max_{\displaystyle(r^{i_{f}},D^{i_{f}})\in\mathcal{F}^{i_{f}}(r^{\neq{i_{f}}}_{t_{0}},D^{\neq{i_{f}}}_{t_{0}})\cap\mathfrak{D}^{i_{f}}}\ \sum_{t=t_{0}}^{T}\beta^{i_{f}}(t)\,\mathbb{M}_{i_{f},t_{0}}\left(\pi^{i_{f},\psi_{t}}_{t}(r^{i_{f}},D^{i_{f}})\right)\,$, (62)
where $\pi^{i_{f},\psi_{t}}_{t}(r^{i_{f}},D^{i_{f}})=r^{i_{f}}(t\mid t_{0})\sum_{j\neq{i_{f}}}d^{i_{f}}_{{i_{f}}j}(t-1\mid t_{0})-\sum_{j\neq{i_{f}}}r^{j,\psi_{t}}(t\mid t_{0})\,d^{i_{f}}_{j{i_{f}}}(t-1\mid t_{0})$, and $\beta^{i}(t)=\prod_{k=t_{0}}^{t}\frac{1}{1+r^{i}(k\mid t_{0})}$ is the actualization factor.
The financial institution’s constraint is then given by:
$\begin{split}\mathcal{F}^{i_{f}}(r^{\neq{i_{f}}}_{t_{0}},D^{\neq{i_{f}}}_{t_{0}})=\Bigg\{&(r^{i_{f}},D^{i_{f}})\in\left([0,1]\times(\mathbb{R}_{+}^{L})^{2n}\right)^{T+1}\ \mid\ d^{i_{f}}(T)\leqslant 0\ ,\ \sum_{j\neq{i_{f}}}d^{i_{f}}_{{i_{f}}j}(t\mid t_{0})\leqslant g_{t}\left({d^{i_{f}}(t\mid t_{0})}^{-}\right)\ ,\\ &d^{i_{f}}_{j{i_{f}}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{f},t_{0}}\left(d^{j,\psi_{t}}_{j{i_{f}}}(t\mid t_{0})\right)\ ,\ d^{i_{f}}_{{i_{f}}j}(t\mid t_{0})\leqslant\mathbb{M}_{i_{f},t_{0}}\left(d^{j,\psi_{t}}_{j{i_{f}}}(t\mid t_{0})\right)\Bigg\}\,.\end{split}$ (63)
The producer’s problem is
$\max_{\displaystyle(p^{i_{s}},r^{i_{s}},X^{i_{s}})\in\mathcal{P}^{i_{s}}(p^{\neq{i_{s}}}_{t_{0}},r^{\neq{i_{s}}}_{t_{0}},X^{\neq{i_{s}}}_{t_{0}},D^{\neq{i_{s}}}_{t_{0}})\cap\mathfrak{T}^{i_{s}}\cap\mathfrak{D}^{i_{s}}}\sum_{t=t_{0}}^{T}\beta^{i_{s}}(t)\,\mathbb{M}_{i_{s},t_{0}}\left(\pi^{i_{s},\psi_{t}}_{t}(p^{i_{s}},r^{i_{s}},X^{i_{s}})\right)\,$, (64)
where
$\pi^{i_{s},\psi_{t}}_{t}(p^{i_{s}},r^{i_{s}},X^{i_{s}})$ $=\left(\sum_{j\neq{i_{s}}}p^{i_{s}}(t\mid t_{0})\cdot x_{{i_{s}}j}^{i_{s}}(t\mid t_{0})-\sum_{j\neq{i_{s}}}p^{j,\psi_{t}}(t\mid t_{0})\cdot x_{j{i_{s}}}^{i_{s}}(t\mid t_{0})\right)$
$+\sum_{j=1}^{n_{f}}\left[(1+r^{i_{s}}(t\mid t_{0}))d^{i_{s}}_{{i_{s}}j}(t-1\mid t_{0})-(1+r^{j,\psi_{t}}(t\mid t_{0}))d^{i_{s}}_{j{i_{s}}}(t-1\mid t_{0})\right]\,$,
and $\mathfrak{D}^{i_{s}}=\left\{D^{i_{s}}\in(\mathbb{R}_{+}^{L})^{2n_{f}}\mid D^{i_{s}}(t)\leqslant\mathbb{M}_{i_{s},t_{0}}\left(F^{i_{s},\psi_{t}}_{t}\right)\right\}$.
The production constraint becomes
$\begin{split}\mathcal{P}^{i_{s}}(p^{\neq{i_{s}}}_{t_{0}},r^{\neq{i_{s}}}_{t_{0}},X^{\neq{i_{s}}}_{t_{0}},D^{\neq{i_{s}}}_{t_{0}})=\Bigg\{&(p^{i_{s}},r^{i_{s}},X^{i_{s}},y^{i_{s}})\in\left(P\times[0,1]\times(\mathbb{R}_{+}^{L})^{2(n_{d}+n_{s})}\times{\mathbb{R}}^{L}\right)^{T+1}\ \Bigg|\ d^{i_{s}}(T)\leqslant 0\ ,\\ &\left(y^{i_{s}}(t+\tau\mid t_{0}),-\tau\right)\in\mathbb{M}_{i_{s},t_{0}}(\mathscr{Y}^{i_{s},\psi_{t}})\ ,\ S^{i_{s}}_{t}\in\mathbb{M}_{i_{s},t_{0}}(\mathscr{S}^{i_{s},\psi_{t}})\ ,\\ &x_{{i_{s}}j}^{i_{s}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{s},t_{0}}(x_{{i_{s}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ x_{j{i_{s}}}^{i_{s}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{s},t_{0}}(x_{j{i_{s}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &d_{{i_{s}}j}^{i_{s}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{s},t_{0}}(d_{{i_{s}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ d_{j{i_{s}}}^{i_{s}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{s},t_{0}}(d_{j{i_{s}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &p^{i_{s}}_{L+1}(t\mid t_{0})=\mathbb{M}_{i_{s},t_{0}}(p^{j,\psi_{t}}_{L+1}(t\mid t_{0}))\ ,\ j\neq i_{s}\Bigg\}\,.\end{split}$ (65)
The consumer’s new inter-temporal utility maximization problem is
$\max_{\displaystyle(p^{i_{d}},r^{i_{d}},X^{i_{d}})\in\mathcal{T}^{i_{d}}(p^{\neq{i_{d}}}_{t_{0}},r^{\neq{i_{d}}}_{t_{0}},X^{\neq{i_{d}}}_{t_{0}},D^{\neq{i_{d}}}_{t_{0}})\cap\mathfrak{T}^{i_{d}}\cap\mathfrak{D}^{i_{d}}}\ \sum_{t=t_{0}}^{T}\beta^{i_{d}}(t)\,\mathbb{M}_{i_{d},t_{0}}\left(u^{i_{d},\psi_{t}}_{t}(x^{i_{d}},m^{i_{d}})\right)\,$, (66)
The intertemporal budget set is given by:
$\begin{split}\mathcal{T}^{i_{d}}(p^{\neq{i_{d}}}_{t_{0}},r^{\neq{i_{d}}}_{t_{0}},X^{\neq{i_{d}}}_{t_{0}},D^{\neq{i_{d}}}_{t_{0}})=\Bigg\{&(p^{i_{d}},r^{i_{d}},X^{i_{d}})\in\left(P\times[0,1]\times(\mathbb{R}_{+}^{L})^{2(n_{d}+n_{s})}\right)^{T+1}\ \Bigg|\ d^{i_{d}}(T)\leqslant 0\ ,\\ &m^{i_{d}}(t\mid t_{0})\in\mathbb{M}_{i_{d},t_{0}}(\mathscr{M}^{i_{d},\psi_{t}})\ ,\\ &\mathbb{M}_{i_{d},t_{0}}(w^{i_{d},\psi_{t}}_{t})+\sum_{j\neq{i_{d}}}x_{j{i_{d}}}^{i_{d}}(t\mid t_{0})-\sum_{j\neq{i_{d}}}x_{{i_{d}}j}^{i_{d}}(t\mid t_{0})\in\mathbb{M}_{i_{d},t_{0}}(\mathscr{X}^{i_{d},\psi_{t}})\ ,\\ &x_{{i_{d}}j}^{i_{d}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{d},t_{0}}(x_{{i_{d}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ x_{j{i_{d}}}^{i_{d}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{d},t_{0}}(x_{j{i_{d}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &d_{{i_{d}}j}^{i_{d}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{d},t_{0}}(d_{{i_{d}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ d_{j{i_{d}}}^{i_{d}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{d},t_{0}}(d_{j{i_{d}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &p^{i_{d}}_{L+1}(t\mid t_{0})=\mathbb{M}_{i_{d},t_{0}}(p^{j,\psi_{t}}_{L+1}(t\mid t_{0}))\ ,\ j\neq i_{d}\Bigg\}\,.\end{split}$ (67)
Considering once more the exogenous and deterministic nature of the action of $\mathfrak{C}$, and fixing the initial state of nature $\psi_{t_{0}}$, the sets $\mathcal{F}$, $\mathcal{P}$, and $\mathcal{T}$ exhibit the requisite properties for the application of Berge’s Maximum Theorem 14.1. Moreover, the generalized game Proposition 2.2 applies and guarantees the existence of a Nash equilibrium for the $L+2$ markets of the economy at time $t=t_{0}$.
It is important to highlight a new phenomenon of time inconsistency: following each transition from period $s$ to period $s+1$, an adjustment is applied to the decision variables $A$ based on the evolving state, updated information sets, and actions. This results in the new decision $a(t\mid s+1)$. The difference between the previous and new decisions, $a(t\mid s+1)-a(t\mid s)$, signifies this correction. While this adjustment ensures equilibrium in transactions, it may also lead to disequilibrium in the supply and demand balance (rationing). In such cases, suppliers might be compelled to reduce their expected sales, while buyers could face constraints on their anticipated purchases. One should consider, for instance, scenarios where producers have already committed to production plans that limit flexibility.
A reexamination of the quantity of money equation deserves careful consideration, especially in light of the evolving nature of expectations and information flows. The equation is given by:
$p(t)\cdot X(t)\cdot\mathbf{1}+(1+r(t))\cdot D^{-}(t-1)\cdot\mathbf{1}$ $=(1+r(t))\cdot D^{+}(t-1)\cdot\mathbf{1}+\overline{M}(t)\left(v_{m}(t)+v_{d}(t)\right),$
It is crucial to observe that, with the changing conception of the future, all the variables are inherently tied to the expectations. These expectations are shaped by the available information at that moment, and as such, they reflect a forward-looking view that evolves with new data and insights. This evolving nature of expectations necessitates a rethinking of how we approach the dynamics of supply, demand, and the interplay of monetary policies over time.
13 Open economy
Now, we introduce foreign exchange. We consider $H$ distinct economies, each represented as $\mathcal{E}^{h}=\left(w^{\psi_{t}}_{t},u^{\psi_{t}}_{t},\mathscr{X}^{\psi_{t}},\mathscr{M}^{\psi_{t}},\mathscr{Y}_{\psi_{t}},C^{\psi_{t}},F^{\psi_{t}},\Psi\right)$, $1\leqslant h\leqslant H$. In each economy $\mathcal{E}^{h}$, let $n_{hd}$, $n_{hs}$, and $n_{hf}$ denote, respectively, the number of consumers, producers, and financial institutions. The total number of agents in economy $h$ is then given by $n_{h}=n_{hd}+n_{hs}+n_{hf}+1$. We further define the aggregate number of each category across all economies as follows:
$N_{H}=\sum_{h=1}^{H}n_{h},\quad N_{HD}=\sum_{h=1}^{H}n_{hd},\quad N_{HS}=\sum_{h=1}^{H}n_{hs},\quad N_{HF}=\sum_{h=1}^{H}n_{hf}.$ (68)
Each economy operates with its own currency, leading to the formation of $H$ distinct currency markets, where exchange rates and cross-border transactions emerge as central mechanisms in the extended equilibrium structure.
We denote the initial endowments of an individual $i_{h}$ in economy $h$ by
$W^{i_{h},\psi_{t}}=\left(w^{i_{h},\psi_{t_{0}}}(t_{0}),\underline{m}^{i_{h},\psi_{t_{0}}}(t_{0}),\dots,w^{i_{h},\psi_{T}}(T),\underline{m}^{i_{h},\psi_{T}}(T)\right)\,$,
where the initial money holdings at time $t$ are given by the vector of holdings across all currencies: $\underline{m}^{i_{h},\psi_{t}}(t)=\left(\underline{m}^{i_{h},\psi_{t}}_{1}(t),\dots,\underline{m}^{i_{h},\psi_{t}}_{H}(t)\right)$. For each period, let $p^{i_{h}}$ represent the price vector faced by consumer $i_{h}$, which consists of both the prices of goods and the exchange rates of currencies. Specifically, $p^{i_{h},G}(t)=\left(p^{i_{h}}_{1}(t),\dots,p^{i_{h}}_{L+H}(t)\right)$ denotes the prices of goods, while $p^{i_{h},M}(t)=\left(p^{i_{h}}_{L+1}(t),\dots,p^{i_{h}}_{L+H}(t)\right)$ corresponds to the exchange rates of currencies.
We assume that economy $h$ has a unique currency, which is normalized to $1$ within that economy at equilibrium. However, this normalization does not necessarily extend to other economies, where relative currency values may fluctuate. Define the matrix of potential goods transactions for agent $i_{h}$ at time $t$ as $X^{i_{h}}(t)=\left(x^{i_{h}}_{j{i_{h}}}(t),x^{i_{h}}_{{i_{h}}j}(t)\right)$. The exchange capacities in this economy are given by: $C^{i,\psi_{t}}_{t}=\left(c^{i,\psi_{t}}_{ji}(t),c^{i,\psi_{t}}_{ij}(t)\right)$, for all $1\leqslant i,j\leqslant N_{HD}+N_{HS}$, $j\neq i$, $t_{0}\leqslant t\leqslant T$.
The goods allocation for agent $i_{h}$ at time $t$ is denoted by $x^{i_{h}}(t)$. The feasible set of transactions is constrained by a topological condition, ensuring that transactions remain within expected exchange capacities: $\mathfrak{T}^{i_{h}}=\left\{X^{i_{h}}\in(\mathbb{R}_{+}^{L})^{2(N_{HD}+N_{HS})}\mid X^{i_{h}}(t)\leqslant\mathbb{M}_{i_{h},t_{0}}\left(C^{i_{h},\psi_{t}}_{t}\right)\right\}$.
At time $t$, we denote the money holdings of individual $i_{h}$} by $m^{i_{h}}(t)=\left(m^{i_{h}}_{1}(t),\dots,m^{i_{h}}_{H}(t)\right)$. We reintroduce the financial transactions matrix of each agent $i_{h}$} as $D^{i_{h}}(t)=\left(d^{i_{h}}_{j{i_{h}}}(t),d^{i_{h}}_{{i_{h}}j}(t)\right)$, where $d^{i_{h}}_{j{i_{h}}}(t)$ represents the financial inflows from agent $j$ to agent $i_{h}$, and $d^{i_{h}}_{{i_{h}}j}(t)$ represents the outflows from $i_{h}$ to $j$. The financial capacities in the economy are given by $F^{i,\psi_{t}}_{t}=\left(f^{i,\psi_{t}}_{ji}(t),f^{i,\psi_{t}}_{ij}(t)\right)$. The corresponding financial constraint for an agent $i_{h}$} ensures that financial transactions remain within expected limits: $\mathfrak{D}^{i_{h}}=\left\{D^{i_{h}}\in(\mathbb{R}_{+}^{L})^{2N_{HF}}\mid D^{i_{h}}(t)\leqslant\mathbb{M}_{i_{h},t_{0}}\left(F^{i_{h},\psi_{t}}_{t}\right)\right\}$, for all $1\leqslant i,j\leqslant N_{HF}$, with $j\neq i$, and $t_{0}\leqslant t\leqslant T$.
Interbank flows are denoted by $(d^{i}_{ji}(t),d^{i}_{ij}(t),d^{j}_{ji}(t),d^{j}_{ij}(t))_{t_{0}\leqslant t\leqslant T,\ 1\leqslant i,j\leqslant N_{HF},\ i\neq j}$, representing financial transactions between different financial institutions. Similarly, flows between financial institutions and the central banks $(\mathfrak{C}_{h})_{1\leqslant h\leqslant H}$ are given by $(d^{i}_{c_{h}i}(t),d^{i}_{ic_{h}}(t),d^{c_{h}}_{c_{h}i}(t),d^{c_{h}}_{ic_{h}}(t))_{t_{0}\leqslant t\leqslant T,\ 1\leqslant i\leqslant N_{HF}}$, where $g_{ht}$ is a continuous function representing the money creation rule determined by the central bank $\mathfrak{C}_{h}$ at time $t$.
To account for currency exchange, we introduce the currency flow matrix $Z^{i_{h}}(t)=\left(z^{i_{h}}_{j{i_{h}}}(t),z^{i_{h}}_{{i_{h}}j}(t)\right)$, $1\leqslant j\leqslant N_{H}$, $j\neq{i_{h}}$, $t_{0}\leqslant t\leqslant T$, where $z^{i_{h}}_{j{i_{h}}}(t)=\left(z^{i_{h}}_{j{i_{h}},1}(t),\dots,z^{i_{h}}_{j{i_{h}},H}(t)\right)$ denotes the vector of currency transactions, with $z^{i_{h}}_{j{i_{h}},l}(t)$ representing the quantity of currency $l$ that agent $i_{h}$ wishes to buy from agent $j$ at time $t$. The capacities for currency transactions are given by $E^{i,\psi_{t}}_{t}=\left(e^{i,\psi_{t}}_{ji}(t),e^{i,\psi_{t}}_{ij}(t)\right)$, and the feasible set of currency exchanges is constrained by $\mathfrak{E}^{i_{h}}=\left\{Z^{i_{h}}\in(\mathbb{R}_{+}^{L})^{2N_{H}\times H}\ \mid\ Z^{i_{h}}(t)\leqslant\mathbb{M}_{i_{h},t_{0}}\left(E^{i_{h},\psi_{t}}_{t}\right)\right\}$.
Each agent $i\neq j_{h}$ can incur debts, which, in the case of consumers and producers, is explicitly specified as
$\begin{split}d^{i}(t)&=\sum_{h=1}^{H}\left(\sum_{j_{h}\neq i}p^{j_{h}}(t)\cdot x_{{j_{h}}i}^{i}(t)-\sum_{{j_{h}}\neq i}p^{i}(t)\cdot x_{i{j_{h}}}^{i}(t)-\sum_{{j_{h}}=1}^{n_{hs}}\pi_{{j_{h}}i,t}-\sum_{{l_{h}}=1}^{n_{hf}}\pi_{{l_{h}}i,t}\right)\\ &+\sum_{h=1}^{H}\left(\sum_{{j_{h}}\neq i}p^{j_{h}}_{L+h}(t)\,z^{i}_{{j_{h}}i,h}(t)-\sum_{{j_{h}}\neq i}p^{i}_{L+h}(t)\,z^{i}_{i{j_{h}},h}(t)\right)\\ &+\sum_{h=1}^{H}\left(\sum_{{j_{h}}=1}^{n_{hf}}\left[(1+r^{j_{h}}(t))d^{i}_{{j_{h}}i}(t-1)-(1+r^{i}(t))d^{i}_{i{j_{h}}}(t-1)\right]\right)\,.\end{split}$ (69)
We retain the uncertainty framework and notations introduced in Section 12. The problem faced by financial institutions is described as follows
$\max_{\displaystyle(r^{i_{hf}},D^{i_{hf}})\in\mathcal{F}^{i_{hf}}(r^{\neq{i_{hf}}}_{t_{0}},D^{\neq{i_{hf}}}_{t_{0}})\cap\mathfrak{D}^{i_{hf}}}\ \sum_{t=t_{0}}^{T}\beta^{i_{hf}}(t)\,\mathbb{M}_{i_{hf},t_{0}}\left(\pi^{i_{hf},\psi_{t}}_{t}(r^{i_{hf}},D^{i_{hf}})\right)\,$, (70)
where $\pi^{i_{hf},\psi_{t}}_{t}(r^{i_{hf}},D^{i_{hf}})=r^{i_{hf}}(t\mid t_{0})\sum_{j\neq{i_{hf}}}d^{i_{hf}}_{{i_{hf}}j}(t-1\mid t_{0})-\sum_{j\neq{i_{hf}}}r^{j,\psi_{t}}(t\mid t_{0})\,d^{i_{hf}}_{j{i_{hf}}}(t-1\mid t_{0})$.
The constraint imposed on the financial institution is given by
$\begin{split}\mathcal{F}^{i_{hf}}(r^{\neq{i_{hf}}}_{t_{0}},D^{\neq{i_{hf}}}_{t_{0}})=\Bigg\{&(r^{i_{hf}},D^{i_{hf}})\in\left([0,1]\times(\mathbb{R}_{+}^{L})^{2N_{H}}\right)^{T+1}\ \mid\ d^{i_{hf}}(T)\leqslant 0\ ,\ \sum_{j\neq{i_{hf}}}d^{i_{hf}}_{{i_{hf}}j}(t\mid t_{0})\leqslant g_{ht}\left({d^{i_{hf}}(t\mid t_{0})}^{-}\right)\ ,\\ &d^{i_{hf}}_{j{i_{hf}}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hf},t_{0}}\left(d^{j,\psi_{t}}_{j{i_{hf}}}(t\mid t_{0})\right)\ ,\ d^{i_{hf}}_{{i_{hf}}j}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hf},t_{0}}\left(d^{j,\psi_{t}}_{j{i_{hf}}}(t\mid t_{0})\right)\ ,\\ &p^{i_{hf}}_{L+h}(t\mid t_{0})=\mathbb{M}_{i_{hf},t_{0}}(p^{j_{h},\psi_{t}}_{L+h}(t\mid t_{0}))\ ,\ j_{h}\neq{i_{hf}}\Bigg\}\,.\end{split}$ (71)
The problem faced by the producer is given by
$\max_{\displaystyle(p^{i_{hs}},r^{i_{hs}},X^{i_{hs}},y^{i_{hs}})\in\mathcal{P}^{i_{hs}}(p^{\neq{i_{hs}}}_{t_{0}},r^{\neq{i_{hs}}}_{t_{0}},X^{\neq{i_{hs}}}_{t_{0}},D^{\neq{i_{hs}}}_{t_{0}})\cap\mathfrak{T}^{i_{hs}}\cap\mathfrak{D}^{i_{hs}}\cap\mathfrak{M}^{i_{hs}}}\sum_{t=t_{0}}^{T}\beta^{i_{hs}}(t)\,\mathbb{M}_{i_{hs},t_{0}}\left(\pi^{i_{hs},\psi_{t}}_{t}(p^{i_{hs}},r^{i_{hs}},X^{i_{hs}})\right)\,$, (72)
where
$\pi^{i_{hs},\psi_{t}}_{t}(p^{i_{hs}},r^{i_{hs}},X^{i_{hs}})$ $=\left(\sum_{j\neq{i_{hs}}}p^{i_{hs}}(t\mid t_{0})\cdot x_{{i_{hs}}j}^{i_{hs}}(t\mid t_{0})-\sum_{j\neq{i_{hs}}}p^{j,\psi_{t}}(t\mid t_{0})\cdot x_{j{i_{hs}}}^{i_{hs}}(t\mid t_{0})\right)$
$+\sum_{j=1}^{n_{f}}\left[(1+r^{i_{hs}}(t\mid t_{0}))d^{i_{hs}}_{{i_{hs}}j}(t-1\mid t_{0})-(1+r^{j,\psi_{t}}(t\mid t_{0}))d^{i_{hs}}_{j{i_{hs}}}(t-1\mid t_{0})\right]\,$,
The production constraint is given by
$\begin{split}\mathcal{P}^{i_{hs}}(p^{\neq{i_{hs}}}_{t_{0}},r^{\neq{i_{hs}}}_{t_{0}},X^{\neq{i_{hs}}}_{t_{0}},D^{\neq{i_{hs}}}_{t_{0}})=\Bigg\{&(p^{i_{hs}},r^{i_{hs}},X^{i_{hs}},y^{i_{hs}})\in\left(P\times[0,1]\times(\mathbb{R}_{+}^{L})^{2(N_{HD}+N_{HS})}\times{\mathbb{R}}^{L}\right)^{T+1}\ \Bigg|\ d^{i_{hs}}(T)\leqslant 0\ ,\\ &\left(y^{i_{hs}}(t+\tau\mid t_{0}),-\tau\right)\in\mathbb{M}_{i_{hs},t_{0}}(\mathscr{Y}^{i_{hs},\psi_{t}})\ ,\ S^{i_{hs}}_{t}\in\mathbb{M}_{i_{hs},t_{0}}(\mathscr{S}^{i_{hs},\psi_{t}})\ ,\\ &x_{{i_{hs}}j}^{i_{hs}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hs},t_{0}}(x_{{i_{hs}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ x_{j{i_{hs}}}^{i_{hs}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hs},t_{0}}(x_{j{i_{hs}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &d_{{i_{hs}}j}^{i_{hs}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hs},t_{0}}(d_{{i_{hs}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ d_{j{i_{hs}}}^{i_{hs}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hs},t_{0}}(d_{j{i_{hs}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &p^{i_{hs}}_{L+h}(t\mid t_{0})=\mathbb{M}_{i_{hs},t_{0}}(p^{j_{h},\psi_{t}}_{L+h}(t\mid t_{0}))\ ,\ j_{h}\neq i_{hs}\Bigg\}\,.\end{split}$ (73)
The consumer’s updated intertemporal utility maximization problem is
$\max_{\displaystyle(p^{i_{hd}},r^{i_{hd}},X^{i_{hd}},Z^{i_{hd}})\in\mathcal{T}^{i_{hd}}(p^{\neq{i_{hd}}}_{t_{0}},r^{\neq{i_{hd}}}_{t_{0}},X^{\neq{i_{hd}}}_{t_{0}},D^{\neq{i_{hd}}}_{t_{0}},Z^{\neq{i_{hd}}}_{t_{0}})\cap\mathfrak{T}^{i_{hd}}\cap\mathfrak{D}^{i_{hd}}\cap\mathfrak{E}^{i_{hd}}}\ \sum_{t=t_{0}}^{T}\beta^{i_{hd}}(t)\,\mathbb{M}_{i_{hd},t_{0}}\left(u^{i_{hd},\psi_{t}}_{t}(x^{i_{hd}},m^{i_{hd}})\right)\,$, (74)
The intertemporal budget constraint is defined as follows
$\begin{split}\mathcal{T}^{i_{hd}}(p^{\neq{i_{hd}}}_{t_{0}},r^{\neq{i_{hd}}}_{t_{0}},X^{\neq{i_{hd}}}_{t_{0}},D^{\neq{i_{hd}}}_{t_{0}},Z^{\neq{i_{hd}}}_{t_{0}})=\Bigg\{&(p^{i_{hd}},r^{i_{hd}},X^{i_{hd}},Z^{i_{hd}})\in\left(P\times[0,1]\times(\mathbb{R}_{+}^{L})^{2(N_{HD}+N_{HS})}\times({\mathbb{R}}_{+})^{2N_{HD}}\right)^{T+1}\ \Bigg|\ \\ &d^{i_{hd}}(T)\leqslant 0\ ,\ m^{i_{hd}}(t\mid t_{0})\in\mathbb{M}_{i_{hd},t_{0}}(\mathscr{M}^{i_{hd},\psi_{t}})\ ,\\ &\mathbb{M}_{i_{hd},t_{0}}(w^{i_{hd},\psi_{t}}_{t})+\sum_{j\neq{i_{hd}}}x_{j{i_{hd}}}^{i_{hd}}(t\mid t_{0})-\sum_{j\neq{i_{hd}}}x_{{i_{hd}}j}^{i_{hd}}(t\mid t_{0})\in\mathbb{M}_{i_{hd},t_{0}}(\mathscr{X}^{i_{hd},\psi_{t}})\ ,\\ &x_{{i_{hd}}j}^{i_{hd}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hd},t_{0}}(x_{{i_{hd}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ x_{j{i_{hd}}}^{i_{hd}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hd},t_{0}}(x_{j{i_{hd}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &d_{{i_{hd}}j}^{i_{hd}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hd},t_{0}}(d_{{i_{hd}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ d_{j{i_{hd}}}^{i_{hd}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hd},t_{0}}(d_{j{i_{hd}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &z_{{i_{hd}}j}^{i_{hd}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hd},t_{0}}(z_{{i_{hd}}j}^{j,\psi_{t}}(t\mid t_{0}))\ ,\ z_{j{i_{hd}}}^{i_{hd}}(t\mid t_{0})\leqslant\mathbb{M}_{i_{hd},t_{0}}(z_{j{i_{hd}}}^{j,\psi_{t}}(t\mid t_{0}))\ ,\\ &p^{i_{hd}}_{L+h}(t\mid t_{0})=\mathbb{M}_{i_{hd},t_{0}}(p^{j_{h},\psi_{t}}_{L+h}(t\mid t_{0}))\ ,\ j_{h}\neq i_{hd}\Bigg\}\,.\end{split}$ (75)
The application of Berge’s Maximum Theorem and the generalized game proposition does not require any additional treatment. It is important to note that we arrive at an initial price system $p^{i_{h}}(t_{0})=(p^{i_{h}}_{1}(t_{0}),\dots,p^{i_{h}}_{L}(t_{0}),p^{i_{h}}_{L+1}(t_{0}),\dots,p^{i_{h}}_{L+H}(t_{0}))$ at equilibrium. The normalization condition for each country is given by $\tilde{p}^{i_{h}}_{L+h}(t_{0})=\frac{p^{i_{h}}_{L+h}(t_{0})}{p^{h}_{L+h}(t_{0})}=\frac{p^{h}_{L+h}(t_{0})}{p^{h}_{L+h}(t_{0})}=1$ for every $h\in H$, which results in the exchange rates $e^{i_{h}}_{hl}=\tilde{p}^{i_{h}}_{L+l}(t_{0})=\frac{p^{i_{h}}_{L+l}(t_{0})}{p^{h}_{L+h}(t_{0})}$.
A general money quantity system can also be derived at equilibrium for an economy $h^{\ast}$} by considering one-way flows, summing over the individuals in the economy, and decomposing the debt equation into internal and external components, as follows;
$-\sum_{h\neq{h^{\ast}}}\sum_{i_{h}=1}^{n_{h}}d^{i_{h}}(t)-d^{c_{h^{\ast}}}(t)-\sum_{h\neq{h^{\ast}}}d^{c_{h}}(t)-\sum_{i_{h^{\ast}f}=1}^{n_{{h^{\ast}}f}}d^{i_{h^{\ast}f}}(t)-\sum_{h\neq{h^{\ast}}}\sum_{{i_{hf}}=1}^{n_{hf}}d^{i_{hf}}(t)$
$=\sum_{i=1}^{n_{{h^{\ast}}d}+n_{{h^{\ast}}s}}\left(\sum_{j_{h^{\ast}}\neq i}p^{j_{h^{\ast}}}(t)\cdot x_{{j_{h^{\ast}}}i}^{i}(t)-\sum_{{j_{h^{\ast}}}\neq i}m_{{j_{h^{\ast}}}i}(t)\right)+\sum_{i=1}^{n_{{h^{\ast}}d}}\left(\sum_{{j_{h^{\ast}}}\neq i}p^{j_{h^{\ast}}}_{L+{h^{\ast}}}(t)\,z^{i}_{{j_{h^{\ast}}}i,{h^{\ast}}}(t)-\sum_{{j_{h^{\ast}}}\neq i}p^{i}_{L+{h^{\ast}}}(t)\,z^{i}_{i{j_{h^{\ast}}},{h^{\ast}}}(t)\right)$
$+\sum_{i=1}^{n_{{h^{\ast}}d}+n_{{h^{\ast}}s}}\left(\sum_{{j_{h^{\ast}}}=1}^{n_{{h^{\ast}}f}}\left[(1+r^{j_{h^{\ast}}}(t))d^{i}_{{j_{h^{\ast}}}i}(t-1)-(1+r^{i}(t))d^{i}_{i{j_{h^{\ast}}}}(t-1)\right]\right)$
$+\sum_{i=1}^{n_{{h^{\ast}}d}+n_{{h^{\ast}}s}}\sum_{h\neq{h^{\ast}}}\left(\sum_{j_{h}\neq i}p^{j_{h}}(t)\cdot x_{{j_{h}}i}^{i}(t)-\sum_{j\neq{i_{d}}}p^{i}(t)\cdot x_{ij_{h}}(t)-\sum_{j_{h}=1}^{n_{hs}}\pi_{j_{h}i,t}-\sum_{l_{h}=1}^{n_{hf}}\pi_{l_{h}i,t}\right)$
$+\sum_{i=1}^{n_{{h^{\ast}}d}}\sum_{h\neq{h^{\ast}}}\left(\sum_{{j_{h}}\neq i}p^{j_{h}}_{L+h}(t)\,z^{i}_{{j_{h}}i,h}(t)-\sum_{{j_{h}}\neq i}p^{i_{h}}_{L+h}(t)\,z^{i}_{i{j_{h}},h}(t)\right)$
$+\sum_{i=1}^{n_{{h^{\ast}}d}+n_{{h^{\ast}}s}}\sum_{h\neq{h^{\ast}}}\left(\sum_{{j_{h}}=1}^{n_{hf}}\left[(1+r^{j_{h}}(t))d^{i}_{{j_{h}}i}(t-1)-(1+r^{i}(t))d^{i}_{i{j_{h}}}(t-1)\right]\right)\,.$
Thus, after normalizing by the money price of the economy $h^{\ast}$}, one can deduce the following equation
[…]
where
$\hat{v}^{l}_{m}(t)=\frac{\sum_{i=1}^{n_{{h^{\ast}}d}+n_{{h^{\ast}}s}}\sum_{{j_{h^{\ast}}}\neq i}m_{{j_{h^{\ast}}}i}(t)}{\overline{M}^{h}(t)}$
is the money velocity of economy $h$ at economy $l$,
$\hat{v}^{l}_{d}(t)=\frac{\displaystyle-\sum_{h\neq{h^{\ast}}}\sum_{i_{h}=1}^{n_{h}}d^{i_{h}}(t)-d^{c_{h^{\ast}}}(t)-\sum_{h\neq{h^{\ast}}}d^{c_{h}}(t)-\sum_{i_{h^{\ast}}=1}^{n_{{h^{\ast}}f}}d^{i_{h^{\ast}f}}(t)-\sum_{h\neq{h^{\ast}}}\sum_{{i_{hf}}=1}^{n_{hf}}d^{i_{hf}}(t)-\sum_{i=1}^{n_{{h^{\ast}}d}+n_{{h^{\ast}}s}}\sum_{h=1}^{H}\sum_{{j_{h}}=1}^{n_{hf}}\left[d^{i}_{{j_{h}}i}(t-1)-d^{i}_{i{j_{h}}}(t-1)\right]}{\overline{M}^{h}(t)}$
is the debt velocity of economy $h$ at economy $l$, and $\overline{M}^{h}(t)=\underline{M}^{h}(t)-\left(d^{c_{h}}(t)+\sum_{i_{h}=1}^{n_{hf}}d^{i_{f}}(t)\right)=\underline{M}(t)+\underline{D}(t)$. We can express the system as:
$p^{I}\cdot C+p^{E}\cdot IM-p^{I}\cdot EX-\Pi^{E}+\Delta(r^{I}\cdot D^{I})+\Delta(r^{E}\cdot D^{E})+e^{I}\cdot\Delta(Z^{I})+e^{E}\cdot\Delta(Z^{E})=Mv\,$,
where
  • $C$: internal consumption.
  • IMIM: imports.
  • EXEX: exports.
  • $\Pi^{E}$: foreign profits.
  • $D^{I}$: internal debt.
  • $D^{E}$: external debt.
  • $r^{I}$: internal interest rate.
  • $r^{E}$: external interest rate.
  • $\Delta(Z^{I})$: internal currency transaction sold.
  • $\Delta(Z^{E})$: external currency transaction sold.
  • $e^{I}$: internal exchange rate.
  • $e^{E}$: external exchange rate.
14 Conclusion
In this paper, we introduced the effective trade model, establishing the fundamental properties of the price-demand correspondence and proving the existence of Nash equilibria. We examined welfare implications, market dynamics, and distortions arising from indivisibilities and market topology. Additionally, we incorporated production and monetary factors, deriving the quantity equation of money and illustrating our theoretical findings with numerical simulations.
Expanding our analysis, we investigated the role of time in economic interactions, demonstrating how a time market naturally emerges as loanable funds suppliers interact with production plans that account for time as a productive resource. We then integrated uncertainty into the model, showing that while trade equilibrium remains attainable, supply and demand rationing may persist. Finally, we extended our framework to open economies, analyzing the formation of exchange rates and their broader implications for economic stability and policy.
Several fundamental conclusions emerge from our analysis:
  • Markets operate based on transaction logic rather than mere supply and demand desires. The feasibility of transactions dictates trade possibilities, challenging conventional equilibrium perspectives.
  • Prices are subjective and determined through simultaneous agent planning. There is no universal market-clearing mechanism but rather a strategic interplay of decisions.
  • Equilibrium uniqueness is not guaranteed, and autarky remains a persistent equilibrium. This reinforces the idea that decentralized coordination does not necessarily lead to optimal trade.
  • The Pareto-Nash dichotomy reshapes the philosophical foundation of general equilibrium. Pareto efficiency and Nash equilibrium emerge as distinct concepts, challenging the traditional view that individual rationality automatically leads to social welfare.
  • Market topology introduces complications that can reduce overall welfare. Trade feasibility depends on network structure and accessibility rather than mere preferences and endowments.
  • Time should be recognized as a production factor. Decisions made today influence future production outcomes, shaping intertemporal resource allocation.
  • The interest rate can be interpreted as the price of time. The interaction between present and future consumption clarifies the role of interest in economic dynamics.
  • Agent anticipation is formed by the mode, incorporating personal information, beliefs, and limited memory. The use of the mean as an expectation does not accurately reflect real-world behavior. Instead, anticipation should reflect subjective judgment, bounded rationality, and memory constraints. In particular, agents are assumed to base their forecasts on the most likely outcome—captured by the mode—which aligns more closely with Keynesian notions of "animal spirits" [Key36].
  • The time gap between production decisions and exchange leads to inherent time inconsistency. Even when trade equilibrium is achieved, delayed responses and adjustments create instability in economic planning.
  • The quantity equation of money requires reinterpretation. Our framework suggests that money circulation depends not only on supply and demand but also on structural constraints, transaction feasibility, and intertemporal adjustments.
These findings challenge classical general equilibrium theory and open new perspectives on economic coordination, strategic interaction, and monetary theory. Future research could further explore the implications of these insights, particularly in dynamic and stochastic environments, to refine our understanding of real-world market mechanisms.
Appendix
14.1 Berge’s Maximum Theorem
Theorem 14.1.
Berge’s Maximum Theorem[Ber63]
Let $H$ and $\Theta$ be topological spaces, $f:H\times\Theta\to{\mathbb{R}}$ a continuous function, and $\mathcal{F}:\Theta\to 2^{H}$ a correspondence such that:
  1. 1.
    $\mathcal{F}(\theta)\subset H$ is non-empty and compact for each $\theta\in\Theta$,
  2. 2.
    $\mathcal{F}$ is continuous.
Define the value function $v:\Theta\to{\mathbb{R}}$ by
$v(\theta)$ $=\max_{x\in\mathcal{F}(\theta)}f(x,\theta).$
Then:
  1. 1.
    $v(\theta)$ is continuous on $\Theta$,
  2. 2.
    The set of maximizers $S(\theta)=\arg\max_{x\in\mathcal{F}(\theta)}f(x,\theta)$ is non-empty, compact, and upper semicontinuous.
Corollary 14.2.
[Sun96]
Let $f:H\times\Theta\to{\mathbb{R}}$ be continuous, and $\mathcal{F}:\Theta\to 2^{H}$ be continuous and compact valued. Define $v$ and $S$ as in Theorem 14.1.
  1. 1.
    Suppose $f(\cdot,\theta)$ is quasi-concave in $x$ for each $\theta$, and $\mathcal{F}$ is convex-valued on $\Theta$. Then $S$ is a convex valued upper-semicontinuous correspondence.
  2. 2.
    If "quasi-concave" is replaced with "strictly quasi-concave", $S$ is single valued everywhere on $\Theta$, and hence defines a continuous function.
Theorem 14.3.
[LRS19]
If a game $G=\{A^{i},\mathcal{F}^{i},u^{i}\}_{1\leqslant i\leqslant n}$ is compact, continuous and quasi-concave, then its set of Nash equilibria is a non-empty and compact subset of $\prod_{i=1}^{n}A^{i}$.
14.2 Generalized Maximum Theorem
Definition 14.1.
$K$-inf-compact functions [FKV14]
Let $f\,:\,H\times\Theta\to{\mathbb{R}}$ and $U$ an open set of $H\times\Theta$. Consider the level sets
$D_{f}(a,U)=\{y\in U\ :\ f(x,y)\leqslant a\}\,$,
and for $X\subset H$
$\text{Gr}_{X}=\{(x,y)\in Z\times\Theta\ :\ y\in\mathcal{F}(x)\}\,.$
$f$ is called $K$-inf-compact on $\text{Gr}_{\Theta}(\mathcal{F})$, if for every compact $K$ of $\Theta$ this function is inf-compact on $\text{Gr}_{K}(\mathcal{F})$.
Definition 14.2.
Compactly generated spaces[FKV14]
A topological space $X$ is compactly generated if it satisfies the following property: each set $A\subset X$ is closed in $X$ if $A\cap K$ is closed in $K$ for each compact $K$ of $X$.
Theorem 14.4.
[FKV14]
Assume that:
  1. 1.
    $H$ is a compactly generated topological space;
  2. 2.
    $\mathcal{F}:\Theta\to 2^{H}$ is lower semicontinuous;
  3. 3.
    $f\,:\,H\times\Theta\to{\mathbb{R}}$ is $K$-inf-compact and upper semicontinuous on $\text{Gr}_{\Theta}(\mathcal{F})$.
Then the value function $v:\Theta\to{\mathbb{R}}$ is continuous and the solution multifunction $S(\theta)=\arg\max_{x\in\mathcal{F}(\theta)}f(x,\theta)$ is upper semicontinuous and compact-valued.

### v2

@@
 Rethinking Arrow–Debreu: A New Framework for Exchange, Time, and Uncertainty
-Rethinking Arrow–Debreu: A New Framework for Exchange, Time, and Uncertainty
-Nizar Riane
+Abstract
+This paper revisits the Arrow-Debreu general equilibrium framework through the lens of effective trade, emphasizing the distinction between theoretical and realizable market interactions.
+We develop the Effective Trade Model (ETM), where transactions arise from bilateral feasibility rather than aggregate supply and demand desires.
+Within this framework, we establish the main properties of the price-demand correspondence and prove the existence of Nash equilibria, incorporating production, money, and network topology.
+The analysis extends to time, uncertainty, and open economies, revealing how loanable funds and exchange rates emerge endogenously.
+Our results show that equilibrium is shaped by transaction constraints, subjective pricing, and decentralized negotiation, rather than by universal market-clearing conditions, and thereby call into question the foundations of welfare theory.
+Anticipation is modeled via the conditional mode, capturing bounded rationality and information limitations in contrast to the rational expectations hypothesis.
+The ETM thus offers a behaviorally and structurally grounded alternative to classical general equilibrium, bridging microfoundations, monetary dynamics, and temporal consistency within a unified framework.
 Université Mohammed V de Rabat, Maroc111nizar.riane@gmail.com
 Keywords: General equilibrium – Welfare – Network – Convergence – Quantity theory of money – Uncertainty – Open economy.
-JEL Classification:
+JEL Classification: C62 – D51 – D52 – D61 – D84
 Acknowledgements: The author would like to thank Michel Lapidus and Michael Maroun for their insightful comments, which greatly contributed to improving this work.
 1 Introduction
-The Arrow–Debreu Model (ADM) [AD54] has long been a cornerstone of modern economic theory, providing a rigorous mathematical framework for analyzing microeconomic decisions.
-It offers a formal resolution to an old question posed by Léon Walras in [Wal74], leveraging advancements in the theory of multifunctions and fixed points.
+The Arrow–Debreu Model (ADM) [Arrow and Debreu, 1954] has long been a cornerstone of modern economic theory, providing a rigorous mathematical framework for analyzing microeconomic decisions.
+It offers a formal resolution to an old question posed by Léon Walras in [Walras, 1874], leveraging advancements in the theory of multifunctions and fixed points.
 While the ADM is meticulously formulated, it has not been universally embraced by economists.
@@
 This has led some scholars, such as Mark Blaug, to express strong reservations.
-Blaug famously characterized Arrow and Debreu’s seminal paper [AD54] as marking "the beginning of what has since become a cancerous growth in the very center of microeconomics" (see [Bla98]).
+Blaug famously characterized Arrow and Debreu’s seminal paper [Arrow and Debreu, 1954] as marking "the beginning of what has since become a cancerous growth in the very center of microeconomics" (see [Blaug, 1998]).
 He further argued that "by the time we got to Arrow and Debreu, general equilibrium theory had ceased to make any descriptive claim about actual economic systems and had become a purely formal apparatus about a quasi-economy".
-His criticism extended to Gérard Debreu’s book, The Theory of Value [Deb59], which he deemed "probably the most arid and pointless book in the entire literature of economics".
+His criticism extended to Gérard Debreu’s book, The Theory of Value [Debreu, 1959], which he deemed "probably the most arid and pointless book in the entire literature of economics".
 Beyond Blaug’s critiques of its theoretical abstraction, it is important to recognize that general equilibrium theory has undergone significant refinements since the publication of Arrow and Debreu’s 1954 paper.
-Debreu’s The Theory of Value [Deb59] further formalized the framework, particularly in economies facing uncertainty.
-The role of money within the model has been extensively debated, with insightful contributions from Robert Clower [Clo67], Frank Hahn [Hah89], and Joseph M.
+Debreu’s The Theory of Value [Debreu, 1959] further formalized the framework, particularly in economies facing uncertainty.
+The role of money within the model has been extensively debated, with insightful contributions from Robert Clower [Clower, 1967], Frank Hahn [Hahn, 1989], and Joseph M.
 Ostroy and Ross M.
-Starr [OS74] and others.
-The question of equilibrium convergence was explored using nontraditional techniques and more realistic approaches (see, for example, [HN62]).
-Additionally, time was explicitly incorporated into the system [Rad72], and market imperfections-such as equilibrium with indivisible goods-were examined to better align the theory with economic realities [Sve84].
-In his book The Methodology of Economics: Or, How Economists Explain [Bla92], Mark Blaug provocatively asks whether general equilibrium theory should be regarded as an empirical theory or merely as a conceptual framework.
+Starr [Ostroy and Starr, 1974] and others.
+The conceptualization of transaction was reformulated by L.
+Shapley and M.
+Shubik through the introduction of trading posts in [ShapleyShubik1977].
+The question of equilibrium convergence was explored using nontraditional techniques and more realistic approaches (see, for example, [Hahn and Negishi, 1962]).
+Additionally, time was explicitly incorporated into the system [Radner, 1972], and market imperfections-such as equilibrium with indivisible goods-were examined to better align the theory with economic realities [Svensson, 1984].
+In his book The Methodology of Economics: Or, How Economists Explain [Blaug, 1992], Mark Blaug provocatively asks whether general equilibrium theory should be regarded as an empirical theory or merely as a conceptual framework.
 To explore this question, one can draw an analogy with classical physics: just as Newtonian mechanics assumes the existence of a Galilean reference frame–a perfectly inertial system in which the laws of motion take their simplest form–general equilibrium theory is grounded in idealized assumptions of complete markets and perfect competition.
@@
 This distinction underscores the importance of addressing the outcomes of exchanges rather than unattainable aspirations.
-Markets are not arenas for idealized redistribution but mechanisms for facilitating feasible transactions shaped by constraints and negotiation.
+Markets are not arenas for idealized redistribution but mechanisms for facilitating feasible transactions shaped by constraints, negotiation, and network interactions.
 Real-world exchanges also separate the roles of buying and selling.
@@
 Moreover, the quantity exchanged in a transaction is inherently constrained.
-It is determined by the minimum that both parties can support, reflecting their respective limitations.
+It is determined by the minimum of what both parties can support, reflecting their respective limitations.
 This contrasts with theoretical models that emphasize optimal allocations without considering such practical constraints.
@@
 In light of these limitations, we propose a departure from the traditional use of expectations and instead define anticipation as the conditional mode—that is, the most likely realization of a variable given the agent’s information, memory, and subjective beliefs.
-This approach better reflects the role of instinctive judgment, focal outcomes, and perceived salience, aligning with Keynes’ notion of "animal spirits" [Key36].
+This approach better reflects the role of instinctive judgment, focal outcomes, and perceived salience, aligning with Keynes’ notion of "animal spirits" [Keynes, 1936].
 By grounding anticipations in the mode rather than the mean, this formulation introduces behavioral realism without sacrificing formal rigor.
@@
 A strict inequality would imply that the customer accepted an exchange in which higher-valued goods were traded for lower-valued ones, which effectively equivalent to setting a price lower than $p^{i}$.
-One might consider a condition similar to that in the Arrow–Debreu model [AD54]:
+One might consider a condition similar to that in the Arrow–Debreu model [Arrow and Debreu, 1954]:
 $p^{j}\cdot x^{i}\leqslant p^{i}\cdot w^{i}.$
@@
 Definition 2.1.
-Definition of a Generalized Game[Kre12]
+Definition of a Generalized Game[Kreps, 2012]
 A $n$-player generalized game $G=\{A^{i},\mathcal{F}^{i},u^{i}\}_{1\leqslant i\leqslant n}$ for a finite $n\in{\mathbb{N}}$ consists of, for $i=1,\ldots,n$,
@@
 Proposition 2.2.
-Equilibrium in Effective Trade Economy[Kre12]
+Equilibrium in Effective Trade Economy[Kreps, 2012]
 Suppose that $G=\{A^{i},\mathcal{F}^{i},u^{i}\}_{1\leqslant i\leqslant n}$ is a generalized game for which
@@
 Proposition 2.4.
-Compactness of $\mathbf{E}$ [LRS19]
+Compactness of $\mathbf{E}$ [Laraki et al., 2019]
 Under the assumptions a, b and c, the set $\mathcal{E}$ of transaction equilibria is non empty and compact.
@@
 Definition 2.2.
-Pareto Optimal Allocation [Sta12]
+Pareto Optimal Allocation [Starr, 2012]
 Let $\mathcal{B}^{i}\subseteq\mathbb{R}^{L}$ be the feasible set of allocations of customer $i$ and $u^{i}:\mathcal{B}^{i}\to\mathbb{R}$ be the utility function of the same agent, for $i=1,\dots,n$.
@@
 3 KKT characterization
-In this section we make the following assumption:
+We now characterize the equilibrium for a smooth economy, that is, when the relevant functions are well-behaved. We make the following assumption:
   1. d
     Differentiability: $u^{i}(\cdot)$ is differentiable on $\mathscr{X}^{i}$.
-And for simplicity, let $\mathscr{X}^{i}={\mathbb{R}}_{+}^{L}$ and denote $x^{i}$ as $x$ and $x^{j}$ as $y$, respectively.
+And for simplicity, let $\mathscr{X}^{i}={\mathbb{R}}_{+}^{L}$ and denote $x^{i}$ as $x$ and $x^{j}$ as differentiable function $y(p^{i},X^{i})$ of $p^{i}$ and $X^{i}$, respectively.
 Define the Lagrangian function $\mathcal{L}(p^{i},X,\lambda,\mu,\eta,\nu,\theta,\kappa)$ as
@@
 We have the following stationary condition
-$\frac{\partial\mathcal{L}}{\partial p^{i}_{k}}$ $=\lambda\sum_{j\neq i}x_{ij,k}+\nu_{k}+\kappa=0$
-$\frac{\partial\mathcal{L}}{\partial x_{ij,k}}$ $=-\frac{\partial u}{\partial x_{k}}+\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}=0$
-$\frac{\partial\mathcal{L}}{\partial x_{ji,k}}$ $=\frac{\partial u}{\partial x_{k}}-\lambda p^{j}_{k}+\mu_{ji,k}-\eta_{ji,k}+\theta_{k}=0\,.$
+$\frac{\partial\mathcal{L}}{\partial p^{i}_{k}}$ $=\lambda\sum_{j\neq i}x_{ij,k}+\nu_{k}+\kappa+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial p^{i}_{k}}y_{ij,l}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial p^{i}_{k}}y_{ji,l}=0$
+$\frac{\partial\mathcal{L}}{\partial x_{ij,k}}$ $=-\frac{\partial u}{\partial x_{k}}+\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial x_{ij,k}}y_{ij,l}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial x_{ij,k}}y_{ji,l}=0$
+$\frac{\partial\mathcal{L}}{\partial x_{ji,k}}$ $=\frac{\partial u}{\partial x_{k}}-\lambda p^{j}_{k}+\mu_{ji,k}-\eta_{ji,k}+\theta_{k}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial x_{ji,k}}y_{ij,l}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial x_{ji,k}}y_{ji,l}=0\,.$
 We obtain the following situations:
 $\sum_{k=1}^{L}p^{i}_{k}\sum_{j\neq i}x_{ij,k}=\sum_{k=1}^{L}\sum_{j\neq i}p^{j}_{k}x_{ji,k}$: thus,
-$\lambda\sum_{j\neq i}x_{ij,k}$ $=-\nu_{k}-\kappa$
-$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}>0$
-$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{j}_{k}-\mu_{ji,k}+\eta_{ji,k}-\theta_{k}>0\,.$
-This implies, for $i\neq j\neq l$:
+$\lambda\sum_{j\neq i}x_{ij,k}$ $=-\nu_{k}-\kappa-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial p^{i}_{k}}y_{ij,l}-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial p^{i}_{k}}y_{ji,l}$
+$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial x_{ij,k}}y_{ij,l}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial x_{ij,k}}y_{ji,l}$
+$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{j}_{k}-\mu_{ji,k}+\eta_{ji,k}-\theta_{k}-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial x_{ji,k}}y_{ij,l}-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial x_{ji,k}}y_{ji,l}\,.$
+  1. 1.
+    This implies, for $i\neq j$, the price-quantity equation:
     $\lambda(p^{i}_{k}-p^{j}_{k})$ $=\eta_{ij,k}+\eta_{ji,k}-(\mu_{ij,k}+\mu_{ji,k})$
-$\mu_{ij,k}-\eta_{ij,k}$ $=\mu_{il,k}-\eta_{il,k}$
-$\eta_{ji,k}-\mu_{ji,k}$ $=\eta_{li,k}-\mu_{li,k}\,.$
+    $-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\left(\frac{\partial}{\partial x_{ji,k}}y_{ij,l}+\frac{\partial}{\partial x_{ji,k}}y_{ij,l}\right)-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\left(\frac{\partial}{\partial x_{ji,k}}y_{ji,l}+\frac{\partial}{\partial x_{ji,k}}y_{ji,l}\right).$
+    Interpretation:
+    3.0.0.1 Shadow valuation of price differences:
+    The difference $p^{i}_{k}-p^{j}_{k}$ measures the relative valuation of good $k$ between the two trading partners.
+The multiplier $\lambda$ is the shadow value of consumer $i$’s trade-balance constraint
+    $\sum_{k=1}^{L}p^{i}_{k}\sum_{j\neq i}x_{ij,k}-\sum_{k=1}^{L}\sum_{j\neq i}p^{j}_{k}x_{ji,k}$
+    Thus, $\lambda(p^{i}_{k}-p^{j}_{k})$ quantifies the utility-valued shadow advantage or disadvantage that arises when consumer $i$ reallocates one marginal unit of good $k$ between “selling at his own price” and “selling at $j$’s price,” ceteris paribus.
+In short, How much does $i$ benefit (in utility) from trading good $k$ at his own price rather than $j$’s price.
+    3.0.0.2 Bounds net direct marginal shadow price.
+    The multipliers $\eta_{ij,k}$ and $\eta_{ji,k}$ are the shadow prices of the upper–bound constraints $x_{ij,k}\leq y_{ij,k}$ and $x_{ji,k}\leq y_{ji,k}$ imposed by consumer $j$.
+A positive $\eta_{ij,k}$ (resp.
+$\eta_{ji,k}$) therefore measures the marginal utility gain for consumer $i$ from relaxing $j$’s cap on how much he is willing to buy (resp.
+sell).
+The multipliers $\mu_{ij,k}$ and $\mu_{ji,k}$ are the shadow prices of the lower–bound constraints $x_{ij,k}\geq 0$ and $x_{ji,k}\geq 0$; they are positive exactly when $i$ is forced to remain at a zero–flow corner and would benefit, at the margin, from allowing a flow in the opposite direction.
+    Thus
+    $\eta_{ij,k}+\eta_{ji,k}-(\mu_{ij,k}+\mu_{ji,k})$
+    is the net direct marginal shadow value of all bilateral quantity bounds on good $k$ in the transaction with $j$.
+It summarizes whether, at the margin, $i$’s welfare is primarily constrained by $j$’s upper trading limits (reflected in the $\eta$’s) or by the zero–flow lower bounds (reflected in the $\mu$’s), thereby capturing the overall tightness of the bilateral quantity restrictions faced by $i$.
+    3.0.0.3 Network indirect adjustment costs:
+    The quantities $y_{ij,\ell}$ and $y_{ji,\ell}$ denote consumer $j$’s own proposed trades.
+The derivatives $\frac{\partial y_{ij,\ell}}{\partial x_{ji,k}}$ and $\frac{\partial y_{ji,\ell}}{\partial x_{ji,k}}$ (resp.
+$\frac{\partial y_{ij,\ell}}{\partial x_{ij,k}}$ and $\frac{\partial y_{ji,\ell}}{\partial x_{ij,k}}$) measure how $j$’s proposals react when $i$ changes $x_{ji,k}$ (resp.
+$x_{ij,k}$).
+The sums
+    $-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\left(\frac{\partial}{\partial x_{ji,k}}y_{ij,l}+\frac{\partial}{\partial x_{ji,k}}y_{ij,l}\right)-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\left(\frac{\partial}{\partial x_{ji,k}}y_{ji,l}+\frac{\partial}{\partial x_{ji,k}}y_{ji,l}\right)$
+    capture the indirect marginal constraint costs that arise because altering $i$’s intended transactions of good $k$ affects the feasibility of other potential transactions through $j$’s strategic adjustments.
+Economically, these terms represent transaction-network externalities: changing one bilateral flow modifies the entire structure of feasible exchanges.
+    3.0.0.4 Economic meaning of the equation:
+    The KKT condition can be summarized as the identity
+[…]
+    Economically, this states that at an optimum consumer $i$ adjusts the bilateral flow of good $k$ with $j$ until the utility-valued marginal gain from reallocating one unit of the good across the price gap $p^{i}_{k}-p^{j}_{k}$ is exactly balanced by (i) the net shadow value of the binding quantity bounds in the $(i,j)$ relationship and (ii) the indirect constraint costs generated by how $j$’s adjustment of other proposed trades reacts to $i$’s change in this particular flow.
+    In short, the marginal price advantage of trading with $j$ at $i$’s valuation rather than $j$’s valuation is driven to equality with the total marginal feasibility cost created by bilateral quantity limits and by the propagation of this adjustment through the transaction network.
+  2. 2.
+    Let’s analyze the price equation:
+    $\lambda\sum_{j\neq i}x_{ij,k}$ $=-\nu_{k}-\kappa-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial p^{i}_{k}}y_{ij,l}-\sum_{k=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial p^{i}_{k}}y_{ji,l}\geqslant 0$
+    Interpretation:
+    3.0.0.5 Shadow impact of the price of good $k$.
+    The multiplier $\lambda$ is the shadow value of consumer $i$’s trade-balance constraint, so it converts any marginal perturbation of this constraint into utility units.
+The coefficient $\sum_{j\neq i}x_{ij,k}$ is the sensitivity of the trade-balance constraint to changes in the price component $p^{i}_{k}$.
+Thus the product $\lambda\sum_{j\neq i}x_{ij,k}$ represents the utility-valued marginal effect of increasing $p^{i}_{k}$ through its impact on consumer $i$’s trade-balance constraint.
+It is the utility-weighted exposure of the trade-balance condition to this price component.
+    3.0.0.6 Nonnegativity constraint cost.
+    The term $\nu_{k}$ is associated with the lower-bound constraint $p^{i}_{k}\geq 0$.
+If $p^{i}_{k}=0$, then $\nu_{k}>0$ indicates that consumer $i$ would strictly prefer to lower $p^{i}_{k}$ further, but the constraint prevents such a move.
+Hence $\nu_{k}$ represents the marginal utility loss created by the infeasibility of decreasing $p^{i}_{k}$ below zero; the negative sign reflects that this is a constraint-induced cost from the perspective of stationarity.
+    3.0.0.7 Normalization constraint cost.
+    The multiplier $\kappa$ enforces the unit-sum constraint $\sum_{r=1}^{L}p^{i}_{r}=1$. Raising $p^{i}_{k}$ necessarily requires reducing the weight placed on other prices. Thus
+    $\kappa$
+    captures the marginal shadow cost of transferring one unit of price weight toward $p^{i}_{k}$ while preserving the normalization constraint.
+It measures the opportunity cost of reallocating price mass across goods.
+    3.0.0.8 Network reaction costs from partners’ demand responses.
+    Consider the term
+    $\sum_{\ell=1}^{L}\sum_{j\neq i}\eta_{ij,\ell}\frac{\partial y_{ij,\ell}}{\partial p^{i}_{k}}.$
+    The quantity $y_{ij,\ell}$ is consumer $j$’s proposed purchase of good $\ell$ from $i$.
+If increasing $p^{i}_{k}$ makes $j$ reduce $y_{ij,\ell}$, then the constraint $x_{ij,\ell}\leq y_{ij,\ell}$ becomes tighter.
+The multiplier $\eta_{ij,\ell}$ is the shadow value of that constraint.
+Thus this sum captures the marginal utility loss generated by the tightening of all purchase-side feasibility constraints caused by partners reducing their intended demands in response to a higher price $p^{i}_{k}$.
+    3.0.0.9 Network reaction costs from partners’ supply responses.
+    Similarly,
+    $\sum_{\ell=1}^{L}\sum_{j\neq i}\eta_{ji,\ell}\frac{\partial y_{ji,\ell}}{\partial p^{i}_{k}}$
+    reflects how raising $p^{i}_{k}$ affects $j$’s proposed sales of good $\ell$ to $i$.
+If these sales decrease, then the constraint $x_{ji,\ell}\leq y_{ji,\ell}$ becomes more binding.
+Weighted by the multipliers $\eta_{ji,\ell}$, this term represents the corresponding marginal utility loss due to the tightening of supply-side feasibility constraints.
+    3.0.0.10 Economic meaning of the full KKT condition.
+    The stationarity condition
+    $\lambda\sum_{j\neq i}x_{ij,k}=-\nu_{k}-\kappa-\sum_{\ell=1}^{L}\sum_{j\neq i}\eta_{ij,\ell}\frac{\partial y_{ij,\ell}}{\partial p^{i}_{k}}-\sum_{\ell=1}^{L}\sum_{j\neq i}\eta_{ji,\ell}\frac{\partial y_{ji,\ell}}{\partial p^{i}_{k}}$
+    states that consumer $i$ sets the price component $p^{i}_{k}$ so that the utility-valued marginal effect of altering $p^{i}_{k}$ through the trade-balance constraint is exactly offset by the aggregate marginal shadow costs that this change induces.
+These costs consist of: (i) the marginal cost of respecting the nonnegativity constraint on prices, (ii) the marginal opportunity cost induced by the normalization constraint, and (iii) the propagated tightening of all bilateral feasibility constraints caused by trading partners’ strategic adjustments to $i$’s price change.
+    In equilibrium, $p^{i}_{k}$ is therefore chosen such that the marginal benefit of adjusting this price is precisely counterbalanced by the total marginal feasibility costs transmitted through the entire trading network.
+4 Convergence
+Next, we propose three convergence processes toward Pareto-optimal and Nash–Pareto states, based on two underlying principles:
   1. 1.
-    If $\lambda=0$, then $\mu_{ij,k}=\eta_{ji,k}>0$, which implies $x_{ij,k}=0$ and $x_{ji,k}=y_{ji,k}$ for all $i,j,k$.
-This is only possible if $x_{ij,k}=x_{ji,k}=y_{ji,k}=0$, a situation that will be excluded in the remainder of the analysis.
+    A gradient-based algorithm representing an exchange dynamic driven by agents’ needs and desires (utility gradient);
   2. 2.
-    $p^{i}_{k}>p^{j}_{k}$: then
-    $\eta_{ij,k}+\eta_{ji,k}>(\mu_{ij,k}+\mu_{ji,k})\,.$
-    The possible ways (or combination) are
-    1. (a)
-      $\eta_{ij,k}>\mu_{ji,k}\geqslant 0$ and $\mu_{ij,k}=0$: $x_{ij,k}=y_{ij,k}$.
-    2. (b)
-      $\eta_{ji,k}>\mu_{ij,k}\geqslant 0$ and $\mu_{ji,k}=0$: $x_{ji,k}=y_{ji,k}$.
-  3. 3.
-    $p^{j}_{k}>p^{i}_{k}$: then
-    $\eta_{ij,k}+\eta_{ji,k}<(\mu_{ij,k}+\mu_{ji,k})\,.$
-    The possible ways (or combination) are
-    1. (a)
-      $0\leqslant\eta_{ji,k}<\mu_{ij,k}$ and $\eta_{ij,k}=0$: $x_{ij,k}=0$.
-    2. (b)
-      $0\leqslant\eta_{ij,k}<\mu_{ji,k}$ and $\eta_{ji,k}=0$: $x_{ji,k}=0$.
-  4. 4.
-    $p^{j}_{k}=p^{i}_{k}\geqslant 0$: then
-    $\eta_{ij,k}+\eta_{ji,k}=(\mu_{ij,k}+\mu_{ji,k})\,.$
-    $\eta_{ij,k}$, $\eta_{ji,k}$, $\mu_{ij,k}$ and $\mu_{ji,k}$ cannot all be positive (a situation previously isolated). The possible ways (or combinations) are
-    1. (a)
-      $\mu_{ij,k}=\eta_{ji,k}>0$ and $\eta_{ij,k}=\mu_{ji,k}=0$: $x_{ij,k}=0$ and $x_{ji,k}=y_{ji,k}$.
-    2. (b)
-      $\mu_{ji,k}=\eta_{ij,k}>0$ and $\eta_{ji,k}=\mu_{ij,k}=0$: $x_{ji,k}=0$ and $x_{ij,k}=y_{ij,k}$.
-    3. (c)
-      $\eta_{ij,k}=\eta_{ji,k}=\mu_{ij,k}=\mu_{ji,k}=0$ and the constraints on $X$ are not satiated.
-Now let us analyze a Nash equilibrium situation: Consider the market of good $k$.
-Given a customer $i$, let designate $\mathbb{E}^{+}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}>p^{i}_{k}\}$, $\mathbb{E}^{-}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}<p^{i}_{k}\}$ and $\mathbb{E}^{=}_{k}=\{j\in\mathbb{N}\mid p^{j}_{k}=p^{i}_{k}\}$.
-  1. 1.
-    For $j\in\mathbb{E}^{-}_{k}$ (resp. $j\in\mathbb{E}^{+}_{k}$), the KKT condition for both customers implies one of the three situations:
-    1. (a)
-      $x^{i}_{ij,k}=x^{j}_{ij,k}=0$ and $x^{i}_{ji,k}=x^{j}_{ji,k}$.
-    2. (b)
-      $x^{i}_{ij,k}=x^{j}_{ij,k}$ and $x^{i}_{ji,k}=x^{j}_{ji,k}=0$.
-    3. (c)
-      $x^{i}_{ij,k}=x^{j}_{ij,k}=0=x^{j}_{ji,k}=x^{i}_{ji,k}$.
-  2. 2.
-    For $j\in\mathbb{E}^{=}_{k}$, the Nash equilibrium condition implies:
-    1. (a)
-      $0\leqslant x^{i}_{ij,k}=x^{j}_{ij,k}$ and $0\leqslant x^{i}_{ji,k}=x^{j}_{ji,k}$.
-Remark 3.1.
-To avoid the indeterminacy caused by the situation $\mathbb{E}^{=}_{k}$, particularly the occurrence of unnecessary transactions in both directions ($0<x^{i}_{ij,k}$ and $0<x^{i}_{ji,k}$) when prices are equal, one could impose the condition $x^{i}_{ij}\cdot x^{i}_{ji}=0$ for all $i,j$.
-This condition ensures that no simultaneous flows occur in opposite directions and will be implicitly assumed when considering the joint feasible set to guarantee boundedness.
-4 Convergence
-Next, we present three convergence processes toward Pareto optimal and Nash–Pareto states.
+    A non-tâtonnement adjustment process where agents trade as long as further improvement is attainable.
 4.1 Gradient direction
@@
 The projection $\pi_{\mathcal{B}(p_{t+1},X_{t+1})}(z)$ refers to the unique closest element to $z$ in the feasible set.
-A classical proof of the projected supergradient descend convergence could be found in [Sho98].
+A classical proof of the projected supergradient descend convergence could be found in [Shor, 1998].
 The convergence point $\left(p^{\star},X^{\star}\right)=\left(p^{1,\star},X^{1,\star},\ldots,p^{n,\star},X^{n,\star}\right)$ satisfies
@@
 In the following, we address the problem of demand indeterminacy by analyzing consumer preferences within an exchange network.
-Our focus is on exchange processes structured by network interactions: each agent $i$ may hold intrinsic preferences over goods, as captured by classical demand theory (see, for instance, [MCWG95]), while also developing affinities toward specific trading partners.
+Our focus is on exchange processes structured by network interactions: each agent $i$ may hold intrinsic preferences over goods, as captured by classical demand theory (see, for instance, [Mas-Colell et al., 1995]), while also developing affinities toward specific trading partners.
 These affinities—shaped by factors such as marketing, customer loyalty, market positioning, or interpersonal relationships—can lead agents to favor transactions with certain peers over others.
@@
 Proposition 7.1.
-[MCWG95]
+[Mas-Colell et al., 1995]
 A preference relation $\succeq$ can be represented by a utility function only if it is rational.
@@
 Proposition 7.2.
-Debreu’s Theorem [Kre12]
+Debreu’s Theorem [Kreps, 2012]
 If a continuous function $u$ represents $\succeq$, then $\succeq$ is continuous. Conversely, if $\succeq$ is continuous, then there exists a continuous utility function $u$ that represents it.
@@
 Proposition 7.3.
-[Kre12]
+[Kreps, 2012]
   1. 1.
@@
 Proposition 7.4.
-[Kre12]
+[Kreps, 2012]
 Suppose preferences $\succeq$ are continuous and strongly separable in $J_{1}$ through $J_{N}$.
@@
 From a trade perspective, one could imagine that the seller, depending on their position in the market, could have different preferences.
-For instance, a large seller might seek to gain more influence in the market and thus maximize their spectral influence (see [Ria24]), while another agent might aim to maximize their prestige by forming relationships with larger sellers and thus maximize their eigenvector centrality.
-For more details on centrality measures, consult [BJT23].
+For instance, a large seller might seek to gain more influence in the market and thus maximize their spectral influence (see [Riane, 2024]), while another agent might aim to maximize their prestige by forming relationships with larger sellers and thus maximize their eigenvector centrality.
+For more details on centrality measures, consult [Bloch et al., 2023].
 Marketing and social determinants can also influence the form of $v$.
@@
 It should not be confused with intrinsic value.
-[^2] The distinction between use value and exchange value can be traced back to Adam Smith [Smi76].
+[^2] The distinction between use value and exchange value can be traced back to Adam Smith [Smith, 1776].
 Furthermore, the topology plays a crucial role in shaping equilibria, influencing welfare losses, and highlighting imperfections caused by intermediaries.
@@
 Here, $w^{i}$ and $\underline{m}^{i}$ represent customer $i$’s initial holdings of goods and money, respectively.
-The realistic nature of transactions makes the Clower cash-in-advance hypothesis unnecessary [Clo67], since each individual only demands goods that they can afford, and every transaction is equivalent to a monetary flow.
+The realistic nature of transactions makes the Clower cash-in-advance hypothesis unnecessary [Clower, 1967], since each individual only demands goods that they can afford, and every transaction is equivalent to a monetary flow.
 When analyzing the monetary flows, the customer satisfies the following transaction balance condition
@@
 where $S^{i_{s}}_{t}=\sum_{j\neq i_{s}}\left(p^{i_{s}}(t)\cdot x^{i_{s}}_{i_{s}j}(t)-p^{j}(t)\cdot x^{i_{s}}_{ji_{s}}(t)\right)+(1-\delta)S^{i_{s}}_{t-1}$ denotes the producer’s stock of commodities at time $t$; $\delta$ is the depreciation rate of the stock; $\mathscr{S}^{i_{s}}_{t}$ is a closed, bounded subset of $\mathbb{R}_{+}^{L}$ representing storage capacity constraints; and $\tau=(\tau_{1},\dots,\tau_{L})$ is a vector of time requirements for producing the outputs $y^{i_{s}}(t+\tau)$, with time being treated as a production factor.
-The time gap between production and exchange results in an intertemporal transfer [^3] A similar idea was introduced by Keynes under the notion of effective demand in his book [Key36]..
+The time gap between production and exchange results in an intertemporal transfer [^3] A similar idea was introduced by Keynes under the notion of effective demand in his book [Keynes, 1936]..
 The production-possibility set $\mathscr{Y}^{i_{s}}_{t}$ of producer $i_{s}$} satisfies, at each time $t$:
@@
 Under uncertainty, agents are assumed to form expectations about future events to guide both their current and future decisions.
-In contrast to the rational expectations hypothesis ([Mut61]), which is grounded in conditional expectations, we argue that expectations should instead be formed based on the mode.
+In contrast to the rational expectations hypothesis ([Muth, 1961]), which is grounded in conditional expectations, we argue that expectations should instead be formed based on the mode.
 The rationale is that individuals are more inclined to act upon the most likely outcome rather than an average.
@@
 So $g(z,x)\to 0$ as $\|z\|\to\infty$.
-One could then apply the generalized maximum Theorem 14.4 (as proved in [FKV14] - theorem 1.2), to prove that the mode is an an upper semicontinuous map $\mathbb{M}(x)$, since $F$ is a constant set-valued map.
+One could then apply the generalized maximum Theorem 14.4 (as proved in [Feinberg et al., 2014] - theorem 1.2), to prove that the mode is an an upper semicontinuous map $\mathbb{M}(x)$, since $F$ is a constant set-valued map.
 Moreover, under the strict quasiconcavity of $g(x,\cdot)$, one could end up with a continuous function.
@@
 Instead, anticipation should reflect subjective judgment, bounded rationality, and memory constraints.
-In particular, agents are assumed to base their forecasts on the most likely outcome—captured by the mode—which aligns more closely with Keynesian notions of "animal spirits" [Key36].
+In particular, agents are assumed to base their forecasts on the most likely outcome—captured by the mode—which aligns more closely with Keynesian notions of "animal spirits" [Keynes, 1936].
   • The time gap between production decisions and exchange leads to inherent time inconsistency.
@@
 Theorem 14.1.
-Berge’s Maximum Theorem[Ber63]
+Berge’s Maximum Theorem[Berge, 1963]
 Let $H$ and $\Theta$ be topological spaces, $f:H\times\Theta\to{\mathbb{R}}$ a continuous function, and $\mathcal{F}:\Theta\to 2^{H}$ a correspondence such that:
@@
 Corollary 14.2.
-[Sun96]
+[Sundaram, 1996]
 Let $f:H\times\Theta\to{\mathbb{R}}$ be continuous, and $\mathcal{F}:\Theta\to 2^{H}$ be continuous and compact valued. Define $v$ and $S$ as in Theorem 14.1.
@@
 Theorem 14.3.
-[LRS19]
+[Laraki et al., 2019]
 If a game $G=\{A^{i},\mathcal{F}^{i},u^{i}\}_{1\leqslant i\leqslant n}$ is compact, continuous and quasi-concave, then its set of Nash equilibria is a non-empty and compact subset of $\prod_{i=1}^{n}A^{i}$.
@@
 Definition 14.1.
-$K$-inf-compact functions [FKV14]
+$K$-inf-compact functions [Feinberg et al., 2014]
 Let $f\,:\,H\times\Theta\to{\mathbb{R}}$ and $U$ an open set of $H\times\Theta$. Consider the level sets
@@
 Definition 14.2.
-Compactly generated spaces[FKV14]
+Compactly generated spaces[Feinberg et al., 2014]
 A topological space $X$ is compactly generated if it satisfies the following property: each set $A\subset X$ is closed in $X$ if $A\cap K$ is closed in $K$ for each compact $K$ of $X$.
 Theorem 14.4.
-[FKV14]
+[Feinberg et al., 2014]
 Assume that:
@@
 Then the value function $v:\Theta\to{\mathbb{R}}$ is continuous and the solution multifunction $S(\theta)=\arg\max_{x\in\mathcal{F}(\theta)}f(x,\theta)$ is upper semicontinuous and compact-valued.
+Declarations
+Conflict of interest
+The author declares that they have no conflict of interest.
+Authors’ contributions
+The author carried out the study conception and design. Material preparation, data generation and analysis were performed by the author. The manuscript was written by the author.
+Funding
+No funding was received for conducting this study.
+Availability of data and materials
+All computations were carried out with Wolfram Mathematica.
+Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work the author used ChatGPT in order to improve the text.
+After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.